### v2

# Acylindrical hyperbolicity and the centers of Artin groups that are not free of infinity

Abstract.
Charney and Morris-Wright showed acylindrical hyperbolicity of Artin groups of infinite type associated with graphs that are not joins, by studying clique-cube complexes and the actions on them. The authors developed their study and clarified when acylindrical hyperbolicity holds for Artin groups of infinite type associated with graphs that are not cones. In this paper, we introduce reduced clique-cube complexes. By using them, we show acylindrical hyperbolicity of irreducible Artin groups associated with graphs that are not cliques, that is, irreducible Artin groups that are not free of infinity. As an application, we see that the centers of such Artin groups are finite, and that actually they are trivial in many cases. Also as a corollary, we see that irreducible Artin groups of infinite type and of type FC are acylindrically hyperbolic.
Key words and phrases:
Artin group, Acylindrical hyperbolicity, WPD contracting element, CAT(0)-cube complex, Center
The first author is supported by JSPS KAKENHI Grant Number 20K14311, and JST, ACT-X Grant Number JPMJAX200A, Japan.
The second author is supported by JSPS KAKENHI Grant Number 20K03590.
1. Introduction
Artin groups, also called Artin-Tits groups, have been widely studied since their introduction by Tits [36]. Let $\Gamma$ be a finite simple graph with the vertex set $V(\Gamma)$ and the edge set $E(\Gamma)$. Each edge $e$ has two end vertices, which we denote as $s_{e}$ and $t_{e}$. We suppose that any edge $e$ is labeled by an integer $\mu(e)\geq 2$. The Artin group $A_{\Gamma}$ associated with $\Gamma$ is defined by the following presentation
(1.1) $A_{\Gamma}=\langle V(\Gamma)\mid\underbrace{s_{e}t_{e}s_{e}t_{e}\cdots}_{\text{length $\mu(e)$}}=\underbrace{t_{e}s_{e}t_{e}s_{e}\cdots}_{\text{length $\mu(e)$}}\quad\text{$\left(e\in E(\Gamma)\right)$ }\rangle.$
Here, $\Gamma$ is called a defining graph of $A_{\Gamma}$ and the rank of $A_{\Gamma}$ is defined as #$\#V(\Gamma)$. Free abelian groups, free groups, and braid groups are typical examples of Artin groups. For $U\subset V(\Gamma)$, we denote the subgroup generated by $U$ as $A_{U}$, and call such a subgroup a standard parabolic subgroup. Adding the relation $v^{2}=1$ for all $v\in V(\Gamma)$ to (1.1) produces the associated Coxeter group $W_{\Gamma}$. Let $\Gamma^{t}$ be the Coxeter graph defined from the presentation of $W_{\Gamma}$. That is, $\Gamma^{t}$ is a finite simple graph with edge labels in $\mathbb{Z}_{\geq 2}\cup\{\infty\}$, which is gained from $\Gamma$ by deleting all edges with label $2$ and adding an edge with label $\infty$ between every pair of vertices that are not connected by an edge in $\Gamma$. Although $\Gamma^{t}$ is used in research related to Coxeter groups and traditional treatments of Artin groups, we mainly use $\Gamma$ in accordance with many recent studies on Artin groups in geometric group theory. We use the associated Coxeter graph $\Gamma^{t}$ as an aid in this paper. See Figures 1 and 2.
Figure 1. The defining graphs $\Gamma=\Gamma_{0}\ast\Gamma_{1}$ and the associated Coxeter graphs $\Gamma^{t}$. In each $\Gamma$, $\Gamma_{0}$ is the maximal clique factor and $\Gamma_{1}$ is an indecomposable factor with two vertices.
Figure 2. The defining graphs $\Gamma=\Gamma_{0}\ast\Gamma_{1}\ast\Gamma_{2}$ and the associated Coxeter graphs $\Gamma^{t}$. In each $\Gamma$, $\Gamma_{0}$ is the maximal clique factor and $\Gamma_{1}$ and $\Gamma_{2}$ are indecomposable factors with two vertices.
In terms of the properties of $W_{\Gamma}$ or $\Gamma^{t}$, we can define several important classes of Artin groups. The Artin group $A_{\Gamma}$ is said to be of spherical type if $W_{\Gamma}$ is finite. Otherwise, it is said to be of infinite type. The Artin group $A_{\Gamma}$ is said to be irreducible if $W_{\Gamma}$ is irreducible, that is, $\Gamma^{t}$ is connected. If $\Gamma^{t}$ does not contain edges labeled by $\infty$, then we say $A_{\Gamma}$ is free of infinity. By definition, $A_{\Gamma}$ is free of infinity if and only if $\Gamma$ is a clique. In this paper, we focus on Artin groups which are not free of infinity, that is, Artin groups associated with graphs that are not cliques.
In general, compared with Artin groups, Coxeter groups are relatively well understood. For example, it is well-known that an infinite Coxeter group $W_{\Gamma}$ is irreducible if and only if $W_{\Gamma}$ cannot be directly decomposed into two nontrivial subgroups ([32] and [36]). However, it is unclear whether $A_{\Gamma}$ is irreducible if and only if $A_{\Gamma}$ cannot be directly decomposed into two nontrivial subgroups. We remark that many basic questions for Artin groups remain open (refer to [12] and [20]).
In order to study Artin groups, it is often effective to investigate hidden nonpositively curved or negatively curved properties of them. In this paper, we consider the following problem [23, Conjecture B].
Problem 1.1.
Are irreducible Artin groups of infinite type acylindrically hyperbolic?
Acylindrical hyperbolicity is an important property in geometric group theory. We can find various applications of acylindrical hyperbolicity in [16], [33], and [34] for example.
We note that reducible Artin groups are not acylindrically hyperbolic, since acylindrically hyperbolic groups cannot be directly decomposed into two infinite subgroups ([33, Corollary 7.3]). We also note that Artin groups with infinite centers are not acylindrically hyperbolic, since acylindrically hyperbolic groups do not permit infinite centers ([33, Corollary 7.3]). Although irreducible Artin groups of spherical type have infinite cyclic center ([4] and [17]) and thus are not acylindrically hyperbolic, the central quotients for them of rank at least $2$ are acylindrically hyperbolic (see [2], [24], [5] for braid groups, and [10] for the general case).
Many affirmative partial answers for Problem 1.1 are known. Indeed, the following irreducible Artin groups of infinite type are known to be acylindrically hyperbolic:
  • Right-angled Artin groups ([11], [30]).
  • Two-dimensional Artin groups such that the associated Coxeter groups are hyperbolic ([31]).
  • Artin groups of XXL-type ([23]).
  • Artin groups of type FC such that the defining graphs have diameter greater than $2$ ([15]). Here, an Artin group is of type FC if every standard parabolic subgroup associated with a clique in the defining graph is of spherical type.
  • Artin groups that are known to be CAT(0) groups according to the result of Brady and McCammond ([6] and [28]).
  • Euclidean Artin groups ([9]).
  • Two-dimensional Artin groups, that is, Artin groups such that every triangle with three vertices $v_{1},v_{2},v_{3}$ of the defining graphs satisfies $\frac{1}{\mu((v_{1},v_{2}))}+\frac{1}{\mu((v_{2},v_{3}))}+\frac{1}{\mu((v_{3},v_{1}))}\leq 1$ ([42]).
  • Artin groups associated with graphs that are not joins ([14]).
  • Artin groups associated with graphs that are not cones ([29]).
  • Relatively extra-large-type Artin groups with mild conditions ([21]).
  • Even Artin groups of type FC and Artin group of type FC that visually splits over a standard parabolic subgroup of spherical type ([13]).
As a general strategy for Problem 1.1, we consider the case that the defining graph is a clique and the case that the defining graph is not a clique separately. Note that Artin groups in the latter case are automatically of infinite type. In this paper we give the affirmative answer for Problem 1.1 for the latter case. In fact, our main theorem is the following:
Theorem 1.2.
Let $A_{\Gamma}$ be an Artin group associated with $\Gamma$. Suppose that $\Gamma$ is not a clique. Then, the following are equivalent:
  1. (1)
    $A_{\Gamma}$ is irreducible;
  2. (2)
    $A_{\Gamma}$ has a WPD contracting element with respect to the isometric action on the reduced clique-cube complex;
  3. (3)
    $A_{\Gamma}$ is acylindrically hyperbolic;
  4. (4)
    $A_{\Gamma}$ is virtually directly indecomposable, that is, it cannot be decomposed as a direct product of two infinite subgroups.
From Theorem 1.2, we find that many irreducible Artin groups of infinite type are acylindrically hyperbolic, e.g., the Artin groups associated with the defining graphs in Figures 1 and 2. We also apply Theorem 1.2 to Artin groups of infinite type and of type FC.
Corollary 1.3.
Let $A_{\Gamma}$ be an irreducible Artin group of infinite type and of type FC. Then $A_{\Gamma}$ is acylindrically hyperbolic.
Note that, when the defining graph of type FC is a clique, the Artin group is of spherical type.
We have an application to the so-called center conjecture, which claims that the center of any irreducible Artin group of infinite type is trivial.
Corollary 1.4.
Let $A_{\Gamma}$ be an irreduicble Artin group associated with $\Gamma$ that is not a clique. Then the center $Z(A_{\Gamma})$ is finite. In addition, $Z(A_{\Gamma})$ is contained in $Z(A_{\Gamma_{0}})$ where $\Gamma_{0}$ is the maximal clique factor of $\Gamma$. In particular, if $Z(A_{\Gamma_{0}})$ is trivial or torsion-free, then $Z(A_{\Gamma})$ is trivial.
A result related to the last part of Corollary 1.4 can be found in [27]. See Remark 4.6.
Also we have a corollary for direct products.
Corollary 1.5.
Let $A_{\Gamma}$ be an irreduicble Artin group associated with $\Gamma$ that is not a clique. Then if $A_{\Gamma_{0}}$ is directly indecomposable or no nontrivial finite subgroup which is a factor of a direct decompositon, then $A_{\Gamma}$ is directly indecomposable.
Example 1.6.
Let $A_{\Gamma}$ be an irreduicble Artin group associated with $\Gamma$ that is not a clique. Suppose that $\Gamma_{0}$ consists of three vertices at most. See Figures 1 and 2. Then $Z(A_{\Gamma_{0}})$ is known to be trivial or torsion-free, and thus $Z(A_{\Gamma})$ is trivial by Corollary 1.4. Also $A_{\Gamma_{0}}$ is directly indecomposable or no nontrivial finite subgroup which is a factor of a direct decompositon, and thus $A_{\Gamma}$ is directly indecomposable by Corollary 1.5.
Remark 1.7.
Goldsborough and Vaskou, who are the authors of [22], pointed us out that results in [22] and Theorem 1.2 induce acylindrical hyperbolicity of random Artin groups. The following is based on their remark. By referring to [22] for some definitions, we can write down the claim precisely as follows: an Artin group picked at random relatively to any non-decreasing divergent function $f:\mathbb{N}\to\mathbb{N}$ is asymptotically almost surely acylindrically hyperbolic. Indeed the case of $f(N)\succ N^{\frac{3}{2}}$ and the case of $f(N)\prec N^{1-\alpha}$ for some $10<\alpha<$ 1 are covered in Corollaries 3.4 and 4.3 in [22], respectively. The missing case was unknown (see Question 1.6 in [22]), but now it is resolved as follows: we consider the case of $f(N)\prec N^{2}$, which contains the missing case. By Lemma 2.7, Theorem 2.9 (2) and Lemma 2.10 in [22], an Artin group picked at random relatively to $f$ asymptotically almost surely is irreducible and not free of infinity. Thus it is acylindrically hyperbolic by Theorem 1.2.
In this paper, for the proof of Theorem 1.2, we develop the argument in Charney and Morris-Wright [14] and the authors [29]. Charney and Morris-Wright [14] showed acylindrical hyperbolicity of Artin groups of infinite type associated with graphs that are not joins, by studying clique-cube complexes, which are CAT(0) cube complexes, and the isometric actions on them. They constructed a WPD contracting element of such an Artin group with respect to the isometric action on the clique-cube complex. The authors [29] developed their study and treated Artin groups associated with graphs that are not cones. To treat the case of Artin groups associated with graphs that are not cliques, we need further development and more delicate arguments. Indeed, we introduce the reduced clique-cube complex. It is a convex subcomplex of the clique-cube complex, and works more appopriately in our case than the clique-cube complex. Combining with careful discussions on the local geometry of the cube complexes, we construct a candidate of a WPD contracting element with respect to the action on the reduced clique-cube complex. We also apply some algebraic results on standard parabolic subgroups of Artin groups when we show that the candidate of a WPD contracting element is actually WPD contracting.
The remainder of this paper is organized as follows. Section 2 contains some preliminaries. Section 2.1 is for preliminaries regarding acylindrically hyperbolic groups, WPD contracting elements, and CAT(0)-cube complexes. Section 2.2 presents preliminaries on defining graphs of Artin groups. Section 2.3 presents facts on standard parabolic subgroups of Artin groups and prove a lemma on intersections of standard parabolic subgroups and their conjugates in Artin groups. Section 2.4 treats clique-cube complexes. Section 2.5 deals with the definition of joins of graphs and some properties of them. In Section 3, we introduce reduced clique-cube complexes and study the local geometry. In Section 4, we give a proof of Theorem 1.2. Our main task is to construct a candidate WPD contracting element and show that it really is a WPD contracting element. Also we show Corollaries 1.4 and 1.5.
2. preliminaries
2.1. Acylindrical hyperbolicity, WPD contracting elements, and CAT(0)-cube complexes
We collect some definitions and properties related to acylindrical hyperbolicity, WPD contracting elements, and CAT(0)-cube complexes that will be used later in the paper. See [18] and the references therein for details.
First, we recall the definition of acylindrical hyperbolicity (see [33]).
Definition 2.1.
A group $G$ is acylindrically hyperbolic if it admits an isometric action on a hyperbolic space $Y$ that is non-elementary (i.e., with an infinite limit set) and acylindrical (i.e., for every $D\geq 0$, there exist some $R,N\geq 0$ such that, for all $y_{1},y_{2}\in Y$, $d_{Y}(y_{1},y_{2})\geq R$ implies #$\#\{g\in G|\ d_{Y}(y_{1},g(y_{1})),d_{Y}(y_{2},g(y_{2}))\leq D\}\leq N$).
Next, we recall the definition of a WPD contracting element.
Definition 2.2.
Let a group $G$ act isometrically on a metric space $X$. For $\gamma\in G$, we say that:
  • $\gamma$ is WPD if, for every $D\geq 0$ and $x\in X$, there exists some $M\geq 1$ such that #$\#\{g\in G|\ d_{X}(x,g(x)),d_{X}(\gamma^{M}(x),g\gamma^{M}(x))\leq D\}<\infty$;
  • $\gamma$ is contracting if $\gamma$ is loxodromic, that is, there exists $x_{0}\in X$ such that $\mathbb{Z}\to X;n\mapsto\gamma^{n}(x_{0})$ is a quasi-isometry onto the image $\gamma^{\mathbb{Z}}x_{0}:=\{\gamma^{n}(x_{0})|n\in\mathbb{Z}\}$, and $\gamma^{\mathbb{Z}}x_{0}$ is contracting, that is, there exists $B\geq 0$ such that the diameter of the nearest-point projection of any ball that is disjoint from $\gamma^{\mathbb{Z}}x_{0}$ onto $\gamma^{\mathbb{Z}}x_{0}$ is bounded by $B$.
The following is a consequence of [1].
Theorem 2.3.
Let a group $G$ act isometrically on a geodesic metric space $X$. Suppose that $G$ is not virtually cyclic. If there exists a WPD contracting element $\gamma\in G$, then $G$ is acylindrically hyperbolic.
CAT(0) cube complexes are considered as generalized trees in higher dimensions. The following is a precise definition (see [7, p.111]).
Definition 2.4.
A cube complex is a $CW$ complex constructed by gluing together cubes of arbitrary (finite) dimension by isometries along their faces. Furthermore, the cube complex is nonpositively curved if the link of any of its vertices is a simplicial flag complex (i.e., $n+1$ vertices span an $n$-simplex if and only if they are pairwise adjacent), and CAT(0) if it is nonpositively curved and simply connected.
Definition 2.5.
Let $X$ be a CAT(0) cube complex. We define an equivalence relation for the edges of $X$ as the transitive closure of the relation identifying two parallel edges of a square. For an equivalence class, a hyperplane is defined as the union of the midcubes transverse to the edges belonging to the equivalence class. Then, for any edge belonging to the equivalence class, the hyperplane is said to be dual to the edge.
For a hyperplane $J$, we denote the union of the cubes intersecting $J$ as $N(J)$, that is, the smallest subcomplex of $X$ containing $J$. We denote the union of the cubes not intersecting $J$ as $X\!\setminus\!\!\setminus J$, that is, the largest subcomplex of $X$ not intersecting $J$.
See [37] for the following.
Theorem 2.6.
Let $X$ be a CAT(0) cube complex and $J$ be a hyperplane. Then, $X\!\setminus\!\!\setminus J$ has exactly two connected components.
The two connected components of $X\!\setminus\!\!\setminus J$ are often denoted as $J^{+}$ and $J^{-}$, respectively.
The following is given in [29, Definition 2.7].
Definition 2.7.
Let $X$ be a CAT(0) cube complex. For two vertices $x$ and $x^{\prime}$ in $X$, we call a sequence of hyperplanes $P_{1},\ldots,P_{M}$ a sequence of separating hyperplanes from $x$ to $x^{\prime}$ if the sequence satisfies
$x\in P_{1}^{-},P_{1}^{+}\supsetneq P_{2}^{+}\supsetneq\cdots\supsetneq P_{M-1}^{+}\supsetneq P_{M}^{+}\ni x^{\prime}$
for some connected components $P_{i}^{+}$ of $X\!\setminus\!\!\setminus P_{i}$ for all $i\in\{1,\ldots,M\}$.
For two hyperplanes $J$ and $J^{\prime}$ in $X$, we call a sequence of hyperplanes $P_{1},\ldots,P_{M}$ a sequence of separating hyperplanes from $J$ to $J^{\prime}$ if the sequence satisfies
$J^{+}\supsetneq P_{1}^{+}\supsetneq P_{2}^{+}\supsetneq\cdots\supsetneq P_{M-1}^{+}\supsetneq P_{M}^{+}\supsetneq J^{\prime+}$
for some connected components $J^{+}$ of $X\!\setminus\!\!\setminus J$, $J^{\prime+}$ of $X\!\setminus\!\!\setminus J^{\prime}$, and $P_{i}^{+}$ of $X\!\setminus\!\!\setminus P_{i}$ for all $i\in\{1,\ldots,M\}$.
Remark 2.8.
When $P_{1},\ldots,P_{M}$ is a sequence of separating hyperplanes from $J$ to $J^{\prime}$, for two vertices $x\in J^{-}\cup N(J)$ and $x^{\prime}\in J^{\prime+}\cup N(J^{\prime})$, $P_{1},\ldots,P_{M}$ is a sequence of separating hyperplanes from $x$ to $x^{\prime}$.
The following is part of [18, Theorem 3.3], and is used in the proof of Theorem 1.2.
Theorem 2.9.
Let a group $G$ act isometrically on a CAT(0) cube complex $X$. Then, $\gamma\in G$ is a WPD contracting element if there exist two hyperplanes $J$ and $J^{\prime}$ satisfying the following:
  1. (i)
    $J$ and $J^{\prime}$ are strongly separated, that is, no hyperplane can intersect both $J$ and $J^{\prime}$;
  2. (ii)
    $\gamma$ skewers $J$ and $J^{\prime}$, that is, we have connected components $J^{+}$ of $X\!\setminus\!\!\setminus J$ and $J^{\prime+}$ of $X\!\setminus\!\!\setminus J^{\prime}$ such that $\gamma^{n}(J^{+})\subsetneq J^{\prime+}\subsetneq J^{+}$ for some $n\in\mathbb{N}$;
  3. (iii)
    $\mathrm{stab}(J)\cap\mathrm{stab}(J^{\prime})$ is finite, where $\mathrm{stab}(J)=\{g\in G\ |\ g(J)=J\}$ and $\mathrm{stab}(J^{\prime})=\{g\in G\ |\ g(J^{\prime})=J^{\prime}\}$.
2.2. Defining graphs of Artin groups
We present a precise description of the defining graph of an Artin group and introduce some related graphs.
Let $V$ be a finite set. Denote the diagonal set as $\mathrm{diag}(V\times V):=\{(v,w)\in V\times V|\ v=w\}$. We consider the involution on the off-diagonal set
$\iota:V\times V\setminus\mathrm{diag}(V\times V)\ni(v,w)\mapsto(w,v)\in V\times V\setminus\mathrm{diag}(V\times V).$
Any $e\in V\times V\setminus\mathrm{diag}(V\times V)$ is often presented as $(s_{e},t_{e})$. Then, for any $e\in V\times V\setminus\mathrm{diag}(V\times V)$, we have $s_{\iota(e)}=t_{e}$ and $t_{\iota(e)}=s_{e}$. We take a symmetric map
$\tilde{\mu}:V\times V\setminus\mathrm{diag}(V\times V)\to\mathbb{Z}_{\geq 2}\cup\{\infty\}.$
Here, ‘symmetric’ means that $\tilde{\mu}\circ\iota=\tilde{\mu}$ is satisfied. Then, we have
$V\times V\setminus\mathrm{diag}(V\times V)=\bigsqcup_{m\in\mathbb{Z}_{\geq 2}\cup\{\infty\}}\tilde{\mu}^{-1}(m).$
We now have a finite simple labeled graph $\Gamma$ with the vertex set $V(\Gamma)=V$, the edge set $E(\Gamma)=\bigsqcup_{m\in\mathbb{Z}_{\geq}2}\tilde{\mu}^{-1}(m)$, and the labeling $\mu:=\tilde{\mu}|_{E(\Gamma)}$. The Artin group $A_{\Gamma}$ associated with $\Gamma$ is then defined by the presentation (1.1) and $\Gamma$ is called the defining graph of $A_{\Gamma}$.
For convenience, we define two other finite simple graphs $\Gamma^{c}$ and $\Gamma^{t}$ as follows. $\Gamma^{c}$ is the finite simple graph with the vertex set $V(\Gamma^{c})=V$ and the edge set $E(\Gamma^{c})=\tilde{\mu}^{-1}(\infty)$. This is the so-called complement graph of $\Gamma$. $\Gamma^{t}$, the Coxeter graph, is the finite simple graph with the vertex set $V(\Gamma^{t})=V$ and the edge set $E(\Gamma^{t})=\bigsqcup_{m\in\mathbb{Z}_{\geq 3}\cup\{\infty\}}\tilde{\mu}^{-1}(m)$.
2.3. Properties of Artin groups
We collect and show some properties of Artin groups. Let $A_{\Gamma}$ be an Artin group associated with a defining graph $\Gamma$ with the vertex set $V(\Gamma)=V$ and the edge set $E(\Gamma)=E$, as in Section 2.2. Lemma 2.11 is essentially used in the proof of Theorem 1.2.
By study of van der Lek (see [35], [41]), the following are well-known, which often used without mention in this paper.
Lemma 2.10.
  1. (1)
    For any subset $U\subset V$, the subgroup of $A_{\Gamma}$ generated by $U$, called a standard parabolic subgroup, is itself an Artin group associated with the full subgraph of $\Gamma$ spanned by $U$. (We denote this subgroup as $A_{U}$. When $U$ is empty, we define $A_{\emptyset}=\{1\}$.)
  2. (2)
    For any subsets $U,U^{\prime}\subset V$, $A_{U}\cap A_{U^{\prime}}=A_{U\cap U^{\prime}}$.
The following is a consequence of a nontrivial result by Godelle [19].
Lemma 2.11.
Take an element $v\in V$ and a subset $V^{\prime}\subset V$. Suppose that $v\not\in V^{\prime}$. Set $V^{\prime\prime}:=\{v^{\prime\prime}\in V^{\prime}\mid\mu(v,v^{\prime\prime})=2\}$. Then $A_{V^{\prime}}\cap vA_{V^{\prime}}v^{-1}=A_{V^{\prime\prime}}$. In particular for any $v^{\prime}\in V^{\prime}\setminus V^{\prime\prime}$, $A_{V^{\prime}}\cap vA_{V^{\prime}}v^{-1}\subset A_{V^{\prime}\setminus\{v^{\prime}\}}$.
Proof.
Clearly we have $A_{V^{\prime\prime}}\subset A_{V^{\prime}}\cap vA_{V^{\prime}}v^{-1}$. By noting that $v$ is $(V^{\prime},V^{\prime})$-reduced in the Coxeter group $W_{\Gamma}$ and applying [19, Proposition 3.3] to our setting, we have some $V^{\prime\prime\prime}\subset V^{\prime}$ such that $vA_{V^{\prime}}v^{-1}\cap A_{V^{\prime}}=A_{V^{\prime\prime\prime}}$. Hence $V^{\prime\prime}\subset V^{\prime\prime\prime}$. Assume that $V^{\prime\prime}\subsetneq V^{\prime\prime\prime}$. Take $v^{\prime}\in V^{\prime\prime\prime}\setminus V^{\prime\prime}$. Then $v^{\prime}\in A_{V^{\prime\prime\prime}}=vA_{V^{\prime}}v^{-1}\cap A_{V^{\prime}}$ and thus $v^{-1}v^{\prime}v\in A_{V^{\prime}}$. By $v^{-1}v^{\prime}v\in A_{\{v,v^{\prime}\}}$ and $\{v,v^{\prime}\}\cap V^{\prime}=\{v^{\prime}\}$, we have $v^{-1}v^{\prime}v\in A_{\{v,v^{\prime}\}}\cap A_{V^{\prime}}=A_{\{v^{\prime}\}}$. Therefore we have $m\in\mathbb{Z}$ such that $v^{-1}v^{\prime}v=(v^{\prime})^{m}$. Since the surjection $A_{\{v,v^{\prime}\}}\to A_{\{v^{\prime}\}}$ defined by $v,v^{\prime}\mapsto v^{\prime}$ gives $v^{-1}v^{\prime}v\mapsto v^{\prime}$ and $(v^{\prime})^{m}\mapsto(v^{\prime})^{m}$, we see $m=1$. Hence we have $v^{-1}v^{\prime}v=v^{\prime}$ in $A_{\Gamma}$. This means $\mu(v,v^{\prime})=2$. This contradicts $v^{\prime}\notin V^{\prime\prime}$. Hence $V^{\prime\prime\prime}=V^{\prime\prime}$. Therefore $vA_{V^{\prime}}v^{-1}\cap A_{V^{\prime}}=A_{V^{\prime\prime}}$. ∎
2.4. Clique-cube complexes and actions on them
We consider clique-cube complexes and the actions on them following [14]. Let $A_{\Gamma}$ be an Artin group associated with a defining graph $\Gamma$ with the vertex set $V(\Gamma)=V$ and the edge set $E(\Gamma)=E$, as in Section 2.2. We say that $U$ spans a clique in $\Gamma$ if any two elements of $U$ are joined by an edge in $\Gamma$.
Definition 2.12 ([14, Definition 2.1]).
Consider the set
$\Delta_{\Gamma}=\{U\subset V|\ U\text{ spans a clique in }\Gamma,\text{ or }U=\emptyset\}.$
The clique-cube complex $C_{\Gamma}$ is the cube complex whose vertices (i.e., 00-dimensional cubes) are cosets $gA_{U}$ ($g\in A_{\Gamma},U\in\Delta_{\Gamma}$), where two vertices $gA_{U}$ and $hA_{U^{\prime}}$ are joined by an edge (i.e., a $1$-dimensional cube) in $C_{\Gamma}$ if and only if $gA_{U}\subset hA_{U^{\prime}}$ and $U$ and $U^{\prime}$ differ by a single generator. Note that, in this case, we can always replace $h$ by $g$, that is, $hA_{U^{\prime}}=gA_{U^{\prime}}$. More generally, two vertices $gA_{U}$ and $gA_{U^{\prime}}$ with $gA_{U}\subset gA_{U^{\prime}}$ span a #$\#(U^{\prime}\setminus U)$-dimensional cube $[gA_{U},gA_{U^{\prime}}]$ in $C_{\Gamma}$.
The group $A_{\Gamma}$ acts on the clique-cube complex $C_{\Gamma}$ by left multiplication, $h\cdot gA_{U}=(hg)A_{U}$. This action preserves the cubical structure and is isometric. The action is also co-compact with a fundamental domain $\bigcup_{U\in\Delta_{\Gamma}}[A_{\emptyset},A_{U}]$, where $[A_{\emptyset},A_{U}]$ is a #$\#U$-dimensional cube spanned by two vertices $A_{\emptyset}$ and $A_{U}$ in $C_{\Gamma}$. However, the action is not proper. In fact, the stabilizer of a vertex $gA_{U}$ is the conjugate subgroup $gA_{U}g^{-1}$, so all vertices except translations of $A_{\emptyset}$ have infinite stabilizers. We also note that $C_{\Gamma}$ is not a proper metric space because it contains infinite valence vertices. Additionally, $C_{\Gamma}$ has infinite diameter if and only if $\Gamma$ itself is not a clique.
Remark 2.13 ([14, Section 2]).
Each edge in $C_{\Gamma}$ can be labeled with a generator in $V$. For example, the edge between $gA_{U}$ and $gA_{U\sqcup\{v\}}$ is labeled by $v$. Any two parallel edges in a cube have the same label, so we can also label the hyperplane dual to such an edge by $v$ and say that such a hyperplane is of $v$-type. Every hyperplane of $v$-type is the translation of a hyperplane dual to the edge between $A_{\emptyset}$ and $A_{\{v\}}$. If a hyperplane of $v$-type crosses another hyperplane of $v^{\prime}$-type, then $(v,v^{\prime})\in E(\Gamma)$. In particular, two different hyperplanes of the same type do not cross each other.
We endow a metric on $C_{\Gamma}$ in the usual way as a cube complex. That is, we identify all cubes in $C_{\Gamma}$ with Euclidean unit cubes and give $C_{\Gamma}$ the path-metric.
Theorem 2.14 ([14, Theorem 2.2]).
The clique-cube complex $C_{\Gamma}$ is a complete CAT(0) space.
Note that completeness follows from finite dimensionality of $C_{\Gamma}$ ([7, Theorem 7.19]).
2.5. Joins of graphs
Definition 2.15.
Let $\Lambda\neq\emptyset$ be an index set. The join $\ast_{\alpha\in\Lambda}\Gamma_{\alpha}$ of simple graphs $\Gamma_{\alpha}$ ($\alpha\in\Lambda$) is defined as a simple graph with the vertex set
$V(\ast_{\alpha\in\Lambda}\Gamma_{\alpha}):=\bigsqcup_{\alpha\in\Lambda}V(\Gamma_{\alpha})$
and the edge set
$E(\ast_{\alpha\in\Lambda}\Gamma_{\alpha}):=\bigsqcup_{\alpha\in\Lambda}E(\Gamma_{\alpha})\sqcup\bigsqcup_{\alpha,\beta\in\Lambda,\alpha\neq\beta}\{(v_{\alpha},v_{\beta})\mid v_{\alpha}\in V(\Gamma_{\alpha}),v_{\beta}\in V(\Gamma_{\beta})\}.$
A simple graph $\Gamma$ is said to be decomposable as a join if there exist an index set $\Lambda$ with #$\#\Lambda\geq 2$ and subgraphs $\Gamma_{\alpha}$ ($\alpha\in\Lambda$) of $\Gamma$ such that $\Gamma=\ast_{\alpha\in\Lambda}\Gamma_{\alpha}$. This is called a join decomposition of $\Gamma$ into factors $\Gamma_{\alpha}$ ($\alpha\in\Lambda$). $\Gamma$ is said to be indecomposable as a join if it is not decomposable as a join.
Remark 2.16.
Any simple graph $\Gamma$ is indecomposable as a join if and only if its complement graph $\Gamma^{c}$ is connected.
The following is a well-known fact (See [29, Lemma 3.4]).
Lemma 2.17.
Let $\Gamma$ be a simple graph. Suppose that $\Gamma$ is decomposable as a join. Then, $\Gamma$ has a unique join decomposition into indecomposable factors
$\Gamma=\ast_{\alpha\in\Lambda}\Gamma_{\alpha},$
and the decomposition of $\Gamma^{c}$ into connected components is given by
$\Gamma^{c}=\bigsqcup_{\alpha\in\Lambda}(\Gamma_{\alpha})^{c}.$
Definition 2.18.
Let $\Gamma$ be a simple graph. Suppose that $\Gamma$ is decomposable as a join. For the unique join decomposition into indecomposable factors $\Gamma=\ast_{\alpha\in\Lambda}\Gamma_{\alpha}$, we set $\Lambda_{0}:=\{\alpha\in\Lambda\mid\ |\Gamma_{\alpha}|=1\}$ and $\Gamma_{0}:=\ast_{\alpha\in\Lambda_{0}}\Gamma_{\alpha}$. We call $\Gamma_{0}$ the maximal clique factor of $\Gamma$. Also we set $\Lambda_{\ast}:=\Lambda\setminus\Lambda_{0}$ and $\Gamma_{\ast}:=\ast_{\alpha\in\Lambda_{\ast}}\Gamma_{\alpha}$. We call $\Gamma$ a cone if $\Lambda_{0}\neq\emptyset$.
See Figures 1 and 2. We have $\Gamma=\Gamma_{0}\ast\Gamma_{\ast}$. Clearly $\Gamma$ is a clique if and only if $\Lambda=\Lambda_{0}$.
3. Reduced clique-cube complexes and actions on them
In this section, we introduce reduced clique-cube complexes and study them. Let $A_{\Gamma}$ be an Artin group associated with a defining graph $\Gamma$ with the vertex set $V(\Gamma)=V$ and the edge set $E(\Gamma)=E$, as in Section 2.2. We consider the unique join decomposition $\Gamma=\Gamma_{0}\ast\Gamma_{\ast}=\Gamma_{0}\ast\left(\ast_{1\leq i\leq k}\Gamma_{i}\right)$ where $\Gamma_{0}$ is the maximal clique factor and $\Gamma_{i}$ $(1\leq i\leq k)$ is an indecomposable factor with at least two vertices (see Definition 2.18). We set $V_{i}=V(\Gamma_{i})$.
Definition 3.1.
Let
$\Delta_{\Gamma,\Gamma_{0}}=\{U\subset V\mid U\in\Delta_{\Gamma},\ U\supset V(\Gamma_{0})\},$
where $\Gamma_{0}$ is the maximal clique factor of $\Gamma$. Let $C_{\Gamma,\Gamma_{0}}$ be a subcomplex of $C_{\Gamma}$ spanned by vertices $gA_{U}$ with $g\in A_{\Gamma}$ and $U\in\Delta_{\Gamma,\Gamma_{0}}$.
When $\Gamma$ is not a cone, that is, $\Gamma=\Gamma_{\ast}$, we have $C_{\Gamma,\Gamma_{0}}=C_{\Gamma}$. When $\Gamma$ is a clique, that is, $\Gamma=\Gamma_{0}$, we have $C_{\Gamma,\Gamma_{0}}=C_{\Gamma,\Gamma}=\{A_{\Gamma}\}$. In this paper, we treat the case where $\Gamma$ is not a clique.
We note that the action of $A_{\Gamma}$ on $C_{\Gamma}$ preserves $C_{\Gamma,\Gamma_{0}}$. In the following, we consider the induced action of $A_{\Gamma}$ on $C_{\Gamma,\Gamma_{0}}$.
Lemma 3.2.
$C_{\Gamma,\Gamma_{0}}$ is a connected subcomplex of $C_{\Gamma}$.
Proof.
If $\Gamma$ is a clique, then $C_{\Gamma,\Gamma_{0}}=\{A_{\Gamma}\}$ is clearly connected. Suppose that $\Gamma$ is not a clique. Let $U\in\Delta_{\Gamma,\Gamma_{0}}$. Let $g\in A_{\Gamma}$. First, we show that there exists a path in $C_{\Gamma,\Gamma_{0}}$ from $A_{V(\Gamma_{0})}$ to $gA_{V(\Gamma_{0})}$. Suppose that $g\notin A_{V(\Gamma_{0})}$, since otherwise $A_{V(\Gamma_{0})}=gA_{V(\Gamma_{0})}$. Let $g=v_{1}^{\varepsilon_{1}}\cdots v_{n}^{\varepsilon_{n}}$ where $v_{i}\in V(\Gamma)$, $\varepsilon_{i}\in\{1,-1\}$ for every $1\leq i\leq n$. By induction on $n$, we assume that there exists a path in $C_{\Gamma,\Gamma_{0}}$ from $A_{V(\Gamma_{0})}$ to $v_{1}^{\varepsilon_{1}}\cdots v_{n-1}^{\varepsilon_{n-1}}A_{V(\Gamma_{0})}$. If $v_{n}\in V(\Gamma_{0})$, then
$v_{1}^{\varepsilon_{1}}\cdots v_{n-1}^{\varepsilon_{n-1}}A_{V(\Gamma_{0})}=v_{1}^{\varepsilon_{1}}\cdots v_{n}^{\varepsilon_{n}}A_{V(\Gamma_{0})},$
and the induction proceeds. If $v_{n}\notin V(\Gamma_{0})$, since $V(\Gamma_{0})\cup\{v_{n}\}\in\Delta_{\Gamma,\Gamma_{0}}$, there exists a path
$\left(v_{1}^{\varepsilon_{1}}\cdots v_{n-1}^{\varepsilon_{n-1}}A_{V(\Gamma_{0})},\ v_{1}^{\varepsilon_{1}}\cdots v_{n-1}^{\varepsilon_{n-1}}A_{V(\Gamma_{0})\cup\{v_{n}\}}=gA_{V(\Gamma_{0})\cup\{v_{n}\}},\ gA_{V(\Gamma_{0})}\right)$
from $A_{V(\Gamma_{0})}$ to $gA_{V(\Gamma_{0})}$ in $C_{\Gamma,\Gamma_{0}}$. Next, we show that there exists a path in $C_{\Gamma,\Gamma_{0}}$ from $gA_{V(\Gamma_{0})}$ to $gA_{U}$. Let $U=V(\Gamma_{0})\cup\{u_{1},\ldots,u_{m}\}$ where $u_{i}\neq u_{j}$ for $1\leq i\neq j\leq m$. Then there exists a path
$\left(gA_{V(\Gamma_{0})},gA_{V(\Gamma_{0})\cup\{u_{1}\}},gA_{V(\Gamma_{0})\cup\{u_{1},u_{2}\}},\ldots,gA_{U}\right)$
in $C_{\Gamma,\Gamma_{0}}$ from $gA_{V(\Gamma_{0})}$ to $gA_{U}$. Therefore, it follows that $C_{\Gamma,\Gamma_{0}}$ is a connected subcomplex of $C_{\Gamma}$. ∎
Definition 3.3.
Let $X$ be a geodesic space and $W$ be a subset of $X$. $W$ is convex in $X$ if for every $w_{1},w_{2}\in W$, any geodesic from $w_{1}$ to $w_{2}$ is contained in $W$.
Definition 3.4.
Let $L$ be a simplicial complex and $L^{\prime}$ be a subcomplex of $L$. $L^{\prime}$ is full if any simplex of $L$ whose vertices are in $L^{\prime}$ is contained in $L^{\prime}$.
Theorem 3.5 (for example [38]).
Let $X$ be a finite dimensional CAT(0) cube complex and $W$ be a connected subcomplex. Then $W$ is convex in $X$ if and only if for every vertex $v$ of $W$, the link $L^{\prime}$ of $v$ in $W$ is a full subcomplex of the link $L$ of $v$ in $X$.
Lemma 3.6.
$C_{\Gamma,\Gamma_{0}}$ is convex in $C_{\Gamma}$. In particular, $C_{\Gamma,\Gamma_{0}}$ is a complete CAT(0) space as a subspace of $C_{\Gamma}$.
Proof.
We show that $C_{\Gamma,\Gamma_{0}}$ is convex in $C_{\Gamma}$. Let $gA_{U}$ be a vertex of $C_{\Gamma,\Gamma_{0}}$. Let $L$ be the link of $gA_{U}$ in $C_{\Gamma}$ and let $L^{\prime}$ be the link of $gA_{U}$ in $C_{\Gamma,\Gamma_{0}}$. According to Theorem 3.5, it is sufficient to show that $L^{\prime}$ is a full subcomplex of $L$. Let $\Delta$ be a $d$-dimensional simplex of $L$ whose vertices are in $L^{\prime}$. Let $C_{\Delta}$ be a $(d+1)$-dimensional cube of $C_{\Gamma}$ which corresponds to $\Delta$. Let $gA_{U_{1}},\ldots,gA_{U_{d+1}}$ be vertices of $C_{\Delta}$ that are connected to $gA_{U}$ by a $1$-dimensional cube of $C_{\Gamma}$. Note that $C_{\Delta}=[gA_{\cap_{i=1}^{d+1}U_{i}},gA_{\cup_{i=1}^{d+1}U_{i}}]$. Since $gA_{U_{i}}\in C_{\Gamma,\Gamma_{0}}$, $U_{i}\supset V(\Gamma_{0})$ for every $i$. Therefore, $\cap_{i=1}^{d+1}U_{i}\supset V(\Gamma_{0})$. Since $C_{\Delta}$ is contained in $C_{\Gamma,\Gamma_{0}}$, $\Delta$ is contained in $L^{\prime}$. It follows that $L^{\prime}$ is a full subcomplex of $L$.
Moreover, $C_{\Gamma,\Gamma_{0}}$ is connected (Lemma 3.2), thus it is CAT(0) as a subspace of a CAT(0) space $C_{\Gamma}$. Completeness follows from the fact that $C_{\Gamma,\Gamma_{0}}$ is a finite dimensional cube complex ([7, Theorem 7.19]). ∎
The following is a generalization of [14, Lemma 3.2].
Lemma 3.7.
The action of $A_{\Gamma}$ on $C_{\Gamma,\Gamma_{0}}$ is a minimal action. That is, for every $x\in C_{\Gamma,\Gamma_{0}}$, the convex hull of the orbit of $x$, $\mathrm{Hull}(orb(x))$, is $C_{\Gamma,\Gamma_{0}}$.
Proof.
When $\Gamma$ is a clique, $C_{\Gamma,\Gamma_{0}}=\{A_{\Gamma_{0}}\}$. In this case, the lemma is trivial. We consider the case where $\Gamma$ is not a clique. We fix a join decomposition
$\Gamma=\Gamma_{0}\ast\Gamma_{1}\ast\cdots\ast\Gamma_{k},$
where $\Gamma_{0}$ is the maximal clique factor of $\Gamma$ and $\Gamma_{i}$ are indecomposable factors with at least two vertices for $i\geq 1$. We note that $\Gamma_{1}\ast\cdots\ast\Gamma_{k}\neq\emptyset$ since $\Gamma$ is not a clique.
Let $x\in C_{\Gamma,\Gamma_{0}}$. Let $C$ be a cube of $C_{\Gamma,\Gamma_{0}}$ such that $x\in C$. Without loss of generality, we set $C=[A_{U},A_{T}]$ where $U\subset T$, $U,T\in\Delta_{\Gamma,\Gamma_{0}}$. By induction on $|T\setminus U|$, we show that $\mathrm{Hull}(orb(x))=C_{\Gamma,\Gamma_{0}}$.
When $|T\setminus U|=0$, then $x=A_{U}$. It is sufficient to show that $\mathrm{Hull}(orb(A_{U}))=\mathrm{Hull}(orb(A_{V(\Gamma_{0})}))$. Indeed, if $\mathrm{Hull}(orb(A_{U}))=\mathrm{Hull}(orb(A_{V(\Gamma_{0})}))$ for every $U\in\Delta_{\Gamma,\Gamma_{0}}$, then $\mathrm{Hull}(orb(A_{V(\Gamma_{0})}))$ contains every vertex of $C_{\Gamma,\Gamma_{0}}$ and thus equals to $C_{\Gamma,\Gamma_{0}}$.
First, we show that $\mathrm{Hull}(orb(A_{U}))\supset\mathrm{Hull}(orb(A_{V(\Gamma_{0})}))$. If $V(\Gamma_{0})\neq U$, we take $u\in U\setminus V(\Gamma_{0})$. There exists $i\in\{1,\ldots,k\}$ such that $u\in\Gamma_{i}$. Since $\Gamma_{i}$ is indecomposable as $\Gamma_{i}=u\ast\Gamma^{\prime}_{i}$, there exists a vertex $s\neq u$ of $\Gamma_{i}$ such that $s$ and $u$ are not connected by an edge in $\Gamma_{i}$. Let
$U^{\prime}=\{v\in V_{\Gamma}\mid\text{ $v$ and $s$ are connected by an edge in$\Gamma$}\}\cap U.$
We note that ${U^{\prime}}$ does not contain $u$ and $s$, and that ${U^{\prime}}\subsetneq U$. We also note that ${U^{\prime}}\cup\{s\}$ spans a clique, but $U\cup\{s\}$ does not. Then,
$A_{U},\ A_{U^{\prime}},\ A_{{U^{\prime}}\cup\{s\}},\ sA_{U^{\prime}},\ sA_{U}$
is a local geodesic, and thus a geodesic, since $C_{\Gamma,\Gamma_{0}}$ is CAT(0) (Lemma 3.6). This geodesic is contained in $\mathrm{Hull}(orb(A_{U}))$, since two endpoints are in $\mathrm{Hull}(orb(A_{U}))$. Therefore, $A_{U^{\prime}}\in\mathrm{Hull}(orb(A_{U}))$, and thus $\mathrm{Hull}(orb(A_{U^{\prime}}))\subset\mathrm{Hull}(orb(A_{U}))$. If $U^{\prime}=V(\Gamma_{0})$, then we have nothing to do. If ${U^{\prime}}\supsetneq V(\Gamma_{0})$, we repeat the same argument until we get $\mathrm{Hull}(orb(A_{U}))\supset\mathrm{Hull}(orb(A_{V(\Gamma_{0})}))$.
Next, we show that $\mathrm{Hull}(orb(A_{U}))\subset\mathrm{Hull}(orb(A_{V(\Gamma_{0})}))$. If $V(\Gamma_{0})\neq U$, we take $u\in U\setminus V(\Gamma_{0})$. Then
$A_{U\setminus\{u\}},\ A_{U},\ uA_{U\setminus\{u\}}$
is a geodesic. It follows that $\mathrm{Hull}(orb(A_{U}))\subset\mathrm{Hull}(orb(A_{U\setminus\{u\}}))$. If $U\setminus\{u\}=V(\Gamma_{0})$, then the proof is finished. If $U\setminus\{u\}\supsetneq V(\Gamma_{0})$, we repeat the same argument until we get $\mathrm{Hull}(orb(A_{U}))\subset\mathrm{Hull}(orb(A_{V(\Gamma_{0})}))$.
When $|T\setminus U|>0$, there exists $s\in T\setminus U$. Then $C\cap sC$ contains a face spanned by $A_{s}$ and $A_{T}$. The geodesic from $x$ to $sx$ passes through this face. Let $y$ be the intersection point. Then $y\in\mathrm{Hull}(orb(x))$, and thus $\mathrm{Hull}(orb(y))\subset\mathrm{Hull}(orb(x))$. By induction, $\mathrm{Hull}(orb(y))=C_{\Gamma,\Gamma_{0}}$. ∎
The following is a generalization of [29, Proposition 4.8].
Proposition 3.8.
Suppose that $\Gamma$ is not a clique. Then, a finite normal subgroup of $A_{\Gamma}$ is in $A_{\Gamma_{0}}$. In particular, any finite subgroup of the center of $A_{\Gamma}$ is in $A_{\Gamma_{0}}$. Also, if $A_{\Gamma}$ is isomorphic to a direct product $A_{1}\times A_{2}$ and $A_{1}$ is finite, then $A_{1}$ is in $A_{\Gamma_{0}}$ and satisfies $A_{\Gamma_{0}}=A_{1}\times(A_{\Gamma_{0}}\cap A_{2})$.
Proof.
Let $N$ be a finite normal subgroup of $A_{\Gamma}$. Set
$Fix(N):=\{x\in C_{\Gamma,\Gamma_{0}}\mid nx=x\text{ for any }n\in N\}.$
Because $N$ is finite and $C_{\Gamma,\Gamma_{0}}$ is a complete CAT(0) space, we have $Fix(N)\neq\emptyset$. Take any $x\in Fix(N)$. Then, $A_{\Gamma}x\subset Fix(N)$. Indeed, the normality of $N$ implies that, for any $g\in A_{\Gamma}$ and $n\in N$, there exists $n^{\prime}\in N$ such that $ng=gn^{\prime}$. Therefore, we have $ngx=gn^{\prime}x=gx$. Because $C_{\Gamma,\Gamma_{0}}$ is CAT(0), $Fix(N)$ is convex. Hence, we have $Hull(A_{\Gamma}x)\subset Fix(N)$. By Lemma 3.7, we have $Hull(A_{\Gamma}x)=C_{\Gamma,\Gamma_{0}}$. Hence, $Fix(N)=C_{\Gamma,\Gamma_{0}}$. In particular, $A_{\Gamma_{0}}\in Fix(N)$. Hence, $N$ must be in $A_{\Gamma_{0}}$. ∎
Finally, we state a lemma related to the local geometry of reduced clique-cube complexes. The following is a nontrivial variant for $C_{\Gamma,\Gamma_{0}}$ of [29, Lemma 5.2] for $C_{\Gamma}$.
Lemma 3.9.
Let $s,t\in V(\Gamma_{\ast})$ belong to mutually different indecomposable factors of $\Gamma_{\ast}$. Suppose that we have $d\in\mathbb{N}$ and mutually different $w_{1},\ldots,w_{d-1}\in V(\Gamma_{0})$ such that $(w_{i-1},w_{i})\in E(\Gamma)$ with label greater than $2$ for any $i\in\{1,\ldots,d\}$ where we set $w_{0}:=s$ and $w_{d}:=t$. Set $\tau:=w_{0}w_{1}\cdots w_{d}\in A_{\Gamma}$. Let $U\in\Delta_{\Gamma,\Gamma_{0}}$ with $s,t\in U$. Then, there exists no square in $C_{\Gamma,\Gamma_{0}}$ containing both edges $[A_{U},A_{U\setminus\{s\}}]$ and $[A_{U},\tau A_{U\setminus\{t\}}]$, that is, there exists no $g\in A_{U\setminus\{s\}}$ such that $gA_{U\setminus\{t\}}=\tau A_{U\setminus\{t\}}$.
See Figure 3.
Figure 3. Squares around $A_{U}$.
Proof.
Denote by $\bar{\tau}$ the image of $\tau$ by the natural projection $A_{\Gamma}\to W_{\Gamma}$. Since $w_{0},w_{1},\ldots,w_{d}$ are mutually different, the word $w_{0}w_{1}\cdots w_{d}$ is a reduced expression of $\bar{\tau}$. Moreover this is a unique word representing $\bar{\tau}$ because $\bar{\tau}$ is regarded as a Coxeter element in $W_{\{w_{0},w_{1},\ldots,w_{d}\}}$, $w_{i-1}$ and $w_{i}$ cannot commute for any $i\in\{1,\ldots,d\}$ (see for example [39]).
Assume that we have $g\in A_{U\setminus\{s\}}$ such that $gA_{U\setminus\{t\}}=\tau A_{U\setminus\{t\}}$. Then we have $\tau\in A_{U\setminus\{s\}}A_{U\setminus\{t\}}$ and thus $\bar{\tau}\in W_{U\setminus\{s\}}W_{U\setminus\{t\}}$. Take $a_{1},\ldots,a_{d_{1}}\in U\setminus\{s\}$ and $b_{1},\ldots,b_{d_{2}}\in U\setminus\{t\}$ such that the word $a_{1}\cdots a_{d_{1}}b_{1}\cdots b_{d_{2}}$ represents $\bar{\tau}$. Then by the deletion property for Coxeter groups (see for example [3]), the word $a_{1}\cdots a_{d_{1}}b_{1}\cdots b_{d_{2}}$ contains a reduced expression for $\tau$ as a subword. In such any subword, the letter $t$ precedes $s$. This contradicts the fact that $w_{0}=s$ precedes $w_{d}=t$ in the unique reduced expresssion $w_{0}w_{1}\cdots w_{d}$ of $\bar{\tau}$. ∎
4. Proof of Theorem 1.2
Let $A_{\Gamma}$ be an Artin group associated with a defining graph $\Gamma$ with the vertex set $V(\Gamma)=V$ and the edge set $E(\Gamma)=E$, as in Section 2.2. In this section, we suppose that $\Gamma$ is not a clique. We show our main theorem (Theorem 1.2), that is, the following are equivalent:
  1. (1)
    $A_{\Gamma}$ is irreducible, that is, $\Gamma$ cannot be decomposed as a join of two subgraphs such that all edges between them are labeled by $2$;
  2. (2)
    $A_{\Gamma}$ has a WPD contracting element with respect to the isometric action on the reduced clique-cube complex;
  3. (3)
    $A_{\Gamma}$ is acylindrically hyperbolic;
  4. (4)
    $A_{\Gamma}$ is virtually directly indecomposable, that is, it cannot be decomposed as a direct product of two infinite subgroups.
4.1. Proof of $(2)\Rightarrow(3),(3)\Rightarrow(4),(4)\Rightarrow(1)$ in Theorem 1.2
We show that $(2)\Rightarrow(3),(3)\Rightarrow(4),(4)\Rightarrow(1)$ in Theorem 1.2.
First, $(2)\Rightarrow(3)$ follows from Theorem 2.3.
Next, we show $(3)\Rightarrow(4)$. Let $A_{\Gamma}$ be acylindrically hyperbolic. If $A_{\Gamma}$ is isomorphic to a direct product $A_{1}\times A_{2}$, then either $A_{1}$ or $A_{2}$ is finite by acylindrical hyperbolicity [33, Corollary 7.3].
Finally, because every non-trivial Artin group is infinite, $(4)\Rightarrow(1)$ is clear.
4.2. Proof of $(1)\Rightarrow(2)$ in Theorem 1.2
We give a proof of $(1)\Rightarrow(2)$ in Theorem 1.2. Suppose that $A_{\Gamma}$ is irreducible, that is, $\Gamma^{t}$ is connected. We will construct an element $\gamma\in A_{\Gamma}$ and show that $\gamma$ is a WPD contracting element with respect to the action on the reduced clique-cube complex $C_{\Gamma,\Gamma_{0}}$. When $\Gamma_{0}$ is empty, a WPD contracting element is given in [29]. Our $\gamma$ in this paper is constructed more delicately than in [29].
We consider the unique join decomposition $\Gamma=\Gamma_{0}\ast\Gamma_{\ast}=\Gamma_{0}\ast(\ast_{1\leq i\leq k}\Gamma_{i})$ where $\Gamma_{0}$ is the maximal clique factor and $\Gamma_{i}$ ($1\leq i\leq k$) is an indecomposable factor with at least two vertices (see Definition 2.18). We set $V_{i}=V(\Gamma_{i})$ and $E_{i}=E(\Gamma_{i})$ for each $i\in\{0,1,\ldots,k\}$. Then, for every $i\neq 0$, the complement graph $(\Gamma_{i})^{c}$ of $\Gamma_{i}$ is connected and the complement graph $(\Gamma_{\ast})^{c}$ of $\Gamma_{\ast}$ is the disjoint union of connected components as $(\Gamma_{\ast})^{c}=\bigsqcup_{1\leq i\leq k}(\Gamma_{i})^{c}$ by Lemma 2.17.
We define $Q(\Gamma,\Gamma_{0})$ as a finite simple graph with
  • the vertex set $V(Q(\Gamma,\Gamma_{0}))=\{V_{i}\}_{1\leq i\leq k}$, and
  • the edge set $E(Q(\Gamma,\Gamma_{0}))=\{(V_{i},V_{j}),(V_{j},V_{i})\mid 1\leq i<j\leq k,\\ (\Gamma_{i})^{t}\text{ and }(\Gamma_{j})^{t}\text{ are connected in }(\Gamma_{0}\ast\Gamma_{i}\ast\Gamma_{j})^{t}\}$.
Note that for $i\neq 0$, $(\Gamma_{i})^{t}$ is connected because $(\Gamma_{i})^{c}$ is connected, $V((\Gamma_{i})^{c})=V((\Gamma_{i})^{t})=V_{i}$, and $E((\Gamma_{i})^{c})\subset E((\Gamma_{i})^{t})$. Note that the following are equivalent for different $i,j\in\{1,\ldots,k\}$:
  • $(\Gamma_{i})^{t}$ and $(\Gamma_{j})^{t}$ are connected in $(\Gamma_{0}\ast\Gamma_{i}\ast\Gamma_{j})^{t}$;
  • there exists a path consisting of edges with label greater than $2$ between a vertex of $\Gamma_{i}$ and a vertex of $\Gamma_{j}$ in $\Gamma_{0}\ast\Gamma_{i}\ast\Gamma_{j}$.
Because $A_{\Gamma}$ is irreducible, i.e., $\Gamma^{t}$ is connected, $Q(\Gamma,\Gamma_{0})$ is connected.
We take a spanning tree $T$ of $Q(\Gamma,\Gamma_{0})$. We regard $T$ as a rooted tree with the root $V_{1}$. By trading indices of $V_{2},\ldots,V_{k}$ if necessary, we suppose that $i<j$ only if $V_{i}$ is not farther than $V_{j}$ from $V_{1}$ in $T$.
For each $i,j$ with $i<j$ and $(V_{i},V_{j})\in E(T)$, take a path $p_{i,j}=(w_{i,j,0},w_{i,j,1},\ldots,w_{i,j,d(i,j)})$ consisting of edges with label greater than $2$ such that $w_{i,j,1},\ldots,w_{i,j,d(i,j)-1}\in\Gamma_{0}$ if $d(i,j)\geq 2$, $w_{i,j,0}\in\Gamma_{i}$ and $w_{i,j,d(i,j)}\in\Gamma_{j}$, and set $p_{j,i}=(w_{j,i,0},w_{j,i,1},\ldots,w_{j,i,d(j,i)}):=(w_{i,j,d(i,j)},\ldots,w_{i,j,1},w_{i,j,0})$. For any $i,j\in\{1,\ldots,k\}$ with $(V_{i},V_{j})\in E(T)$, set $s_{i,j}:=w_{i,j,0}$ in $\Gamma_{i}$, $t_{i,j}:=w_{i,j,d(i,j)}$ in $\Gamma_{j}$ and $\tau_{i,j}=w_{i,j,0}w_{i,j,1}\cdots w_{i,j,d(i,j)}$ in $A_{\Gamma}$, thus $s_{j,i}=t_{i,j}$ in $\Gamma_{j}$, $t_{j,i}=s_{i,j}$ in $\Gamma_{i}$ and $\tau_{j,i}=w_{i,j,d(i,j)}\cdots w_{i,j,1}w_{i,j,0}$ in $A_{\Gamma}$.
The following is almost same with [29, Lemma 6.1].
Lemma 4.1.
There exist $n\in\mathbb{N}$, for each $i\in\{1,\ldots,k\}$, a closed path $(v_{i,1},\ldots,v_{i,n},v_{i,n+1})$ with $v_{i,1}=v_{i,n+1}$ on $(\Gamma_{i})^{c}$ passing through every vertex at least once, and for any $i,j\in\{1,\ldots,k\}$ with $(V_{i},V_{j})\in E(T)$, $l(i,j)\in\{1,\dots,n\}$ such that $v_{i,l(i,j)}=s_{i,j}$, $v_{j,l(i,j)}=t_{i,j}$ and $l(j,i)=l(i,j)$.
Proof.
For any $i\in\{1,\ldots,k\}$, we consider the minimum length $n_{i}$ of closed paths on $(\Gamma_{i})^{c}$ passing through every vertex at least once. Set $n=\prod_{1\leq i\leq k}n_{i}$. Then, for any $i\in\{1,\ldots,k\}$, by concatenating $\frac{n}{n_{i}}$ copies of a closed path of length $n_{i}$ on $(\Gamma_{i})^{c}$ passing through every vertex at least once, we have a closed path $(v^{\prime}_{i,1},\ldots,v^{\prime}_{i,n},v^{\prime}_{i,n+1})$ with $v^{\prime}_{i,1}=v^{\prime}_{i,n+1}$ on $(\Gamma_{i})^{c}$ passing through every vertex at least $\frac{n}{n_{i}}$ times (in particular, at least once).
We set $(v_{1,1},\ldots,v_{1,n},v_{1,n+1}):=(v^{\prime}_{1,1},\ldots,v^{\prime}_{1,n},v^{\prime}_{1,n+1})$. For any $j\in\{2,\ldots,k\}$, we define $(v_{j,1},\ldots,v_{j,n},v_{j,n+1})$ inductively as follows. Take $j\in\{2,\ldots,k\}$. Assume that $(v_{i,1},\ldots,v_{i,n},v_{i,n+1})$ is defined for $i\in\{1,\ldots,j-1\}$ with $(V_{i},V_{j})\in E(T)$. Then, we set $l(i,j)$ as the minimum $l$ such that $v_{i,l}=s_{i,j}$. By a cyclic permutation of $v^{\prime}_{j,1},\ldots,v^{\prime}_{j,n}$, we have $v_{j,1},\ldots,v_{j,n}$ such that $v_{j,l(i,j)}=t_{i,j}$. By setting $v_{j,n+1}:=v_{j,1}$, we have $(v_{j,1},\ldots,v_{j,n},v_{j,n+1})$. For any $i,j$ with $1\leq i<j\leq k$ and $(V_{i},V_{j})\in E(T)$, we set $l(j,i):=l(i,j)$. ∎
We take $n\in\mathbb{N}$, $(v_{i,1},\ldots,v_{i,n},v_{i,n+1})$ for each $i\in\{1,\ldots,k\}$, and $l(i,j)\in\{1,\dots,n\}$ for any $i,j\in\{1,\ldots,k\}$ with $(V_{i},V_{j})\in E(T)$ as in Lemma 4.1. For $1\leq l\leq n$, set
$\lambda_{l}:=v_{1,l}v_{2,l}\cdots v_{k,l}.$
For any $i,j\in\{1,\ldots,k\}$ with $(V_{i},V_{j})\in E(T)$, we define $\lambda_{l}(i,j)$ as $\lambda_{l}(i,j):=\lambda_{l}$ if $l\neq l(i,j)$. In addition, we define $\lambda_{l(i,j)}(i,j)$ as
$\lambda_{l(i,j)}(i,j):=\begin{array}[]{lll}\tau_{i,j}v_{1,l(i,j)}v_{2,l(i,j)}\cdots v_{i-1,l(i,j)}v_{i+1,l(i,j)}&&\\ \ \ \cdots v_{j-1,l(i,j)}v_{j+1,l(i,j)}\cdots v_{k,l(i,j)}\end{array}\text{if }i<j,\text{ and}$
$\lambda_{l(i,j)}(i,j):=\begin{array}[]{lll}\tau_{i,j}v_{1,l(i,j)}v_{2,l(i,j)}\cdots v_{j-1,l(i,j)}v_{j+1,l(i,j)}&&\\ \ \ \cdots v_{i-1,l(i,j)}v_{i+1,l(i,j)}\cdots v_{k,l(i,j)}\end{array}\text{if }i>j,$
where we note that $l(j,i)=l(i,j)$. Moreover, we define $\gamma_{\natural}(i,j)$ as
$\gamma_{\natural}(i,j):$ $=\lambda_{1}(i,j)\lambda_{2}(i,j)\cdots\lambda_{n}(i,j)$
$=\lambda_{1}\lambda_{2}\cdots\lambda_{l(i,j)-1}\lambda_{l(i,j)}(i,j)\lambda_{l(i,j)+1}\cdots\lambda_{n}.$
Take a closed path
(4.1) $(V_{i_{1}},\dots,V_{i_{r}},V_{i_{r+1}})$
with $i_{1}=i_{r+1}=1$ on the spanning tree $T$ of $Q(\Gamma,\Gamma_{0})$ passing through every vertex at least once. We define $\gamma_{\natural}$ as
(4.2) $\gamma_{\natural}:=\gamma_{\natural}(i_{1},i_{2})\gamma_{\natural}(i_{2},i_{3})\cdots\gamma_{\natural}(i_{r},i_{r+1}).$
For $h\geq 1$, we set
(4.3) $W(h):=\{w\in V_{0}\mid\min_{v\in V_{\ast}}d_{\Gamma^{t}}(w,v)=h\}.$
Since $\Gamma^{t}$ is connected, we have $\mathfrak{d}\in\mathbb{N}$ such that $W(\mathfrak{d})\neq\emptyset$ and $V_{0}=W(1)\sqcup\cdots\sqcup W(\mathfrak{d})$. We define a map $V_{0}\to\{1,\ldots,k\}\times\{1,\ldots,n\};w\mapsto(i(w),l(w))$ as $(i(w),l(w))$ is the minimum of $(i,l)$ indexing $v_{i,l}\in V_{*}$ such that $d_{\Gamma^{t}}(w,v_{i,l})=\min_{v\in V_{\ast}}d_{\Gamma^{t}}(w,v)$ with respect to the dictionary order on $\{1,\ldots,k\}\times\{1,\ldots,n\}$. In particular for $w\in W(1)$, $(i(w),l(w))$ is the minimum of $(i,l)$ with $\mu(w,v_{i,l})>2$ with respect to the dictionary order on $\{1,\ldots,k\}\times\{1,\ldots,n\}$. We set
(4.4) $W(0):=\{v_{i(w),l(w)}\in V_{\ast}\mid w\in W(1)\},$
which equals $\{v_{i(w),l(w)}\in V_{\ast}\mid w\in V_{0}\}$. Note that for $w\in W(0)$, $i$ and $l$ with $w=v_{i,l}$, $i$ is unique, but $l$ is not unique. Therefore we define for $w\in W(0)$, $(i(w),l(w))$ is the minimum of $(i,l)$ with $w=v_{i,l}$ with respect to the dictionary order on $\{1,\ldots,k\}\times\{1,\ldots,n\}$.
We define a total order $<$ on $W(0)\sqcup V_{0}=W(0)\sqcup\cdots\sqcup W(\mathfrak{d})$ as follows. First we take a total order $<$ on $W(0)$ such that for any $w,w^{\prime}\in W(0)$, $(i(w),l(w))<(i(w^{\prime}),l(w^{\prime}))$ if and only if $w<w^{\prime}$. Next suppose that for $h\geq 1$, $W(0)\sqcup\cdots\sqcup W(h-1)$ is equipped with a total order $<$. Then we define a map
(4.5) $W(h)\to W(h-1);w\mapsto w(h-1)$
as $w(h-1)$ is the minimum of $w^{\prime}\in W(h-1)$ with $\mu(w,w^{\prime})>2$ with respect to the given total order $<$. We take a total order $<$ on $W(h)$ such that for any $w,w^{\prime}\in W(h)$, $w(h-1)\leq w^{\prime}(h-1)$ if $w<w^{\prime}$. Moreover for any $w\in W(0)\sqcup\cdots\sqcup W(h-1)$ and $w^{\prime}\in W(h)$, we set $w<w^{\prime}$. Then $<$ on $W(0)\sqcup\cdots\sqcup W(h-1)$ is extended on $W(0)\sqcup\cdots\sqcup W(h)$. Finally we have the order $<$ on $W(0)\sqcup V_{0}=W(0)\sqcup\cdots\sqcup W(\mathfrak{d})$. For any $0\leq h\leq\mathfrak{d}$ and $w\in W(h)$, we take the path $(w(h),w(h-1),\ldots,w(0))$ such that $w(h)=w$, $w(\mathfrak{j}-1)=w(\mathfrak{j})(\mathfrak{j}-1)$ for any $2\leq\mathfrak{j}\leq h$, and $w(0)=v_{i(w),l(w)}$, where $(i(w),l(w))=(i(w(1)),l(w(1)))$. See Figure 4.
Figure 4. $W(0),\ldots,W(\mathfrak{d})$ in the case of $\mathfrak{d}=3$.
Now we name all elements of $W(0)\sqcup\cdots\sqcup W(\mathfrak{d}-1)=W(0)\sqcup V_{0}\setminus W(\mathfrak{d})$ as
(4.6) $w_{1},\ldots,w_{\mathfrak{m}}$
such that $w_{1}<\cdots<w_{\mathfrak{m}}$ where $\mathfrak{m}:=\#(W(0)\sqcup\cdots\sqcup W(\mathfrak{d}-1))$. For any $\mathfrak{j}\in\{1,\ldots,\mathfrak{m}\}$, we define $\lambda_{l}(\mathfrak{j})$ as $\lambda_{l}(\mathfrak{j}):=\lambda_{l}$ if $l\neq l(w_{\mathfrak{j}})$. In addition, we take $h\in\{0,1,\ldots,\mathfrak{d}-1\}$ with $w_{\mathfrak{j}}\in W(h)$ and define $\lambda_{l(w_{\mathfrak{j}})}(\mathfrak{j})$ as
$\lambda_{l(w_{\mathfrak{j}})}(\mathfrak{j}):=v_{1,l(w_{\mathfrak{j}})}\cdots v_{i(w_{\mathfrak{j}})-1,l(w_{\mathfrak{j}})}v_{i(w_{\mathfrak{j}})+1,l(w_{\mathfrak{j}})}\cdots v_{k,l(w_{\mathfrak{j}})}$
$w_{\mathfrak{j}}(h)w_{\mathfrak{j}}(h-1)\cdots w_{\mathfrak{j}}(1)w_{\mathfrak{j}}(0),$
where we note $w_{\mathfrak{j}}(h)=w_{\mathfrak{j}}$ and $w_{\mathfrak{j}}(0)=v_{i(w_{\mathfrak{j}}),l(w_{\mathfrak{j}})}$. We set
(4.7) $\gamma_{\flat}(\mathfrak{j}):=\lambda_{1}(\mathfrak{j})\cdots\lambda_{n}(\mathfrak{j})=\lambda_{1}\cdots\lambda_{l(w_{\mathfrak{j}})-1}\lambda_{l(w_{\mathfrak{j}})}(\mathfrak{j})\lambda_{l(w_{\mathfrak{j}})+1}\cdots\lambda_{n}.$
We define $\gamma_{\flat}$ as
(4.8) $\gamma_{\flat}:=\gamma_{\flat}(1)\cdots\gamma_{\flat}(\mathfrak{m}).$
By using (4.8) and (4.2), we define
(4.9) $\gamma:=\gamma_{\flat}\gamma_{\natural}=\gamma_{\flat}(1)\cdots\gamma_{\flat}(\mathfrak{m})\gamma_{\natural}(i_{1},i_{2})\gamma_{\natural}(i_{2},i_{3})\cdots\gamma_{\natural}(i_{r},i_{r+1}).$
We set
$\gamma(0):=1,\gamma(1):=\lambda_{1}(1),\gamma(2):=\lambda_{1}(1)\lambda_{2}(1),\ldots,\gamma(n):=\gamma_{\flat}(1),$
for $a\in\{2,\ldots,\mathfrak{m}\}$ and $l\in\{1,\ldots,n\}$,
$\gamma((a-1)n+l):=\gamma_{\flat}(1)\cdots\gamma_{\flat}(a-1)(\lambda_{1}(a)\cdots\lambda_{l}(a)).$
In particular, we have $\gamma(\mathfrak{m}n)=\gamma_{\flat}$. Moreover we set
$\gamma(\mathfrak{m}n+1):=\gamma_{\flat}\lambda_{1}(i_{1},i_{2}),\gamma(\mathfrak{m}n+2):=\gamma_{\flat}\lambda_{1}(i_{1},i_{2})\lambda_{2}(i_{1},i_{2}),\ldots,\gamma(\mathfrak{m}(n+1)):=\gamma_{\flat}\gamma_{\natural}(i_{1},i_{2}),$
for $a\in\{2,\ldots,r\}$ and $l\in\{1,\ldots,n\}$,
$\gamma((\mathfrak{m}+a-1)n+l):=\gamma_{\flat}\gamma_{\natural}(i_{1},i_{2})\cdots\gamma_{\natural}(i_{a-1},i_{a})(\lambda_{1}(i_{a},i_{a+1})\cdots\lambda_{l}(i_{a},i_{a+1})).$
In particular, we have $\gamma((\mathfrak{m}+r)n)=\gamma$.
We confirm the following for convenience:
  1. (1)
    $k$ is the number of indecomposable factors of a unique join decomposition of $\Gamma_{\ast}$;
  2. (2)
    $n$ is the common length of the closed paths on $(\Gamma_{i})^{c}$ for all $i\in\{1,\ldots,k\}$ taken in Lemma 4.1;
  3. (3)
    $r$ is the length of the closed path (4.1) on $T$ taken above.
For $1\leq l\leq n$, set
$U_{l}:=V_{0}\sqcup\{v_{1,l},v_{2,l},\ldots,v_{k,l}\}.$
Then, $U_{l}$ spans a clique containing $\Gamma_{0}$ in $\Gamma$, that is, $U_{l}\in\Delta_{\Gamma,\Gamma_{0}}$. Hence, we have a $k$-dimensional cube $[A_{V_{0}},A_{U_{l}}]$ in $C_{\Gamma,\Gamma_{0}}$. The hyperplanes dual to edges of $[A_{V_{0}},A_{U_{l}}]$ are of $v_{i,l}$-type ($i\in\{1,\ldots,k\}$). We denote such hyperplanes as $H_{i,l}$ ($i\in\{1,\ldots,k\}$).
The following is a variant for $C_{\Gamma,\Gamma_{0}}$ of [29, Lemma 6.2] for $C_{\Gamma}$.
Lemma 4.2.
For $i\in\{1,\ldots,k\}$, $l\in\{1,\ldots,n\}$, $\mathfrak{j}\in\{1,\ldots,\mathfrak{m}\}$, and $a\in\{1,\ldots,r\}$, we have the following:
  1. (1)
    $H_{i,l}\cap H_{i,l+1}=\emptyset$, where we set $H_{i,n+1}:=H_{i,1}$;
  2. (2)
    $H_{i,l}\cap\lambda_{l}H_{i,l}=\emptyset$;
  3. (3)
    $H_{i,l(i_{a},i_{a+1})}\cap\lambda_{l(i_{a},i_{a+1})}(i_{a},i_{a+1})H_{i,l(i_{a},i_{a+1})}=\emptyset$;
  4. (4)
    $H_{i,l(w_{\mathfrak{j}})}\cap\lambda_{l(w_{\mathfrak{j}})}(\mathfrak{j})H_{i,l(w_{\mathfrak{j}})}=\emptyset$;
  5. (5)
    $[A_{V_{0}},A_{U_{l}}]\cap[A_{V_{0}},A_{U_{l+1}}]=\{A_{V_{0}}\}$, where we set $U_{n+1}:=U_{1}$;
  6. (6)
    $[A_{V_{0}},A_{U_{l}}]\cap\lambda_{l}[A_{V_{0}},A_{U_{l}}]=\{A_{U_{l}}\}$;
  7. (7)
    $[A_{V_{0}},A_{U_{l(i_{a},i_{a+1})}}]\cap\lambda_{l(i_{a},i_{a+1})}(i_{a},i_{a+1})[A_{V_{0}},A_{U_{l(i_{a},i_{a+1})}}]=\{A_{U_{l(i_{a},i_{a+1})}}\}$.
  8. (8)
    $[A_{V_{0}},A_{U_{l(w_{\mathfrak{j}})}}]\cap\lambda_{l(w_{\mathfrak{j}})}(\mathfrak{j})[A_{V_{0}},A_{U_{l(w_{\mathfrak{j}})}}]=\{A_{U_{l(w_{\mathfrak{j}})}}\}$.
Proof.
(1) $H_{i,l}$ and $H_{i,l+1}$ are of $v_{i,l}$-type and $v_{i,l+1}$-type, respectively. Note that $v_{i,l}\neq v_{i,l+1}$ implies $H_{i,l}\neq H_{i,l+1}$. Because $(v_{i,l},v_{i,l+1})\notin E(\Gamma)$, we have $\{v_{i,l},v_{i,l+1}\}\notin\Delta_{\Gamma}$, and thus $H_{i,l}\cap H_{i,l+1}=\emptyset$.
(2) Because $[A_{U_{l}\setminus\{v_{i,l}\}},A_{U_{l}}]\subset N(H_{i,l})$, we have $[\lambda_{l}A_{U_{l}\setminus\{v_{i,l}\}},\lambda_{l}A_{U_{l}}]\subset\lambda_{l}N(H_{i,l})$. Note that $\lambda_{l}A_{U_{l}}=A_{U_{l}}$ and $\lambda_{l}A_{U_{l}\setminus\{v_{i,l}\}}=v_{1,l}\cdots v_{i,l}A_{U_{l}\setminus\{v_{i,l}\}}$. Now, assume that $H_{i,l}\cap\lambda_{l}H_{i,l}\neq\emptyset$. Because $H_{i,l}$ and $\lambda_{l}H_{i,l}$ are of $v_{i,l}$-type, we see that $H_{i,l}=\lambda_{l}H_{i,l}$ (see Remark 2.13). Then, we have
$[\lambda_{l}A_{U_{l}\setminus\{v_{i,l}\}},\lambda_{l}A_{U_{l}}]=[v_{1,l}\cdots v_{i,l}A_{U_{l}\setminus\{v_{i,l}\}},A_{U_{l}}]\subset N(H_{i,l}).$
Hence, we obtain $A_{U_{l}\setminus\{v_{i,l}\}}=v_{1,l}\cdots v_{i,l}A_{U_{l}\setminus\{v_{i,l}\}}$. This means that $v_{i,l}\in A_{U_{l}\setminus\{v_{i,l}\}}$. However, from Lemma 2.10 (2), $\{v_{i,l}\}\cap(U_{l}\setminus\{v_{i,l}\})=\emptyset$ implies that $A_{\{v_{i,l}\}}\cap A_{U_{l}\setminus\{v_{i,l}\}}=A_{\emptyset}=\{1\}$. This contradicts $v_{i,l}\neq 1$ in $A_{\Gamma}$.
(3) When $i\neq i_{a},i_{a+1}$, by the same argument as in (2), we see that
$H_{i,l(i_{a},i_{a+1})}\cap\lambda_{l(i_{a},i_{a+1})}(i_{a},i_{a+1})H_{i,l(i_{a},i_{a+1})}=\emptyset.$
Now, assume that $i=i_{a}$ or $i=i_{a+1}$ and
$H_{i,l(i_{a},i_{a+1})}\cap\lambda_{l(i_{a},i_{a+1})}(i_{a},i_{a+1})H_{i,l(i_{a},i_{a+1})}\neq\emptyset.$
Note that $\lambda_{l(i_{a},i_{a+1})}(i_{a},i_{a+1})A_{U_{l(i_{a},i_{a+1})}}=A_{U_{l(i_{a},i_{a+1})}}$ and
$\lambda_{l(i_{a},i_{a+1})}(i_{a},i_{a+1})A_{U_{l(i_{a},i_{a+1})}\setminus\{v_{i,l(i_{a},i_{a+1})}\}}=\tau_{i_{a},i_{a+1}}A_{U_{l(i_{a},i_{a+1})}\setminus\{v_{i,l(i_{a},i_{a+1})}\}}.$
Because $H_{i,l(i_{a},i_{a+1})}$ and $\lambda_{l(i_{a},i_{a+1})}H_{i,l(i_{a},i_{a+1})}$ are of $v_{i,l(i_{a},i_{a+1})}$-type, we see that $H_{i,l(i_{a},i_{a+1})}=\lambda_{l(i_{a},i_{a+1})}H_{i,\lambda_{l(i_{a},i_{a+1})}}$. Then, we have
$\begin{split}[\lambda_{l(i_{a},i_{a+1})}A_{U_{l(i_{a},i_{a+1})}\setminus\{v_{i,l(i_{a},i_{a+1})}\}},\lambda_{l(i_{a},i_{a+1})}A_{U_{l(i_{a},i_{a+1})}}]\\ =[\tau_{i_{a},i_{a+1}}A_{U_{l(i_{a},i_{a+1})}\setminus\{v_{i,{l(i_{a},i_{a+1})}}\}},A_{U_{l(i_{a},i_{a+1})}}]\subset N(H_{i,{l(i_{a},i_{a+1})}}).\end{split}$
Hence, we have
$\tau_{i_{a},i_{a+1}}A_{U_{l(i_{a},i_{a+1})}\setminus\{v_{i,l(i_{a},i_{a+1})}\}}=A_{U_{l(i_{a},i_{a+1})}\setminus\{v_{i,l(i_{a},i_{a+1})}\}}.$
This contradicts Lemma 3.9.
(4) is shown by a similar way as in (2).
Parts (5), (6), (7) and (8) follow from (1), (2), (3) and (4), respectively. ∎
The following is a variant for $C_{\Gamma,\Gamma_{0}}$ of [29, Lemma 6.3] for $C_{\Gamma}$, and its proof is same.
Lemma 4.3.
For $a\in\{1,\ldots,r\}$, we have
$H_{i_{a},l(i_{a},i_{a+1})}\cap\lambda_{l(i_{a},i_{a+1})}(i_{a},i_{a+1})H_{i_{a+1},l(i_{a},i_{a+1})}=\emptyset.$
Noting Lemma 4.2, for any $i\in\{1,\ldots,k\}$, we define a sequence of hyperplanes
$\ldots,J_{i,-1},J_{i,0},J_{i,1},\ldots,J_{i,2(\mathfrak{m}+r)n},J_{i,2(\mathfrak{m}+r)n+1},\ldots,$
a sequence of $k$-dimensional cubes
$\ldots,K_{-1},K_{0},K_{1},\cdots,K_{2(\mathfrak{m}+r)n},K_{2(\mathfrak{m}+r)n+1},\ldots,$
and a sequence of vertices of $C_{\Gamma,\Gamma_{0}}$
$\ldots,w_{-1},w_{0},w_{1},\ldots,w_{2(\mathfrak{m}+r)n},w_{2(\mathfrak{m}+r)n+1},\ldots$
as follows. First, for $a\in\{1,\ldots,\mathfrak{m}+r\}$ and $l\in\{1,\dots,n\}$, we define
$J_{i,2((a-1)n+l)-1}$ $:=\gamma((a-1)n+l-1)H_{i,l},$
$J_{i,2((a-1)n+l)}$ $:=\gamma((a-1)n+l)H_{i,l},$
$K_{2((a-1)n+l)-1}$ $:=\gamma((a-1)n+l-1)[A_{V_{0}},A_{U_{l}}],$
$K_{2((a-1)n+l)}$ $:=\gamma((a-1)n+l)[A_{V_{0}},A_{U_{l}}],$
$w_{2((a-1)n+l)-1}$ $:=\gamma((a-1)n+l-1)A_{U_{l}},$
$w_{2((a-1)n+l)}$ $:=\gamma((a-1)n+l)A_{V_{0}},$
where we note that both $J_{i,2((a-1)n+l)-1}$ and $J_{i,2((a-1)n+l)}$ are of the same $v_{i,l}$-type.
Second, for any $b\in\{1,\dots,2(\mathfrak{m}+r)n\}$ and $c\in\mathbb{Z}$, we set
$J_{i,2(\mathfrak{m}+r)nc+b}:=\gamma^{c}J_{i,b},K_{2(\mathfrak{m}+r)nc+b}:=\gamma^{c}K_{b},w_{2(\mathfrak{m}+r)nc+b}:=\gamma^{c}w_{b}.$
Then, we have
$K_{2(\mathfrak{m}+r)nc+b}=[w_{2(\mathfrak{m}+r)nc+b-1},w_{2(\mathfrak{m}+r)nc+b}].$
Additionally, we have the two connected components $J_{i,2(\mathfrak{m}+r)nc+b}^{-}$ and $J_{i,2(\mathfrak{m}+r)nc+b}^{+}$ such that
$C_{\Gamma}\!\setminus\!\!\setminus J_{i,2(\mathfrak{m}+r)nc+b}=J_{i,2(\mathfrak{m}+r)nc+b}^{-}\sqcup J_{i,2(\mathfrak{m}+r)nc+b}^{+},$
$w_{2(\mathfrak{m}+r)nc+b-1}\in J_{i,2(\mathfrak{m}+r)nc+b}^{-},w_{2(\mathfrak{m}+r)nc+b}\in J_{i,2(\mathfrak{m}+r)nc+b}^{+}.$
Then, Lemma 4.2 implies that, for any $i\in\{1,\ldots,k\}$,
$\cdots\subsetneq J_{i,-1}^{-}\subsetneq J_{i,0}^{-}\subsetneq J_{i,1}^{-}\subsetneq\cdots\subsetneq J_{i,2(\mathfrak{m}+r)n}^{-}\subsetneq J_{i,2(\mathfrak{m}+r)n+1}^{-}\subsetneq\cdots,$
$\cdots\supsetneq J_{i,-1}^{+}\supsetneq J_{i,0}^{+}\supsetneq J_{i,1}^{+}\supsetneq\cdots\supsetneq J_{i,2(\mathfrak{m}+r)n}^{+}\supsetneq J_{i,2(\mathfrak{m}+r)n+1}^{+}\supsetneq\cdots.$
Note that
$J_{i,0}^{-}\not\ni w_{0}=A_{V_{0}}\in J_{i,1}^{-},$
$J_{i,2(\mathfrak{m}+r)n}^{+}\ni w_{2(\mathfrak{m}+r)n}=\gamma A_{V_{0}}\not\in J_{i,2(\mathfrak{m}+r)n+1}^{+}.$
We have the following, which is a variant for $C_{\Gamma,\Gamma_{0}}$ of [29, Lemma 6.4] for $C_{\Gamma}$.
Lemma 4.4.
The set $\{J_{i,d}\}_{i\in\{1,\ldots,k\},d\in\{1,\ldots,2(\mathfrak{m}+r)n\}}$ is the set of all hyperplanes separating $w_{0}=A_{V_{0}}$ and $w_{2(\mathfrak{m}+r)n}=\gamma A_{V_{0}}$.
For the proof, we consider a path from $w_{0}=A_{V_{0}}$ to $w_{2(\mathfrak{m}+r)n}=\gamma A_{V_{0}}$ that diagonally penetrates each of the cubes $K_{1},\ldots,K_{2(\mathfrak{m}+r)n}$ in order. The set of all hyperplanes intersecting the path is $\{J_{i,d}\}_{i\in\{1,\ldots,k\},d\in\{1,\ldots,2(\mathfrak{m}+r)n\}}$.
We now state the final lemma required for the proof of Theorem 1.2, where we recall Definition 2.7 and Remark 2.8. The following is a variant for $C_{\Gamma,\Gamma_{0}}$ of [29, Lemma 6.5] for $C_{\Gamma}$.
Lemma 4.5.
  1. (1)
    For any $i\in\{1,\ldots,k\}$, the sequence of hyperplanes
    $J_{i,1},\ldots,J_{i,2(\mathfrak{m}+r)n}$
    is a sequence of separating hyperplanes from $J_{i,0}$ to $J_{i,2(\mathfrak{m}+r)n+1}$, where $J_{i,0}$ and $J_{i,2(\mathfrak{m}+r)n+1}$ are of $v_{i,n}$-type and $v_{i,1}$-type, respectively (in particular, from $w_{0}=A_{V_{0}}$ to $w_{2(\mathfrak{m}+r)n}=\gamma A_{V_{0}}$).
  2. (2)
    The sequence of hyperplanes
    $J_{i_{1},1},J_{i_{1},2},\ldots,J_{i_{1},2\mathfrak{m}n},$
    $J_{i_{1},2\mathfrak{m}n+1},J_{i_{1},2\mathfrak{m}n+2},\ldots,J_{i_{1},2\mathfrak{m}n+2l(i_{1},i_{2})-1},$
    $J_{i_{2},2\mathfrak{m}n+2l(i_{1},i_{2})},J_{i_{2},2\mathfrak{m}n+2l(i_{1},i_{2})+1},\ldots,J_{i_{2},2\mathfrak{m}n+2n},$
    $J_{i_{2},2\mathfrak{m}n+2n+1},J_{i_{2},2\mathfrak{m}n+2(n+1)},\ldots,J_{i_{2},2\mathfrak{m}n+2(n+l(i_{2},i_{3}))-1},$
    $\cdots$
    $J_{i_{r},2\mathfrak{m}n+2((r-2)n+l(i_{r-1},i_{r}))},J_{i_{r},2\mathfrak{m}n+2((r-2)n+l(i_{r-1},i_{r}))+1},\ldots,J_{i_{r},2\mathfrak{m}n+2(r-1)n},$
    $J_{i_{r},2\mathfrak{m}n+2(r-1)n+1},J_{i_{r},2\mathfrak{m}n+2((r-1)n+1)},\ldots,J_{i_{r},2\mathfrak{m}n+2((r-1)n+l(i_{r},i_{r+1}))-1},$
    $J_{i_{r+1},2\mathfrak{m}n+2((r-1)n+l(i_{r},i_{r+1}))},J_{i_{r+1},2\mathfrak{m}n+2((r-1)n+l(i_{r},i_{r+1}))+1},\ldots,J_{i_{r+1},2(\mathfrak{m}+r)n}$
    is a sequence of separating hyperplanes from $J_{1,0}$ to $J_{1,2(\mathfrak{m}+r)n+1}$, where $J_{1,0}$ and $J_{1,2(\mathfrak{m}+r)n+1}$ are of $v_{1,n}$-type and $v_{1,1}$-type, respectively (in particular, from $w_{0}=A_{V_{0}}$ to $w_{2(\mathfrak{m}+r)n}=\gamma A_{V_{0}}$). Here note that $i_{1}=1$. (See Figure 5.)
    Moreover, the sequence contains a hyperplane of $v_{i}$-type for any $i\in\{1,\ldots,k\}$ and $v_{i}\in V(\Gamma_{i})$.
Figure 5. Part of a sequence of hyperplanes for the case where $\Gamma=\Gamma_{0}\ast\Gamma_{1}\ast\Gamma_{2}$.
Proof.
(1) The assertion is clear from parts (1), (2), and (3) of Lemma 4.2.
(2) Parts (1), (2), (3) and (4) of Lemma 4.2 and Lemma 4.3 imply that the sequence specified in the assertion is a sequence of separating hyperplanes from $J_{i_{1},0}$ to $J_{i_{r+1},2(\mathfrak{m}+r)n+1}$. Note that $i_{1}=1$ and $i_{r+1}=1$ by definition.
We show that the sequence contains a hyperplane of $v_{i}$-type for any $i\in\{1,\ldots,k\}$ and $v_{i}\in V(\Gamma_{i})$. Recall that $(V_{i_{1}},\dots,V_{i_{r}},V_{i_{r+1}})$ (4.1) is a closed path on the tree $T$. Thus, if an edge $(V_{i},V_{j})\in E(T)$ is contained in the closed path, then so is the inverse edge $(V_{j},V_{i})$. In addition, because $T$ is a spanning tree of $Q(\Gamma)$ and the closed path $(V_{i_{1}},\dots,V_{i_{r}},V_{i_{r+1}})$ passes through every vertex at least once, any $i\in\{1,\ldots,k\}$, $V_{i}$ is contained in the closed path as a vertex. Additionally, recall that $T$ has the root $V_{1}$ and that $i<j$ only if $V_{i}$ is not farther than $V_{j}$ from $V_{1}$ in $T$.
Now, take any $i\in\{1,\ldots,k\}$. We consider the two cases of $i=1$ and $i\neq 1$.
First, suppose that $i=1$. Note that $i_{1}=1$ and set $j=i_{2}$. Take $a\in\{1,2,\ldots,r\}$ such that $a\neq 1$, $i_{a}=1$, and $i_{a-1}=j$. Then, the sequence in the assertion contains the two subsequences
$J_{i_{1},2\mathfrak{m}n+1},J_{i_{1},2\mathfrak{m}n+2},\ldots,J_{i_{1},2\mathfrak{m}n+2l(i_{1},i_{2})-1},$
which are of $v_{1,1}$-type, $v_{1,1}$-type, $\ldots$, $v_{1,l(1,j)}$-type, and
$J_{i_{a},2\mathfrak{m}n+2((a-2)n+l(i_{a-1},i_{a}))},J_{i_{a},2\mathfrak{m}n+2((a-2)n+l(i_{a-1},i_{a}))+1},\ldots,J_{i_{a},2\mathfrak{m}n+2(a-1)n},$
which are of $v_{1,l(j,1)}$-type, $v_{1,l(j,1)+1}$-type, $\ldots$, $v_{1,n}$-type. Note that $l(j,1)=l(1,j)$ by Lemma 4.1. Take any vertex $v_{1}\in V(\Gamma_{1})$. Then, there exists some $l\in\{1,\ldots,n\}$ such that $v_{1,l}=v_{1}$ by Lemma 4.1. Hence, the sequence in the assertion contains a hyperplane of $v_{1}$-type.
Next, suppose that $i\neq 1$. Take the smallest $a\in\{1,\ldots,r\}$ such that $i_{a}=i$. Because $i_{1}=1$ and $i\neq 1$, we have $a\neq 1$. We set $j=i_{a-1}$. Then, we have $a^{\prime}\in\{1,\ldots,r\}$ such that $a\leq a^{\prime}$, $i_{a^{\prime}}=i$, and $i_{a^{\prime}+1}=j$. The sequence contains the two subsequences
$J_{i_{a},2\mathfrak{m}n+2((a-2)n+l(i_{a-1},i_{a}))},J_{i_{a},2\mathfrak{m}n+2((a-2)n+l(i_{a-1},i_{a}))+1},\ldots,J_{i_{a},2\mathfrak{m}n+2(a-1)n},$
which are of $v_{i,l(j,i)}$-type, $v_{i,l(j,i)+1}$-type, $\ldots$, $v_{i,n}$-type, and
$J_{i_{a^{\prime}},2\mathfrak{m}n+2(a^{\prime}-1)n+1},J_{i_{a^{\prime}},2\mathfrak{m}n+2((a^{\prime}-1)n+1)},\ldots,J_{i_{a^{\prime}},2\mathfrak{m}n+2((a^{\prime}-1)n+l(i_{a^{\prime}},i_{a^{\prime}+1}))-1},$
which are of $v_{i,1}$-type, $v_{i,1}$-type, $\ldots$, $v_{i,l(j,i)}$-type. Note that $l(j,i)=l(i,j)$ by Lemma 4.1. Take any vertex $v_{i}\in V(\Gamma_{i})$. Then, there exists some $l\in\{1,\ldots,n\}$ such that $v_{i,l}=v_{i}$ by Lemma 4.1. Hence, the sequence in the assertion contains a hyperplane of $v_{i}$-type. ∎
We can now complete the proof of Theorem 1.2.
Proof of $(1)\Rightarrow(2)$ in Theorem 1.2.
We show $(1)\Rightarrow(2)$ in Theorem 1.2. Consider $\gamma\in A_{\Gamma}$ defined by (4.9) and hyperplanes
$J:=J_{1,0}\text{ and }J^{\prime}:=J_{1,2(\mathfrak{m}+r)n+1},$
which are of $v_{1,n}$-type and $v_{1,1}$-type, respectively. We will confirm conditions (i), (ii), and (iii) in Theorem 2.9.
(i) $\gamma$ skewers $(J,J^{\prime})$. Indeed, it is clear that
$J^{+}\supsetneq(\gamma^{-1}(J^{\prime+})\supsetneq\gamma(J^{+})\supsetneq)J^{\prime+}\supsetneq\gamma^{2}(J^{+}).$
(ii) We show that $J$ and $J^{\prime}$ are strongly separated. Take any hyperplane $H$ with $J\cap H\neq\emptyset$. When $H$ is of $v_{i}$-type for some $i$ and $v_{i}\in V(\Gamma_{i})$, take a hyperplane $H^{\prime}$ of $v_{i}$-type separating $A_{V_{0}}$ and $\gamma A_{V_{0}}$ such that $J\cap H^{\prime}=\emptyset$ by part (2) of Lemma 4.5. Then, $H\cap H^{\prime}=\emptyset$ by Remark 2.13. Because $J\subset H^{\prime-},H^{\prime+}\supset J^{\prime}$, $J\cap H\neq\emptyset$ and $H\cap H^{\prime}=\emptyset$, we have $J^{\prime}\cap H=\emptyset$.
(iii) We show $Stab(J)\cap Stab(J^{\prime})=\{1\}$. We devide the proof into three parts: (iii-1) we show that $Stab(J)\cap Stab(J^{\prime})\subset A_{V_{0}}$; (iii-2) we show that $Stab(J)\cap Stab(J^{\prime})\subset\gamma A_{V_{0}}\gamma^{-1}$; (iii-3) we show that $Stab(J)\cap Stab(J^{\prime})=\{1\}$ using (iii-1) and (iii-2).
(iii-1) Note that for any $i\in\{1,\ldots,k\}$ and any $v_{i}\in V(\Gamma_{i})$, we have at least one sequence of separating hyperplanes $P^{\prime}_{1},\ldots,P^{\prime}_{M^{\prime}}$ from $A_{V_{0}}$ to $\gamma A_{V_{0}}$ such that $P^{\prime}_{1}$ is of $v_{i}$-type and $P^{\prime}_{M^{\prime}}$ is of $v_{1,n}$-type. For example, we can take such a sequence by considering a subsequence of the sequence in part (2) of Lemma 4.5. For any $i\in\{1,\ldots,k\}$ and any $v_{i}\in V(\Gamma_{i})$, we can take a longest sequence of separating hyperplanes $P_{1},\ldots,P_{M}$ from $A_{V_{0}}$ to $\gamma A_{V_{0}}$ such that $P_{1}$ is of $v_{i}$-type and $P_{M}$ is of $v_{1,n}$-type, where
$A_{V_{0}}\in P_{1}^{-},P_{1}^{+}\supsetneq P_{2}^{+}\supsetneq\cdots\supsetneq P_{M-1}^{+}\supsetneq P_{M}^{+}\ni\gamma A_{V_{0}}$
by taking decompositions by appropriate connected components
$C_{\Gamma,\Gamma_{0}}\!\setminus\!\!\setminus P_{1}=P_{1}^{-}\sqcup P_{1}^{+},\ldots,C_{\Gamma,\Gamma_{0}}\!\setminus\!\!\setminus P_{M}=P_{M}^{-}\sqcup P_{M}^{+}.$
Note that $P_{1},\ldots,P_{M}\in\{J_{j,d}\}_{j\in\{1,\ldots,k\},d\in\{1,\ldots,2(\mathfrak{m}+r)n\}}$ by Lemma 4.4. By noting $P_{M}\in\{J_{1,d}\}_{d\in\{1,\ldots,2(\mathfrak{m}+r)n\}}$ and part (1) of Lemma 4.5 for the case $i=1$, we have $P_{M}^{+}\supsetneq J^{\prime+}$.
Now assume that $Stab(J)\cap Stab(J^{\prime})\not\subset A_{V_{0}}$. Take $g\in Stab(J)\cap Stab(J^{\prime})$ with $g\not\in A_{V_{0}}$. Note that $g^{-1}\in Stab(J)\cap Stab(J^{\prime})$. Then, we have a hyperplane $H$ of $v_{i}$-type for some $i\in\{1,\ldots,k\}$ and some $v_{i}\in V(\Gamma_{i})$, separating $A_{V_{0}}$ and $gA_{V_{0}}$. We take a longest sequence of separating hyperplanes $P_{1},\ldots,P_{M}$ from $A_{V_{0}}$ to $\gamma A_{V_{0}}$ such that $P_{1}$ is of $v_{i}$-type and $P_{M}$ is of $v_{1,n}$-type. Then we have
$g\gamma A_{V_{0}},g^{-1}\gamma A_{V_{0}}\in P_{M}^{+}$
by $P_{M}^{+}\supsetneq J^{\prime+}$. We take a connected component $H^{-}$ of $C_{\Gamma,\Gamma_{0}}\!\setminus\!\!\setminus H$ such that $A_{V_{0}}t\in H^{-}$. Then the other connected component $H^{+}$ satisfies $gA_{V_{0}}\in H^{+}$. Because $H\cap J\neq\emptyset$, we have $H\cap J^{\prime}=\emptyset$ by (ii). Thus $H$ can not separate $\gamma A_{V_{0}}$ and $g\gamma A_{V_{0}}$. Hence we have either
  1. (iii-1-a)
    $A_{V_{0}},\gamma A_{V_{0}},g\gamma A_{V_{0}}\in H^{-}\text{ and }gA_{V_{0}}\in H^{+}$ or
  2. (iii-1-b)
    $A_{V_{0}}\in H^{-}\text{ and }gA_{V_{0}},\gamma A_{V_{0}},g\gamma A_{V_{0}}\in H^{+}$.
Assume that the case (iii-1-a) occurs. Then $H$ does not separate $A_{V_{0}}$ and $\gamma A_{V_{0}}$ and thus $H\notin\{J_{i,d}\}_{d\in\{1,\ldots,2(\mathfrak{m}+r)n\}}$ by Lemma 4.4. Hence $H$ and $P_{1}$ are different hyperplanes of the same $v_{i}$-type, and thus they can not intersect by Remark 2.13. Then we have either $H^{-}\supsetneq P_{1}^{+}$ or $H^{+}\supsetneq P_{1}^{+}$. However $H^{+}\supsetneq P_{1}^{+}$ can not occur because $\gamma A_{V_{0}}\notin H^{+}$ and $\gamma A_{V_{0}}\in P_{1}^{+}$. Therefore
$H^{-}\supsetneq P_{1}^{+}.$
Then we have
$gA_{V_{0}}\in H^{+},H^{-}\supsetneq P_{1}^{+}\supsetneq P_{2}^{+}\supsetneq\cdots\supsetneq P_{M-1}^{+}\supsetneq P_{M}^{+}\ni g\gamma A_{V_{0}}.$
Then, $Q_{1}=H,Q_{2}=P_{1},\ldots,Q_{M+1}=P_{M}$ is a sequence of separating hyperplanes from $gA_{V_{0}}$ to $g\gamma A_{V_{0}}$ such that $Q_{1}$ is of $v_{i}$-type and $Q_{M+1}$ is of $v_{1,n}$-type. Thus, $g^{-1}Q_{1},g^{-1}Q_{2},\ldots,g^{-1}Q_{M+1}$ is a sequence of separating hyperplanes from $A_{V_{0}}$ to $\gamma A_{V_{0}}$ such that $g^{-1}Q_{1}$ is of $v_{i}$-type and $g^{-1}Q_{M+1}$ is of $v_{1,n}$-type, which contradicts the fact that the sequence $P_{1},\ldots,P_{M}$ is longest.
Next, assume that the case (iii-1-b) occurs. Then $H$ does not separate $gA_{V_{0}}$ and $g\gamma A_{V_{0}}$, that is, $g^{-1}H$ does not separate $A_{V_{0}}$ and $\gamma A_{V_{0}}$ and thus $g^{-1}H\notin\{J_{i,d}\}_{d\in\{1,\ldots,2(\mathfrak{m}+r)n\}}$ by Lemma 4.4. Hence $g^{-1}H$ and $P_{1}$ are different hyperplanes of the same $v_{i}$-type, and thus they can not intersect by Remark 2.13. Then we have either $g^{-1}H^{+}\supsetneq P_{1}^{+}$ or $g^{-1}H^{-}\supsetneq P_{1}^{+}$. However $g^{-1}H^{-}\supsetneq P_{1}^{+}$ can not occur because $\gamma A_{V_{0}}\notin g^{-1}H^{-}$ and $\gamma A_{V_{0}}\in P_{1}^{+}$. Therefore
$g^{-1}H^{+}\supsetneq P_{1}^{+}.$
Then we have
$g^{-1}A_{V_{0}}\in g^{-1}H^{-},g^{-1}H^{+}\supsetneq P_{1}^{+}\supsetneq P_{2}^{+}\supsetneq\cdots\supsetneq P_{M-1}^{+}\supsetneq P_{M}^{+}\ni g^{-1}\gamma A_{V_{0}}.$
Then, $Q_{1}=H,Q_{2}=gP_{1},\ldots,Q_{M+1}=gP_{M}$ is a sequence of separating hyperplanes from $A_{V_{0}}$ to $\gamma A_{V_{0}}$ such that $Q_{1}$ is of $v_{i}$-type and $Q_{M+1}$ is of $v_{1,n}$-type, which contradicts the fact that the sequence $P_{1},\ldots,P_{M}$ is longest.
We now see that $Stab(J)\cap Stab(J^{\prime})\subset A_{V_{0}}$.
(iii-2) We show $Stab(J)\cap Stab(J^{\prime})\subset\gamma A_{V_{0}}\gamma^{-1}$. This can be shown by a similar way with (iii-1). Indeed we can lead a contradiction by assuming $Stab(J)\cap Stab(J^{\prime})\not\subset\gamma A_{V_{0}}\gamma^{-1}$.
[…]
Suppose that $g\in A_{V_{0}\setminus\{w^{\prime}\in V_{0}\mid w^{\prime}<w\}}=A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}$. Let $\ell$ be a path from $w_{0}=A_{V_{0}}$ to $w_{2(\mathfrak{m}+r)n}=\gamma A_{V_{0}}$ that diagonally penetrates each of the cubes $K_{1},\ldots,K_{2(\mathfrak{m}+r)n}$ in order. We note that the set of all hyperplanes intersecting the path $\ell$ is $\{J_{i,d}\}_{i\in\{1,\ldots,k\},d\in\{1,\ldots,2(\mathfrak{m}+r)n\}}$. Therefore, $\ell$ contains a vertex
(4.10) $\gamma_{\flat}(1)\cdots\gamma_{\flat}(\mathfrak{j}(w)-1)\lambda_{1}\cdots\lambda_{l(w_{\mathfrak{j}(w)})-1}\lambda_{l(w_{\mathfrak{j}(w)})}(\mathfrak{j}(w))A_{V_{0}},$
[…]
$\gamma_{\flat}(1)\cdots\gamma_{\flat}(\mathfrak{j}(w)-1)\lambda_{1}\cdots\lambda_{l(w_{\mathfrak{j}(w)})-1}$
$v_{1,l(w_{\mathfrak{j}(w)})}\cdots v_{i(w_{\mathfrak{j}(w)})-1,l(w_{\mathfrak{j}(w)})}v_{i(w_{\mathfrak{j}(w)})+1,l(w_{\mathfrak{j}(w)})}\cdots v_{k,l(w_{\mathfrak{j}(w)})}$
$w_{\mathfrak{j}(w)}(h-1)\cdots w_{\mathfrak{j}(w)}(1)w_{\mathfrak{j}(w)}(0)A_{V_{0}}$
$=\gamma_{\flat}(1)\cdots\gamma_{\flat}(\mathfrak{j}(w)-1)\lambda_{1}\cdots\lambda_{l(w_{\mathfrak{j}(w)})-1}$
$v_{1,l(w_{\mathfrak{j}(w)})}\cdots v_{i(w_{\mathfrak{j}(w)})-1,l(w_{\mathfrak{j}(w)})}v_{i(w_{\mathfrak{j}(w)})+1,l(w_{\mathfrak{j}(w)})}\cdots v_{k,l(w_{\mathfrak{j}(w)})}$
$w(h-1)\cdots w(1)w(0)A_{V_{0}}$
where we note $w(h-1)=w_{\mathfrak{j}(w)}$ and $w(0)=v_{i(w),l(w)}$. Hence
(4.11) $\begin{split}g\in&\gamma_{\flat}(1)\cdots\gamma_{\flat}(\mathfrak{j}(w)-1)\lambda_{1}\cdots\lambda_{l(w_{\mathfrak{j}(w)})-1}\lambda_{l(w_{\mathfrak{j}(w)})}(\mathfrak{j}(w))\\ &A_{V_{0}}(\gamma_{\flat}(1)\cdots\gamma_{\flat}(\mathfrak{j}(w)-1)\lambda_{1}\cdots\lambda_{l(w_{\mathfrak{j}(w)})-1}\lambda_{l(w_{\mathfrak{j}(w)})}(\mathfrak{j}(w)))^{-1}.\end{split}$
Since $g$ commutes with
$\gamma_{\flat}(1)\cdots\gamma_{\flat}(\mathfrak{j}(w)-1)\lambda_{1}\cdots\lambda_{l(w_{\mathfrak{j}(w)})-1}$
$v_{1,l(w_{\mathfrak{j}(w)})}\cdots v_{i(w_{\mathfrak{j}(w)})-1,l(w_{\mathfrak{j}(w)})}v_{i(w_{\mathfrak{j}(w)})+1,l(w_{\mathfrak{j}(w)})}\cdots v_{k,l(w_{\mathfrak{j}(w)})},$
$g\in(w(h-1)\cdots w(0))A_{V_{0}}(w(h-1)\cdots w(0))^{-1}$. By noting $w(0)<w(1)<\cdots<w(h-1)<w(h)=w$, we can show $g\in A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}>w\}}$. Indeed it follows from Lemma 2.11 that $g$ belongs to
$A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}$
$\cap(w(h-1)\cdots w(0))A_{V_{0}}(w(h-1)\cdots w(0))^{-1}$
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}\cap A_{V_{0}}$
$\cap(w(h-1)\cdots w(1)w(0))A_{V_{0}}(w(h-1)\cdots w(1)w(0))^{-1}$
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}\cap(w(h-1)\cdots w(1))A_{V_{0}}(w(h-1)\cdots w(1))^{-1}$
$\cap(w(h-1)\cdots w(1)w(0))A_{V_{0}}(w(h-1)\cdots w(1)w(0))^{-1}$
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}$
$\cap(w(h-1)\cdots w(1))(A_{V_{0}}\cap w(0)A_{V_{0}}w(0)^{-1})(w(h-1)\cdots w(1))^{-1}$
$\subset$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}$
$\cap(w(h-1)\cdots w(1))A_{V_{0}\setminus\{w(1)\}}(w(h-1)\cdots w(1))^{-1}$
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}\cap A_{V_{0}\setminus\{w(1)\}}$
$\cap(w(h-1)\cdots w(2)w(1))A_{V_{0}\setminus\{w(1)\}}(w(h-1)\cdots w(2)w(1))^{-1}$
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}\cap(w(h-1)\cdots w(2))A_{V_{0}\setminus\{w(1)\}}(w(h-1)\cdots w(2))^{-1}$
$\cap(w(h-1)\cdots w(2)w(1))A_{V_{0}\setminus\{w(1)\}}(w(h-1)\cdots w(2)w(1))^{-1}$
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}$
$\cap(w(h-1)\cdots w(2))(A_{V_{0}\setminus\{w(1)\}}\cap w(1)A_{V_{0}\setminus\{w(1)\}}w(1)^{-1})(w(h-1)\cdots w(2))^{-1}$
$\subset$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}$
$\cap(w(h-1)\cdots w(2))A_{V_{0}\setminus\{w(1),w(2)\}}(w(h-1)\cdots w(2))^{-1}$
$\cdots$
$\subset$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}$
$\cap w(h-1)A_{V_{0}\setminus\{w(1),\ldots,w(h-1)\}}w(h-1)^{-1}$
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}$
$\cap(A_{V_{0}\setminus\{w(1),\ldots,w(h-1)\}}\cap w(h-1)A_{V_{0}\setminus\{w(1),\ldots,w(h-1)\}}w(h-1)^{-1})$
$\subset$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}$
$\cap A_{V_{0}\setminus\{w(1),\ldots,w(h-1),w(h)\}}$
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}>w\}}.$
Since $<$ is the total order on $V_{0}=W(1)\sqcup\cdots\sqcup W(\mathfrak{d})$, which is a finite set, we repeat the same argument and have $g=1$. ∎
Proof of Corollaries 1.4 and 1.5.
Let $A_{\Gamma}$ be an irreduicble Artin group associated with $\Gamma$ that is not a clique. Then, Theorem 1.2 $(1)\Rightarrow(2)$ implies that $A_{\Gamma}$ is acylindrically hyperbolic. Therefore, the center of $A_{\Gamma}$ is finite by acylindrical hyperbolicity ([33, Corollary 7.3]). Proposition 3.8 then implies Corollary 1.4. We can show Corollary 1.5 by a similar way. ∎
Remark 4.6.
There are some versions of the so-called center conjecture. Note that the center of any irreducible Artin group of spherical type is known to be infinite cyclic [4], [17]. Then as a version, the center conjecture containing this known fact can be stated as follows: the center of any irreducible Artin group of spherical type and of infinite type is infinite cyclic and trivial, respectively. Since the center of any possibly reducible Artin group with each irreducible factor satisfying the center conjecture is trivial or torsion-free, the last part of Corollary 1.4 obviously implies the following: in the setting of Corollary 1.4, if each irreducible factor of $A_{\Gamma_{0}}$ satisfies the center conjecture, then $A_{\Gamma}$ satisfies the center conjecture. This claim is independently given in [27] by a different way.
Acknowledgments
The authors would like to thank Antoine Goldsborough, Nicolas Vaskou, Kasia Jankiewicz and MurphyKate Montee for useful conversations. Especially the authors thank Antoine Goldsborough and Nicolas Vaskou for their pointing us out that Theorem 1.2 has an application to random Artin groups.