### v3

# Acylindrical hyperbolicity and the centers of Artin groups that are not free of infinity

Abstract.
Charney and Morris-Wright showed acylindrical hyperbolicity of Artin groups of infinite type associated with graphs that are not joins, by studying clique-cube complexes and the actions on them. The authors developed their study and clarified when acylindrical hyperbolicity holds for Artin groups of infinite type associated with graphs that are not cones. In this paper, we introduce reduced clique-cube complexes. Using these complexes, we show acylindrical hyperbolicity of irreducible Artin groups associated with graphs that are not cliques, i.e., irreducible Artin groups that are not free of infinity. One application of our main result is that the centers of such Artin groups are finite, and that actually they are trivial in many cases. Another corollary is that irreducible Artin groups of infinite type and of type FC are acylindrically hyperbolic.
Key words and phrases:
Artin group, Acylindrical hyperbolicity, WPD contracting element, CAT(0)-cube complex, Center
2000 Mathematics Subject Classification:
20F36, 20F65, 20F67
The first author is supported by JSPS KAKENHI Grant Number 20K14311, and JST, ACT-X Grant Number JPMJAX200A, Japan.
The second author is supported by JSPS KAKENHI Grant Number 20K03590.
1. Introduction
Artin groups, also called Artin-Tits groups, have been widely studied since their introduction by Tits [36]. Let $\Gamma$ be a finite simple graph with the vertex set $V(\Gamma)$ and the edge set $E(\Gamma)$. Each edge $e$ has two endpoints, which we denote by $s_{e}$ and $t_{e}$. We suppose that any edge $e$ is labeled by an integer $\mu(e)\geq 2$. The Artin group $A_{\Gamma}$ associated with $\Gamma$ is defined by the following presentation:
(1.1) $A_{\Gamma}=\langle V(\Gamma)\mid\underbrace{s_{e}t_{e}s_{e}t_{e}\cdots}_{\text{length $\mu(e)$}}=\underbrace{t_{e}s_{e}t_{e}s_{e}\cdots}_{\text{length $\mu(e)$}}\quad\text{$\left(e\in E(\Gamma)\right)$ }\rangle.$
[…]
Figure 1. An example of a defining graph $\Gamma$ and the associated Coxeter graph $\Gamma^{t}$. Thick solid lines indicate edges labeled by $3$, and dotted lines indicate edges labeled by $2$. In $\Gamma^{t}$, thin solid lines denote edges labeled by $\infty$.
In terms of the properties of $W_{\Gamma}$ or $\Gamma^{t}$, we can define several important classes of Artin groups. The Artin group $A_{\Gamma}$ is said to be of spherical type if $W_{\Gamma}$ is finite. Otherwise, it is said to be of infinite type. The Artin group $A_{\Gamma}$ is said to be irreducible if $W_{\Gamma}$ is irreducible; that is, when $\Gamma^{t}$ is connected. If $\Gamma^{t}$ does not contain edges labeled by $\infty$, then we say $A_{\Gamma}$ is free of infinity. By definition, $A_{\Gamma}$ is free of infinity if and only if $\Gamma$ is a clique. In this paper, we focus on Artin groups that are not free of infinity; that is, Artin groups associated with graphs that are not cliques.
In general, compared with Artin groups, Coxeter groups are relatively well understood. For example, it is well known that an infinite Coxeter group $W_{\Gamma}$ is irreducible if and only if $W_{\Gamma}$ cannot be directly decomposed into two nontrivial subgroups ([32] and [36]). However, it is unclear whether $A_{\Gamma}$ is irreducible if and only if $A_{\Gamma}$ cannot be directly decomposed into two nontrivial subgroups. We remark that many basic questions about Artin groups remain open (refer to [12] and [20]).
To study Artin groups, it is often effective to investigate their hidden nonpositively curved or negatively curved properties. In this paper, we consider the following problem [23, Conjecture B].
Problem 1.1.
Are irreducible Artin groups of infinite type acylindrically hyperbolic?
Acylindrical hyperbolicity is an important property in geometric group theory. We can find various applications of acylindrical hyperbolicity in [16], [33] and [34] for example.
It is known that reducible Artin groups are not acylindrically hyperbolic, since acylindrically hyperbolic groups cannot be directly decomposed into two infinite subgroups ([33, Corollary 7.3]). We also note that Artin groups with infinite centers are not acylindrically hyperbolic, since acylindrically hyperbolic groups do not admit infinite centers ([33, Corollary 7.3]). Although irreducible Artin groups of spherical type have infinite cyclic center ([4] and [17]) and thus are not acylindrically hyperbolic, their central quotients are acylindrically hyperbolic when the rank is at least $2$ (see [2], [24], [5] for braid groups and [10] for the general case).
Many affirmative partial answers to Problem 1.1 are known. Indeed, the following irreducible Artin groups of infinite type are known to be acylindrically hyperbolic:
  • Right-angled Artin groups ([11], [30]).
  • Two-dimensional Artin groups such that the associated Coxeter groups are hyperbolic ([31]).
  • Artin groups of XXL-type ([23]).
  • Artin groups of type FC such that the defining graphs have diameter greater than $2$ ([15]). Here, an Artin group is of type FC if every standard parabolic subgroup associated with a clique in the defining graph is of spherical type.
  • Artin groups that are known to be CAT(0) groups according to the result of Brady and McCammond ([6] and [28]).
  • Euclidean Artin groups ([9]).
  • Two-dimensional Artin groups; that is, Artin groups for which every triangle with three edges $e_{1},e_{2},e_{3}$ in the defining graphs satisfy $\frac{1}{\mu(e_{1})}+\frac{1}{\mu(e_{2})}+\frac{1}{\mu(e_{3})}\leq 1$ ([42]).
  • Artin groups of which defining graphs that are not joins ([14]).
  • Artin groups of which defining graphs that are not cones ([29]).
  • Relatively extra-large-type Artin groups under mild conditions ([21]).
  • Even Artin groups of type FC, as well as Artin groups of type FC that visually split over standard parabolic subgroups of spherical type ([13]).
As a general strategy for Problem 1.1, we separately consider the case in which the defining graph is a clique and the case in which the defining graph is not a clique. Note that Artin groups in the latter case are automatically of infinite type. In this paper, we give an affirmative answer to Problem 1.1 in the latter case. Our main theorem is the following:
Theorem 1.2.
Let $A_{\Gamma}$ be an irreducible Artin group associated with $\Gamma$. Suppose that $\Gamma$ is not a clique. Then $A_{\Gamma}$ is acylindrically hyperbolic.
From Theorem 1.2, we find that many irreducible Artin groups of infinite type are acylindrically hyperbolic, e.g., the Artin group associated with the defining graph in Figure 1. We also apply Theorem 1.2 to Artin groups that are both of infinite type and of type FC.
Corollary 1.3.
Let $A_{\Gamma}$ be an irreducible Artin group of infinite type and of type FC. Then $A_{\Gamma}$ is acylindrically hyperbolic.
Note that, an Artin group of type FC is of spherical type if the defining graph is a clique.
Before we give other applications, we define the maximal clique factor of $\Gamma$. Let $\Gamma$ be a simple graph that is not a clique. If a factor of a join decomposition of $\Gamma$ is a clique, then it is called a clique factor. If $\Gamma$ has a clique factor, then it has a unique maximal one, which is called the maximal clique factor and usually denoted by $\Gamma_{0}$ in this paper. See Figures 1 and 2. For convenience, if $\Gamma$ has no nontrivial clique factor, we set $\Gamma_{0}=\emptyset$ and refer to it as the maximal clique factor, as well. See Section 2.5 for details related to join decompositions, and Figures 4 and 5 for other examples.
Figure 2. $\Gamma$ in Figure 1 has a join decomposition $\Gamma=\Gamma_{0}\ast\Gamma_{1}\ast\Gamma_{2}$, and $\Gamma_{0}$ is the maximal clique factor. Later we denote $\Gamma_{\ast}=\Gamma_{1}\ast\Gamma_{2}$. See Section 2.5.
We apply Theorem 1.2 to the so-called center conjecture, which states that the center of any irreducible Artin group of infinite type is trivial.
Corollary 1.4.
Let $A_{\Gamma}$ be an irreducible Artin group associated with $\Gamma$ that is not a clique. Then the center $Z(A_{\Gamma})$ is finite. Moreover, $Z(A_{\Gamma})$ is contained in $Z(A_{\Gamma_{0}})$. In particular, if $Z(A_{\Gamma_{0}})$ is trivial or torsion-free, then $Z(A_{\Gamma})$ is trivial. Here, $\Gamma_{0}$ is the maximal clique factor of $\Gamma$.
A result related to the last part of Corollary 1.4 can be found in [27]. See Remark 4.8.
We also obtain a corollary concerning direct products.
Corollary 1.5.
Let $A_{\Gamma}$ be an irreducible Artin group associated with $\Gamma$ that is not a clique. Then $A_{\Gamma}$ is virtually directly indecomposable; that is, it cannot be decomposed as a direct product of two infinite subgroups. Moreover, if $A_{\Gamma_{0}}$ contains no nontrivial finite subgroup that is a factor of a direct decomposition of $A_{\Gamma_{0}}$, then $A_{\Gamma}$ is directly indecomposable. Here, $\Gamma_{0}$ is the maximal clique factor of $\Gamma$.
Example 1.6.
Let $A_{\Gamma}$ be an irreducible Artin group associated with $\Gamma$ that is not a clique. Suppose that the maximal clique factor $\Gamma_{0}$ consists of three vertices or fewer. See Figures 1 and 2. Then $Z(A_{\Gamma_{0}})$ is known to be trivial or torsion-free, and thus $Z(A_{\Gamma})$ is trivial by Corollary 1.4. Also, $A_{\Gamma_{0}}$ is known to be torsion-free, and thus $A_{\Gamma}$ is directly indecomposable by Corollary 1.5.
Remark 1.7.
Goldsborough and Vaskou, who are the authors of [22], pointed to us out that the results in [22] and Theorem 1.2 imply acylindrical hyperbolicity of random Artin groups. The following is based on their remark. By referring to [22] for some definitions, we can state the resulting implication precisely: an Artin group picked at random relatively to any non-decreasing divergent function $f:\mathbb{N}\to\mathbb{N}$ is asymptotically almost surely acylindrically hyperbolic. Indeed the case of $f(N)\succ N^{\frac{3}{2}}$ and the case of $f(N)\prec N^{1-\alpha}$ for some $10<\alpha<$ 1 are covered in Corollaries 3.4 and 4.3 in [22], respectively. The missing case was unknown (see Question 1.6 in [22]), but now it is resolved as follows: we consider the case of $f(N)\prec N^{2}$, which contains the missing case. By Lemma 2.7, Theorem 2.9 (2) and Lemma 2.10 in [22], an Artin group picked at random relatively to $f$ asymptotically almost surely is irreducible and not free of infinity. Thus, it is acylindrically hyperbolic by Theorem 1.2.
In this paper, for the proof of Theorem 1.2, we develop arguments from [14] and [29]. Charney and Morris-Wright [14] established the acylindrical hyperbolicity of Artin groups of infinite type associated with graphs that are not joins, by studying clique-cube complexes, which are CAT(0) cube complexes, and the isometric actions on them. They constructed a WPD contracting element of such an Artin group with respect to the isometric action on the clique-cube complex. We note that the existence of a WPD contracting element is a sufficient condition for acylindrical hyperbolicity [1]. The authors [29] developed this approach and treated Artin groups associated with graphs that are not cones. To treat the case of Artin groups associated with graphs that are not cliques, we need additional development and more refined arguments. Indeed, we introduce the reduced clique-cube complex. It is a convex subcomplex of the clique-cube complex, and works more appropriately in our case than the clique-cube complex. Together with careful analysis of the local geometry of the cube complexes, we construct a candidate of a WPD contracting element with respect to the action on the reduced clique-cube complex. We also apply some algebraic results on standard parabolic subgroups of Artin groups to show that the candidate is actually a WPD contracting element.
The remainder of this paper is organized as follows. Section 2 contains some preliminaries. Section 2.1 is for preliminaries regarding acylindrically hyperbolic groups, WPD contracting elements, and CAT(0) cube complexes. Section 2.2 presents preliminaries on defining graphs of Artin groups. Section 2.3 presents facts on standard parabolic subgroups of Artin groups and gives a lemma on intersections of standard parabolic subgroups and their conjugates in Artin groups. Section 2.4 treats clique-cube complexes. Section 2.5 deals with the definition of joins of graphs and some related properties. In Section 2.6, we prepare some operations on matrices over a group, which are used when we construct a WPD contracting element in Section 4. In Section 3, we introduce reduced clique-cube complexes and study the local geometry. In Section 4, we state Theorem 4.1 which contains Theorem 1.2 and adds information, and prove it. Our main task is to construct a candidate WPD contracting element and show that it is really a WPD contracting element. We also prove Corollaries 1.4 and 1.5. In Section 5, we treat an example to aid understanding of Section 4.
2. preliminaries
2.1. Acylindrical hyperbolicity, WPD contracting elements, and CAT(0) cube complexes
We collect some definitions and properties related to acylindrical hyperbolicity, WPD contracting elements, and CAT(0) cube complexes for later use. See [18] and the references therein for details.
First, we recall the definition of acylindrical hyperbolicity (see [33]).
Definition 2.1.
A group $G$ is acylindrically hyperbolic if it admits an isometric action on a hyperbolic space $Y$ that is non-elementary (i.e., with an infinite limit set) and acylindrical (i.e., for every $D\geq 0$, there exist some $R,N\geq 0$ such that, for all $y_{1},y_{2}\in Y$, $d_{Y}(y_{1},y_{2})\geq R$ implies #$\#\{g\in G|\ d_{Y}(y_{1},g(y_{1})),d_{Y}(y_{2},g(y_{2}))\leq D\}\leq N$).
Next, we recall the definition of a WPD contracting element.
Definition 2.2.
Let a group $G$ act isometrically on a metric space $X$. For $\gamma\in G$, we say that:
  • $\gamma$ is weakly properly discontinuous (WPD) if, for every $D\geq 0$ and $x\in X$, there exists some $M\geq 1$ such that
    #$\#\{g\in G|\ d_{X}(x,g(x)),d_{X}(\gamma^{M}(x),g\gamma^{M}(x))\leq D\}<\infty;$
  • $\gamma$ is contracting if $\gamma$ is loxodromic, that is, there exists $x_{0}\in X$ such that the map $\mathbb{Z}\to X;n\mapsto\gamma^{n}(x_{0})$ is a quasi-isometry onto the image $\gamma^{\mathbb{Z}}x_{0}:=\{\gamma^{n}(x_{0})|n\in\mathbb{Z}\}$, and $\gamma^{\mathbb{Z}}x_{0}$ is contracting; that is, there exists $B\geq 0$ such that the diameter of the nearest-point projection of any ball that is disjoint from $\gamma^{\mathbb{Z}}x_{0}$ onto $\gamma^{\mathbb{Z}}x_{0}$ is bounded by $B$.
The following is a consequence of [1].
Theorem 2.3.
Let a group $G$ act isometrically on a geodesic metric space $X$. Suppose that $G$ is not virtually cyclic. If there exists a WPD contracting element $\gamma\in G$, then $G$ is acylindrically hyperbolic.
CAT(0) cube complexes are regarded as generalized trees in higher dimensions. The following is a precise definition (see [7, p.111]).
Definition 2.4.
A cube complex is a $CW$ complex constructed by gluing together cubes of arbitrary (finite) dimension by isometries along their faces. Furthermore, the cube complex is nonpositively curved if the link of any of its vertices is a simplicial flag complex (i.e., $n+1$ vertices span an $n$-simplex if and only if they are pairwise adjacent), and CAT(0) if it is nonpositively curved and simply connected.
Definition 2.5.
Let $X$ be a CAT(0) cube complex. We define an equivalence relation on the edges of $X$ as the transitive closure of the relation identifying two parallel edges of a square. For an equivalence class, a hyperplane is defined as the union of the midcubes transverse to the edges belonging to the equivalence class. Then, for any edge belonging to the equivalence class, the hyperplane is said to be dual to the edge.
For a hyperplane $J$, we denote the union of the cubes intersecting $J$ as $N(J)$; that is, the smallest subcomplex of $X$ containing $J$. We denote the union of the cubes not intersecting $J$ as $X\!\setminus\!\!\setminus J$; that is, the largest subcomplex of $X$ not intersecting $J$.
See [37] for the following.
Theorem 2.6.
Let $X$ be a CAT(0) cube complex and $J$ a hyperplane. Then, $X\!\setminus\!\!\setminus J$ has exactly two connected components.
The two connected components of $X\!\setminus\!\!\setminus J$ are often denoted by $J^{+}$ and $J^{-}$, respectively.
The following is adapted from [29, Definition 2.7].
Definition 2.7.
Let $X$ be a CAT(0) cube complex. For two vertices $x$ and $x^{\prime}$ in $X$, we call a sequence of hyperplanes $P_{1},\ldots,P_{M}$ a sequence of separating hyperplanes from $x$ to $x^{\prime}$ if the sequence satisfies
$x\in P_{1}^{-},P_{1}^{+}\supsetneq P_{2}^{+}\supsetneq\cdots\supsetneq P_{M-1}^{+}\supsetneq P_{M}^{+}\ni x^{\prime}$
for some connected components $P_{i}^{+}$ of $X\!\setminus\!\!\setminus P_{i}$ for all $i\in\{1,\ldots,M\}$.
For two hyperplanes $J$ and $J^{\prime}$ in $X$, we call a sequence of hyperplanes $P_{1},\ldots,P_{M}$ a sequence of separating hyperplanes from $J$ to $J^{\prime}$ if the sequence satisfies
$J^{+}\supsetneq P_{1}^{+}\supsetneq P_{2}^{+}\supsetneq\cdots\supsetneq P_{M-1}^{+}\supsetneq P_{M}^{+}\supsetneq J^{\prime+}$
for some connected components $J^{+}$ of $X\!\setminus\!\!\setminus J$, $J^{\prime+}$ of $X\!\setminus\!\!\setminus J^{\prime}$, and $P_{i}^{+}$ of $X\!\setminus\!\!\setminus P_{i}$ for all $i\in\{1,\ldots,M\}$.
Remark 2.8.
When $P_{1},\ldots,P_{M}$ is a sequence of separating hyperplanes from $J$ to $J^{\prime}$, for two vertices $x\in J^{-}\cup N(J)$ and $x^{\prime}\in J^{\prime+}\cup N(J^{\prime})$, $P_{1},\ldots,P_{M}$ is a sequence of separating hyperplanes from $x$ to $x^{\prime}$.
The following is based on part of [18, Theorem 3.3] and is used in the proof of Theorem 1.2.
Theorem 2.9.
Let a group $G$ act isometrically on a CAT(0) cube complex $X$. Then, $\gamma\in G$ is a WPD contracting element if there exist two hyperplanes $J$ and $J^{\prime}$ satisfying the following:
  1. (i)
    $J$ and $J^{\prime}$ are strongly separated; that is, no hyperplane intersects both $J$ and $J^{\prime}$;
  2. (ii)
    $\gamma$ skewers $J$ and $J^{\prime}$; that is, we have connected components $J^{+}$ of $X\!\setminus\!\!\setminus J$ and $J^{\prime+}$ of $X\!\setminus\!\!\setminus J^{\prime}$ such that $\gamma^{n}(J^{+})\subsetneq J^{\prime+}\subsetneq J^{+}$ for some $n\in\mathbb{N}$;
  3. (iii)
    $\mathrm{stab}(J)\cap\mathrm{stab}(J^{\prime})$ is finite, where $\mathrm{stab}(J)=\{g\in G\ |\ g(J)=J\}$ and $\mathrm{stab}(J^{\prime})=\{g\in G\ |\ g(J^{\prime})=J^{\prime}\}$.
2.2. Defining graphs of Artin groups
We present a precise description of the defining graph of an Artin group and introduce some related graphs.
Let $V$ be a finite set. Denote the diagonal set as $\mathrm{diag}(V\times V):=\{(v,w)\in V\times V|\ v=w\}$. We consider the involution on the off-diagonal set
$\iota:V\times V\setminus\mathrm{diag}(V\times V)\ni(v,w)\mapsto(w,v)\in V\times V\setminus\mathrm{diag}(V\times V).$
Any $e\in V\times V\setminus\mathrm{diag}(V\times V)$ is often presented as $(s_{e},t_{e})$. Then, for any $e\in V\times V\setminus\mathrm{diag}(V\times V)$, we have $s_{\iota(e)}=t_{e}$ and $t_{\iota(e)}=s_{e}$. We take a symmetric map
$\tilde{\mu}:V\times V\setminus\mathrm{diag}(V\times V)\to\mathbb{Z}_{\geq 2}\cup\{\infty\}.$
Here, ‘symmetric’ means that $\tilde{\mu}\circ\iota=\tilde{\mu}$ is satisfied. Then, we have
$V\times V\setminus\mathrm{diag}(V\times V)=\bigsqcup_{m\in\mathbb{Z}_{\geq 2}\cup\{\infty\}}\tilde{\mu}^{-1}(m).$
We now have a finite simple labeled graph $\Gamma$ with the vertex set $V(\Gamma)=V$, the edge set $E(\Gamma)=\bigsqcup_{m\in\mathbb{Z}_{\geq}2}\tilde{\mu}^{-1}(m)$, and the labeling $\mu:=\tilde{\mu}|_{E(\Gamma)}$. The Artin group $A_{\Gamma}$ associated with $\Gamma$ is defined by the presentation (1.1), and $\Gamma$ is called the defining graph of $A_{\Gamma}$.
For convenience, we define two other finite simple graphs, $\Gamma^{c}$ and $\Gamma^{t}$, as follows. $\Gamma^{c}$ is the finite simple graph with the vertex set $V(\Gamma^{c})=V$ and the edge set $E(\Gamma^{c})=\tilde{\mu}^{-1}(\infty)$. This is called the complement graph of $\Gamma$. The Coxeter graph $\Gamma^{t}$ is the finite simple graph with the vertex set $V(\Gamma^{t})=V$ and the edge set $E(\Gamma^{t})=\bigsqcup_{m\in\mathbb{Z}_{\geq 3}\cup\{\infty\}}\tilde{\mu}^{-1}(m)$.
2.3. Properties of Artin groups
We summarize some properties of Artin groups. Let $A_{\Gamma}$ be an Artin group associated with a defining graph $\Gamma$ with the vertex set $V(\Gamma)=V$ and the edge set $E(\Gamma)=E$, as in Section 2.2. Lemma 2.11 is essentially used in the proof of Theorem 1.2.
By the work of van der Lek (see [35], [41]), the following are well known and often used without mention in this paper.
Lemma 2.10.
  1. (1)
    For any subset $U\subset V$, the subgroup of $A_{\Gamma}$ generated by $U$, called a standard parabolic subgroup, is itself an Artin group associated with the full subgraph of $\Gamma$ spanned by $U$. (We denote this subgroup by $A_{U}$. When $U$ is empty, we define $A_{\emptyset}=\{1\}$.)
  2. (2)
    For any subsets $U,U^{\prime}\subset V$, $A_{U}\cap A_{U^{\prime}}=A_{U\cap U^{\prime}}$.
The following is a consequence of a nontrivial result by Godelle [19].
Lemma 2.11.
Take an element $v\in V$ and a subset $U\subset V$. Suppose that $v\not\in U$. Set $W:=\{w\in U\mid\mu(v,w)=2\}$. Then $A_{U}\cap vA_{U}v^{-1}=A_{W}$. In particular, for any $u\in U\setminus W$, $A_{U}\cap vA_{U}v^{-1}\subset A_{U\setminus\{u\}}$.
Proof.
Clearly we have $A_{W}\subset A_{U}\cap vA_{U}v^{-1}$. In terms of [19], $v$ is $(U,U)$-reduced in the Coxeter group $W_{\Gamma}$. Applying [19, Proposition 3.3] to our setting, we have some $U^{\prime}\subset U$ such that $vA_{U}v^{-1}\cap A_{U}=A_{U^{\prime}}$. Hence $W\subset U^{\prime}$. Assume that $W\subsetneq U^{\prime}$. Take $u^{\prime}\in U^{\prime}\setminus W$. Then $u^{\prime}\in A_{U^{\prime}}=vA_{U}v^{-1}\cap A_{U}$ and thus $v^{-1}u^{\prime}v\in A_{U}$. By $v^{-1}u^{\prime}v\in A_{\{v,u^{\prime}\}}$ and $\{v,u^{\prime}\}\cap U=\{u^{\prime}\}$, we have $v^{-1}u^{\prime}v\in A_{\{v,u^{\prime}\}}\cap A_{U}=A_{\{u^{\prime}\}}$. Therefore, we have $m\in\mathbb{Z}$ such that $v^{-1}u^{\prime}v=(u^{\prime})^{m}$. Since the surjection $A_{\{v,u^{\prime}\}}\to A_{\{u^{\prime}\}}$ defined by $v,u^{\prime}\mapsto u^{\prime}$ gives $v^{-1}u^{\prime}v\mapsto u^{\prime}$ and $(u^{\prime})^{m}\mapsto(u^{\prime})^{m}$, we see $u^{\prime}={(u^{\prime})}^{m}$ and $m=1$. Hence we have $v^{-1}u^{\prime}v=u^{\prime}$ in $A_{\Gamma}$. This means $\mu(v,u^{\prime})=2$, contradicting $u^{\prime}\notin W$. Hence $U^{\prime}=W$. Therefore, $vA_{U}v^{-1}\cap A_{U}=A_{W}$. ∎
2.4. Clique-cube complexes and actions on them
We consider clique-cube complexes and the actions on them following [14]. Let $A_{\Gamma}$ be an Artin group associated with a defining graph $\Gamma$ with the vertex set $V(\Gamma)=V$ and the edge set $E(\Gamma)=E$, as in Section 2.2. We say that $U$ spans a clique in $\Gamma$ if any two elements of $U$ are joined by an edge in $\Gamma$.
Definition 2.12 ([14, Definition 2.1]).
Consider the set
$\Delta_{\Gamma}=\{U\subset V|\ U\text{ spans a clique in }\Gamma,\text{ or }U=\emptyset\}.$
The clique-cube complex $C_{\Gamma}$ is the cube complex whose vertices (i.e., 00-dimensional cubes) are cosets $gA_{U}$ ($g\in A_{\Gamma},U\in\Delta_{\Gamma}$), where two vertices $gA_{U}$ and $hA_{U^{\prime}}$ are joined by an edge (i.e., a $1$-dimensional cube) in $C_{\Gamma}$ if $gA_{U}\subset hA_{U^{\prime}}$ and $U$ and $U^{\prime}$ differ by a single generator. Note that, in this case, we can always replace $h$ by $g$ so that $hA_{U^{\prime}}=gA_{U^{\prime}}$. More generally, two vertices $gA_{U}$ and $gA_{U^{\prime}}$ with $gA_{U}\subset gA_{U^{\prime}}$ span a #$\#(U^{\prime}\setminus U)$-dimensional cube $[gA_{U},gA_{U^{\prime}}]$ in $C_{\Gamma}$.
See Figure 3.
Figure 3. For the above graph $\Gamma$, the below cube complex is a part of $C_{\Gamma}$ that is a fundamental domain of $C_{\Gamma}$ with respect to the action of $A_{\Gamma}$. We define a subcomplex $C_{\Gamma,\Gamma_{0}}$ of $C_{\Gamma}$ in Section 3. The subcomplex consisting of colored vertices and thick edges is a part of $C_{\Gamma,\Gamma_{0}}$, which is a fundamental domain of $C_{\Gamma,\Gamma_{0}}$ with respect to the action of $A_{\Gamma}$.
We endow a metric on $C_{\Gamma}$ in the standard way for a cube complex. That is, we identify all cubes in $C_{\Gamma}$ with Euclidean unit cubes and give $C_{\Gamma}$ the path metric. Note that $C_{\Gamma}$ is connected.
Theorem 2.13 ([14, Theorem 2.2]).
The clique-cube complex $C_{\Gamma}$ is a finite dimensional CAT(0) cube complex. In particular, $C_{\Gamma}$ is a complete CAT(0) space.
Note that completeness follows from finite dimensionality of $C_{\Gamma}$ ([7, Theorem 7.19]).
The group $A_{\Gamma}$ acts on the clique-cube complex $C_{\Gamma}$ by left multiplication, $h\cdot gA_{U}=(hg)A_{U}$. This action preserves the cubical structure and is isometric. The action is also cocompact with a fundamental domain $\bigcup_{U\in\Delta_{\Gamma}}[A_{\emptyset},A_{U}]$, where $[A_{\emptyset},A_{U}]$ is a #$\#U$-dimensional cube spanned by the two vertices $A_{\emptyset}$ and $A_{U}$ in $C_{\Gamma}$. See Figure 3. However, the action is not proper. In fact, the stabilizer of a vertex $gA_{U}$ is the conjugate subgroup $gA_{U}g^{-1}$, so all vertices other than those in the orbit of $A_{\emptyset}$ have infinite stabilizers. We also note that $C_{\Gamma}$ is not a proper metric space because it contains vertices of infinite valence. Additionally, $C_{\Gamma}$ has infinite diameter if and only if $\Gamma$ itself is not a clique.
Remark 2.14 ([14, Section 2]).
Each edge in $C_{\Gamma}$ can be labeled with a generator in $V$. Indeed, the edge between $gA_{U}$ and $gA_{U\sqcup\{v\}}$ is labeled by $v$. Any two parallel edges in a cube have the same label, so we can also label the hyperplane dual to such an edge by $v$ and say that such a hyperplane is of $v$-type. Every hyperplane of $v$-type is a translation of the hyperplane dual to the edge between $A_{\emptyset}$ and $A_{\{v\}}$. If a hyperplane of $v$-type crosses another hyperplane of $v^{\prime}$-type, then $(v,v^{\prime})\in E$. In particular, two different hyperplanes of the same type do not cross each other.
2.5. Joins of graphs
Definition 2.15.
Let $\Lambda\neq\emptyset$ be an index set. The join $\ast_{\alpha\in\Lambda}\Gamma_{\alpha}$ of simple graphs $\Gamma_{\alpha}$ ($\alpha\in\Lambda$) is defined as a simple graph with the vertex set
$V(\ast_{\alpha\in\Lambda}\Gamma_{\alpha}):=\bigsqcup_{\alpha\in\Lambda}V(\Gamma_{\alpha})$
and the edge set
$E(\ast_{\alpha\in\Lambda}\Gamma_{\alpha}):=\bigsqcup_{\alpha\in\Lambda}E(\Gamma_{\alpha})\sqcup\bigsqcup_{\alpha,\beta\in\Lambda,\alpha\neq\beta}\{(v_{\alpha},v_{\beta})\mid v_{\alpha}\in V(\Gamma_{\alpha}),v_{\beta}\in V(\Gamma_{\beta})\}.$
A simple graph $\Gamma$ is said to be decomposable as a join if there exists an index set $\Lambda$ with #$\#\Lambda\geq 2$ and subgraphs $\Gamma_{\alpha}$ ($\alpha\in\Lambda$) of $\Gamma$ such that $\Gamma=\ast_{\alpha\in\Lambda}\Gamma_{\alpha}$. This is called a join decomposition of $\Gamma$ into factors $\Gamma_{\alpha}$ ($\alpha\in\Lambda$). $\Gamma$ is said to be indecomposable as a join if it is not decomposable as a join.
Remark 2.16.
A simple graph $\Gamma$ is indecomposable as a join if and only if its complement graph $\Gamma^{c}$ is connected.
The following is a well-known fact (See [29, Lemma 3.4]).
Lemma 2.17.
Let $\Gamma$ be a simple graph. Suppose that $\Gamma$ is decomposable as a join. Then, $\Gamma$ has a unique join decomposition into indecomposable factors
$\Gamma=\ast_{\alpha\in\Lambda}\Gamma_{\alpha},$
and the decomposition of $\Gamma^{c}$ into connected components is given by
$\Gamma^{c}=\bigsqcup_{\alpha\in\Lambda}(\Gamma_{\alpha})^{c}.$
For convenience, even if $\Gamma$ is indecomposable as a join, we take $\Lambda$ consisting of a single index $\alpha$, define $\Gamma_{\alpha}=\Gamma$ and regard $\Gamma=\ast_{\alpha\in\Lambda}\Gamma_{\alpha}$ as the unique join decomposition of $\Gamma$ (by a single indecomposable factor).
Figure 4. The defining graphs $\Gamma$ and the associated Coxeter graphs $\Gamma^{t}$. Each $\Gamma$ has a join decomposition $\Gamma=\Gamma_{0}\ast\Gamma_{1}$, where $\Gamma_{0}$ is the maximal clique factor and $\Gamma_{1}$ is an indecomposable factor with two vertices.
Figure 5. The defining graphs $\Gamma$ and the associated Coxeter graphs $\Gamma^{t}$. Each $\Gamma$ has a join decomposition $\Gamma=\Gamma_{0}\ast\Gamma_{1}\ast\Gamma_{2}$, where $\Gamma_{0}$ is the maximal clique factor and each of $\Gamma_{1}$ and $\Gamma_{2}$ is an indecomposable factor with two vertices.
Definition 2.18.
Let $\Gamma$ be a simple graph. For the unique join decomposition into indecomposable factors $\Gamma=\ast_{\alpha\in\Lambda}\Gamma_{\alpha}$, we set $\Lambda_{0}:=\{\alpha\in\Lambda\mid\ \#\Gamma_{\alpha}=1\}$ and $\Gamma_{0}:=\ast_{\alpha\in\Lambda_{0}}\Gamma_{\alpha}$. We call $\Gamma_{0}$ the maximal clique factor of $\Gamma$. Also we set $\Lambda_{\ast}:=\Lambda\setminus\Lambda_{0}$ and $\Gamma_{\ast}:=\ast_{\alpha\in\Lambda_{\ast}}\Gamma_{\alpha}$. We call $\Gamma$ a cone if $\Lambda_{0}\neq\emptyset$.
See Figures 1, 2, 4 and 5.
We have $\Gamma=\Gamma_{0}\ast\Gamma_{\ast}$. Clearly, $\Gamma$ is a clique if and only if $\Lambda=\Lambda_{0}$.
2.6. Notation on matrices
For natural numbers $k$, $m$, and a set $X$, let $\mathcal{M}(k,m;X)$ be the set of all $k\times m$ matrices whose entries are in $X$. For $i\in\{1,\ldots,k\}$ and $l\in\{1,\ldots,m\}$, let $M[i,l]$ be the $(i,l)$-component of $M$. Let $M[i,\ ]$ be the $i$-th row and $M[\ ,l]$ be the $l$-th column of the matrix $M$. For $l_{1},l_{2}\in\{1,\ldots,m\}$ with $l_{1}\leq l_{2}$, we define $M[\ ,l_{1}:l_{2}]\in\mathcal{M}(k,l_{2}-l_{1}+1;X)$ by
(2.1) $\left(M[\ ,l_{1}:l_{2}]\right)[\ ,l]=M[\ ,l_{1}-1+l]\quad(1\leq l\leq l_{2}-l_{1}+1).$
For two matrices $M_{j}\in\mathcal{M}(k,m_{j};X)$ $(j=1,2)$ with the same number of rows, we define $cbind(M_{1},M_{2})\in\mathcal{M}(k,m_{1}+m_{2};X)$ as a matrix constructed from $M_{1},M_{2}$ by binding the columns. That is,
(2.2) $cbind(M_{1},M_{2})[\ ,l]=\begin{cases}M_{1}[\ ,l]\quad(1\leq l\leq m_{1})\\ M_{2}[\ ,l-m_{1}]\quad(m_{1}+1\leq l\leq m_{1}+m_{2}).\end{cases}$
For matrices $M_{j}\in\mathcal{M}(k,m_{j};X)$ $(1\leq j\leq r)$ with the same number of rows, we define $cbind(M_{1},\ldots,M_{r})\in\mathcal{M}(k,m_{1}+\cdots+m_{r};X)$ inductively by
(2.3) $cbind(M_{1},\ldots,M_{r})=cbind\left(cbind(M_{1},\ldots,M_{r-1}),M_{r}\right).$
Let $A$ be a group and let $M\in\mathcal{M}(k,m;A)$. For any subset $B\subset A$, which is typically a subgroup or a generating set, we naturally regard $\mathcal{M}(k,m;B)\subset\mathcal{M}(k,m;A)$. For two matrices $M_{1},M_{2}\in\mathcal{M}(k,m;A)$ of the same size, $M_{1}\odot M_{2}\in\mathcal{M}(k,m;A)$ denotes the Hadamard product of $M_{1}$ and $M_{2}$. That is,
(2.4) $(M_{1}\odot M_{2})[i,l]=M_{1}[i,l]\cdot M_{2}[i,l]\in A.$
We define a map $prod:\mathcal{M}(k,m;A)\to A$ inductively as follows. When $m=1$, let
(2.5) $prod(M)=M[1,1]\cdot M[2,1]\cdot\cdots\cdot M[k,1]\in A.$
If $prod$ is defined on $\mathcal{M}(k,m-1;A)$, we define $prod:\mathcal{M}(k,m;A)\to A$ by
(2.6) $prod(M)=prod(M[\ ,1:m-1])\cdot prod(M[\ ,m])\in A.$
3. Reduced clique-cube complexes and actions on them
In this section, we introduce and study reduced clique-cube complexes. Let $A_{\Gamma}$ be an Artin group associated with a defining graph $\Gamma$ with the vertex set $V(\Gamma)=V$ and the edge set $E(\Gamma)=E$, as in Section 2.2. Let $\Gamma_{0}$ be the maximal clique factor of $\Gamma$. We consider the join decomposition $\Gamma=\Gamma_{0}\ast\Gamma_{\ast}$. We set $V_{0}=V(\Gamma_{0})$.
Definition 3.1.
Let
$\Delta_{\Gamma,\Gamma_{0}}=\{U\subset V\mid U\in\Delta_{\Gamma},\ U\supset V_{0}\}.$
Let $C_{\Gamma,\Gamma_{0}}$ be the subcomplex of $C_{\Gamma}$ spanned by vertices $gA_{U}$ with $g\in A_{\Gamma}$ and $U\in\Delta_{\Gamma,\Gamma_{0}}$. We call $C_{\Gamma,\Gamma_{0}}$ the reduced clique-cube complex.
See Figure 3.
When $\Gamma$ is not a cone; that is, $\Gamma=\Gamma_{\ast}$, we have $C_{\Gamma,\Gamma_{0}}=C_{\Gamma}$. When $\Gamma$ is a clique; that is, $\Gamma=\Gamma_{0}$, we have $C_{\Gamma,\Gamma_{0}}=C_{\Gamma,\Gamma}=\{A_{\Gamma}\}$. In this paper, we are mainly interested in the case where $\Gamma$ is not a clique.
We note that the action of $A_{\Gamma}$ on $C_{\Gamma}$ preserves $C_{\Gamma,\Gamma_{0}}$. We consider the induced action of $A_{\Gamma}$ on $C_{\Gamma,\Gamma_{0}}$. Then $\bigcup_{U\in\Delta_{\Gamma,\Gamma_{0}}}[A_{V_{0}},A_{U}]$ is a fundamental domain of $C_{\Gamma,\Gamma_{0}}$ with respect to the action of $A_{\Gamma}$.
Lemma 3.2.
$C_{\Gamma,\Gamma_{0}}$ is a connected subcomplex of $C_{\Gamma}$.
Proof.
If $\Gamma$ is a clique, then $C_{\Gamma,\Gamma_{0}}=\{A_{\Gamma}\}$ is clearly connected. Suppose that $\Gamma$ is not a clique. Let $U\in\Delta_{\Gamma,\Gamma_{0}}$. Let $g\in A_{\Gamma}$. First, we show that there exists a path in $C_{\Gamma,\Gamma_{0}}$ from $A_{V_{0}}$ to $gA_{V_{0}}$. Suppose that $g\notin A_{V_{0}}$; otherwise, we would have $A_{V_{0}}=gA_{V_{0}}$. Let $g=v_{1}^{\varepsilon_{1}}\cdots v_{n}^{\varepsilon_{n}}$, where $n\in\mathbb{N}$, $v_{i}\in V$, $\varepsilon_{i}\in\{1,-1\}$ for every $1\leq i\leq n$. By induction on $n$, we suppose that there exists a path in $C_{\Gamma,\Gamma_{0}}$ from $A_{V_{0}}$ to $v_{1}^{\varepsilon_{1}}\cdots v_{n-1}^{\varepsilon_{n-1}}A_{V_{0}}$. If $v_{n}\in V_{0}$, then
$v_{1}^{\varepsilon_{1}}\cdots v_{n-1}^{\varepsilon_{n-1}}A_{V_{0}}=v_{1}^{\varepsilon_{1}}\cdots v_{n}^{\varepsilon_{n}}A_{V_{0}},$
and the induction proceeds. If $v_{n}\notin V_{0}$, since $V_{0}\sqcup\{v_{n}\}\in\Delta_{\Gamma,\Gamma_{0}}$, there exists a path
$\left(v_{1}^{\varepsilon_{1}}\cdots v_{n-1}^{\varepsilon_{n-1}}A_{V_{0}},\ v_{1}^{\varepsilon_{1}}\cdots v_{n-1}^{\varepsilon_{n-1}}A_{V_{0}\sqcup\{v_{n}\}}=gA_{V_{0}\sqcup\{v_{n}\}},\ gA_{V_{0}}\right)$
in $C_{\Gamma,\Gamma_{0}}$ from $v_{1}^{\varepsilon_{1}}\cdots v_{n-1}^{\varepsilon_{n-1}}A_{V_{0}}$ to $gA_{V_{0}}$. By concatenating these two paths, we obtain a path from $A_{V_{0}}$ to $gA_{V_{0}}$ in $C_{\Gamma,\Gamma_{0}}$. Next, we show the existence of a path in $C_{\Gamma,\Gamma_{0}}$ from $gA_{V_{0}}$ to $gA_{U}$. Let $U=V_{0}\sqcup\{u_{1},\ldots,u_{m}\}$ where $u_{i}\neq u_{j}$ for $1\leq i\neq j\leq m$. Then there exists a path
$\left(gA_{V_{0}},gA_{V_{0}\sqcup\{u_{1}\}},gA_{V_{0}\sqcup\{u_{1},u_{2}\}},\ldots,gA_{U}\right)$
in $C_{\Gamma,\Gamma_{0}}$ from $gA_{V_{0}}$ to $gA_{U}$. Therefore, it follows that $C_{\Gamma,\Gamma_{0}}$ is a connected subcomplex of $C_{\Gamma}$. ∎
We endow $C_{\Gamma,\Gamma_{0}}$ with a metric in the standard way for a cube complex. That is, we identify all cubes in $C_{\Gamma,\Gamma_{0}}$ with Euclidean unit cubes and give $C_{\Gamma,\Gamma_{0}}$ the path metric. Note that $C_{\Gamma,\Gamma_{0}}$ is connected.
Definition 3.3.
Let $X$ be a geodesic space and $W$ be a subset of $X$. $W$ is convex in $X$ if for every $w_{1},w_{2}\in W$, every geodesic from $w_{1}$ to $w_{2}$ is contained in $W$.
Definition 3.4.
Let $L$ be a simplicial complex and $L^{\prime}$ be a subcomplex of $L$. $L^{\prime}$ is full if any simplex of $L$ whose vertices are in $L^{\prime}$ is contained in $L^{\prime}$.
Theorem 3.5 (see, e.g., [38]).
Let $X$ be a finite dimensional CAT(0) cube complex and $W$ be a connected subcomplex. Then $W$ is convex in $X$ if and only if for every vertex $v$ of $W$, the link $L^{\prime}$ of $v$ in $W$ is a full subcomplex of the link $L$ of $v$ in $X$.
Lemma 3.6.
$C_{\Gamma,\Gamma_{0}}$ is convex in $C_{\Gamma}$. In particular, $C_{\Gamma,\Gamma_{0}}$ is a finite dimensional CAT(0) cube complex and a complete CAT(0) space.
Proof.
We show that $C_{\Gamma,\Gamma_{0}}$ is convex in $C_{\Gamma}$. Let $gA_{U}$ be a vertex of $C_{\Gamma,\Gamma_{0}}$. Let $L$ be the link of $gA_{U}$ in $C_{\Gamma}$ and let $L^{\prime}$ be the link of $gA_{U}$ in $C_{\Gamma,\Gamma_{0}}$. According to Theorem 3.5 and Lemma 3.2, it is sufficient to show that $L^{\prime}$ is a full subcomplex of $L$. Let $\Delta$ be a $d$-dimensional simplex of $L$ whose vertices are in $L^{\prime}$. Let $C_{\Delta}$ be the $(d+1)$-dimensional cube of $C_{\Gamma}$ corresponding to $\Delta$. Let $gA_{U_{1}},\ldots,gA_{U_{d+1}}$ be vertices of $C_{\Delta}$ that are connected to $gA_{U}$ by a $1$-dimensional cube of $C_{\Gamma}$. Note that $C_{\Delta}=[gA_{\cap_{i=1}^{d+1}U_{i}},gA_{\cup_{i=1}^{d+1}U_{i}}]$. Since $gA_{U_{i}}\in C_{\Gamma,\Gamma_{0}}$, we have $U_{i}\supset V_{0}$ for every $i$. Therefore, $\cap_{i=1}^{d+1}U_{i}\supset V_{0}$. Since $C_{\Delta}$ is contained in $C_{\Gamma,\Gamma_{0}}$, $\Delta$ is contained in $L^{\prime}$. It follows that $L^{\prime}$ is a full subcomplex of $L$.
The other assertions follow from the convexity of $C_{\Gamma,\Gamma_{0}}$ in $C_{\Gamma}$ and Theorem 2.13. ∎
The following is a generalization of [14, Lemma 3.2].
Lemma 3.7.
The action of $A_{\Gamma}$ on $C_{\Gamma,\Gamma_{0}}$ is a minimal action. That is, for every $x\in C_{\Gamma,\Gamma_{0}}$, the convex hull of the orbit of $x$, $\mathrm{Hull}(orb(x))$, is $C_{\Gamma,\Gamma_{0}}$.
Proof.
When $\Gamma$ is a clique, $C_{\Gamma,\Gamma_{0}}=\{A_{V_{0}}\}$. In this case, the lemma is trivial. We now treat the case where $\Gamma$ is not a clique. We fix a join decomposition
$\Gamma=\Gamma_{0}\ast\Gamma_{1}\ast\cdots\ast\Gamma_{k},$
where $\Gamma_{0}$ is the maximal clique factor of $\Gamma$ and $\Gamma_{i}$ are indecomposable factors with at least two vertices for $i\geq 1$. We note that $\Gamma_{1}\ast\cdots\ast\Gamma_{k}\neq\emptyset$, since $\Gamma$ is not a clique.
Let $x\in C_{\Gamma,\Gamma_{0}}$. Let $C$ be a cube of $C_{\Gamma,\Gamma_{0}}$ with $x\in C$. Without loss of generality, we set $C=[A_{U},A_{T}]$ where $U\subset T$, $U,T\in\Delta_{\Gamma,\Gamma_{0}}$. By induction on #$\#(T\setminus U)$, we show that $\mathrm{Hull}(orb(x))=C_{\Gamma,\Gamma_{0}}$.
When #$\#(T\setminus U)=0$, then $x=A_{U}$. It is sufficient to show that $\mathrm{Hull}(orb(A_{U}))=\mathrm{Hull}(orb(A_{V_{0}}))$. Indeed, if $\mathrm{Hull}(orb(A_{U}))=\mathrm{Hull}(orb(A_{V_{0}}))$ for every $U\in\Delta_{\Gamma,\Gamma_{0}}$, then $\mathrm{Hull}(orb(A_{V_{0}}))$ contains every vertex of $C_{\Gamma,\Gamma_{0}}$ and thus equals $C_{\Gamma,\Gamma_{0}}$.
First, we show that $\mathrm{Hull}(orb(A_{U}))\supset\mathrm{Hull}(orb(A_{V_{0}}))$. If $V_{0}\neq U$, we take $u\in U\setminus V_{0}$. There exists $i\in\{1,\ldots,k\}$ such that $u\in V(\Gamma_{i})$. Since $\Gamma_{i}$ is indecomposable, we have a vertex $s\neq u$ of $\Gamma_{i}$ such that $s$ and $u$ are not connected by an edge in $\Gamma_{i}$. Note that $s\notin U$. Let
$U^{\prime}=\{v\in U\mid\text{ $v$ and $s$ are connected by an edge in $\Gamma$}\}.$
We note that ${U^{\prime}}$ does not contain $u$. Hence we have $V_{0}\subset{U^{\prime}}\subsetneq U$. We also note that ${U^{\prime}}\sqcup\{s\}$ spans a clique but $U\sqcup\{s\}$ does not. Then, the sequence of vertices
$A_{U},\ A_{U^{\prime}},\ A_{{U^{\prime}}\sqcup\{s\}},\ sA_{U^{\prime}},\ sA_{U}$
forms a local geodesic, which is a geodesic, since $C_{\Gamma,\Gamma_{0}}$ is CAT(0) (Lemma 3.6). This geodesic is contained in $\mathrm{Hull}(orb(A_{U}))$, since the two endpoints are in $\mathrm{Hull}(orb(A_{U}))$. Therefore, $A_{U^{\prime}}\in\mathrm{Hull}(orb(A_{U}))$, and thus $\mathrm{Hull}(orb(A_{U^{\prime}}))\subset\mathrm{Hull}(orb(A_{U}))$. If $U^{\prime}=V_{0}$, then we have nothing further to show. If ${U^{\prime}}\supsetneq V_{0}$, we repeat the same argument until we obtain $\mathrm{Hull}(orb(A_{U}))\supset\mathrm{Hull}(orb(A_{V_{0}}))$.
Next, we show that $\mathrm{Hull}(orb(A_{U}))\subset\mathrm{Hull}(orb(A_{V_{0}}))$. If $V_{0}\neq U$, we take $u\in U\setminus V_{0}$. Then
$A_{U\setminus\{u\}},\ A_{U},\ uA_{U\setminus\{u\}}$
is a geodesic. It follows that $\mathrm{Hull}(orb(A_{U}))\subset\mathrm{Hull}(orb(A_{U\setminus\{u\}}))$. If $U\setminus\{u\}=V_{0}$, then the proof is finished. If $U\setminus\{u\}\supsetneq V_{0}$, we repeat the same argument until we obtain $\mathrm{Hull}(orb(A_{U}))\subset\mathrm{Hull}(orb(A_{V_{0}}))$.
When #$\#(T\setminus U)>0$, there exists $s\in T\setminus U$. Then $C\cap sC$ contains a face spanned by $A_{U\sqcup\{s\}}$ and $A_{T}$. The geodesic from $x$ to $sx$ passes through this face. Let $y$ be the intersection point. Then $y\in\mathrm{Hull}(orb(x))$, and thus $\mathrm{Hull}(orb(y))\subset\mathrm{Hull}(orb(x))$. By induction, $\mathrm{Hull}(orb(y))=C_{\Gamma,\Gamma_{0}}$. ∎
The following is a generalization of [29, Proposition 4.8].
Proposition 3.8.
Suppose that $\Gamma$ is not a clique. Then, any finite normal subgroup of $A_{\Gamma}$ is contained in $A_{V_{0}}$. In particular, any finite subgroup of the center of $A_{\Gamma}$ is contained in $A_{V_{0}}$. Also, if $A_{\Gamma}$ is decomposed as a direct product $A_{\Gamma}=A_{1}\times A_{2}$ and $A_{1}$ is finite, then $A_{1}$ is contained in $A_{V_{0}}$ and we have $A_{V_{0}}=A_{1}\times(A_{V_{0}}\cap A_{2})$.
Proof.
Let $N$ be a finite normal subgroup of $A_{\Gamma}$. Set
$Fix(N):=\{x\in C_{\Gamma,\Gamma_{0}}\mid nx=x\text{ for any }n\in N\}.$
Because $N$ is finite and $C_{\Gamma,\Gamma_{0}}$ is a complete CAT(0) space, we have $Fix(N)\neq\emptyset$. Take any $x\in Fix(N)$. Then, $orb(x)=A_{\Gamma}x\subset Fix(N)$. Indeed, the normality of $N$ implies that, for any $g\in A_{\Gamma}$ and $n\in N$, there exists $n^{\prime}\in N$ such that $ng=gn^{\prime}$. Therefore, we have $ngx=gn^{\prime}x=gx$. Because $C_{\Gamma,\Gamma_{0}}$ is CAT(0), $Fix(N)$ is convex. Hence, we have $Hull(orb(x))\subset Fix(N)$. By Lemma 3.7, we have $Hull(orb(x))=C_{\Gamma,\Gamma_{0}}$. Hence, $Fix(N)=C_{\Gamma,\Gamma_{0}}$. In particular, $A_{V_{0}}\in Fix(N)$. Hence, $N$ must be contained in $A_{V_{0}}$. ∎
Finally, we state a lemma related to the local geometry of reduced clique-cube complexes. The following is a nontrivial variant in the setting of $C_{\Gamma,\Gamma_{0}}$ of [29, Lemma 5.2] concerning $C_{\Gamma}$.
Lemma 3.9.
Let $s,t\in V(\Gamma_{\ast})$ belong to mutually different indecomposable factors of $\Gamma_{\ast}$. Suppose that we have $d\in\mathbb{N}$ and mutually different $w_{0},w_{1},\ldots,w_{d}\in V$ such that $w_{0}=s$, $w_{d}=t$ and $(w_{i-1},w_{i})\in E$ with label greater than $2$ for any $i\in\{1,\ldots,d\}$. Moreover, suppose that $w_{i}\in V_{0}$ if $0<i<d$. Set $\tau:=w_{0}w_{1}\cdots w_{d}\in A_{\Gamma}$. Let $U\in\Delta_{\Gamma,\Gamma_{0}}$ with $s,t\in U$. Then, there exists no square in $C_{\Gamma,\Gamma_{0}}$ containing both edges $[A_{U},A_{U\setminus\{s\}}]$ and $[A_{U},\tau A_{U\setminus\{t\}}]$. That is, there exists no $g\in A_{U\setminus\{s\}}$ such that $gA_{U\setminus\{t\}}=\tau A_{U\setminus\{t\}}$.
See Figure 6.
Figure 6. Squares around $A_{U}$ where $s=w_{0}$, $t=w_{d}$ and $\tau=w_{0}w_{1}\cdots w_{d}$.
Proof.
Denote by $\bar{\tau}$ the image of $\tau$ by the natural projection $A_{\Gamma}\to W_{\Gamma}$. Since $w_{0},w_{1},\ldots,w_{d}$ are mutually different, the word $w_{0}w_{1}\cdots w_{d}$ is a reduced expression of $\bar{\tau}$. Moreover this is a unique reduced word representing $\bar{\tau}$ because $\bar{\tau}$ is regarded as a Coxeter element in $W_{\{w_{0},w_{1},\ldots,w_{d}\}}$ and two elements $w_{i-1}$ and $w_{i}$ cannot commute for any $i\in\{1,\ldots,d\}$ (see, for example, [39]).
Assume that we have $g\in A_{U\setminus\{s\}}$ such that $gA_{U\setminus\{t\}}=\tau A_{U\setminus\{t\}}$. Then we have $\tau\in A_{U\setminus\{s\}}A_{U\setminus\{t\}}$, and thus $\bar{\tau}\in W_{U\setminus\{s\}}W_{U\setminus\{t\}}$. Take $a_{1},\ldots,a_{d_{1}}\in U\setminus\{s\}$ and $b_{1},\ldots,b_{d_{2}}\in U\setminus\{t\}$ such that the word $a_{1}\cdots a_{d_{1}}b_{1}\cdots b_{d_{2}}$ represents $\bar{\tau}$. Then, by the deletion property for Coxeter groups (see, for example, [3]), the word $a_{1}\cdots a_{d_{1}}b_{1}\cdots b_{d_{2}}$ contains a reduced expression for $\bar{\tau}$ as a subword. In any such subword, the letter $t$ precedes $s$. This contradicts the fact that $w_{0}=s$ precedes $w_{d}=t$ in the unique reduced expression $w_{0}w_{1}\cdots w_{d}$ of $\bar{\tau}$. ∎
4. Proof of Theorem 1.2
We show our main theorem (Theorem 1.2). In fact, we prove the following:
Theorem 4.1.
Let $A_{\Gamma}$ be an Artin group associated with $\Gamma$. Suppose that $\Gamma$ is not a clique. Then, the following are equivalent:
  1. (1)
    $A_{\Gamma}$ is irreducible;
  2. (2)
    $A_{\Gamma}$ has a WPD contracting element with respect to the isometric action on the reduced clique-cube complex;
  3. (3)
    $A_{\Gamma}$ is acylindrically hyperbolic;
  4. (4)
    $A_{\Gamma}$ is virtually directly indecomposable; that is, it cannot be decomposed as a direct product of two infinite subgroups.
In this section, let $A_{\Gamma}$ be an Artin group associated with a defining graph $\Gamma$ with the vertex set $V(\Gamma)=V$ and the edge set $E(\Gamma)=E$, as in Section 2.2. We suppose that $\Gamma$ is not a clique.
4.1. Proof of $(2)\Rightarrow(3),(3)\Rightarrow(4),(4)\Rightarrow(1)$ in Theorem 4.1
We show that $(2)\Rightarrow(3),(3)\Rightarrow(4),(4)\Rightarrow(1)$ in Theorem 4.1.
First, $(2)\Rightarrow(3)$ follows from Theorem 2.3.
Next, we show $(3)\Rightarrow(4)$. Let $A_{\Gamma}$ be acylindrically hyperbolic. If $A_{\Gamma}$ is isomorphic to a direct product $A_{1}\times A_{2}$, then either $A_{1}$ or $A_{2}$ is finite by acylindrical hyperbolicity [33, Corollary 7.3].
Finally, because every non-trivial Artin group is infinite, $(4)\Rightarrow(1)$ is clear.
4.2. Proof of $(1)\Rightarrow(2)$ in Theorem 4.1
We give a proof of $(1)\Rightarrow(2)$ in Theorem 4.1. Suppose that $A_{\Gamma}$ is irreducible; that is, $\Gamma^{t}$ is connected. By noting that $\Gamma$ is not a clique, we consider the unique join decomposition $\Gamma=\Gamma_{0}\ast\Gamma_{\ast}=\Gamma_{0}\ast(\ast_{1\leq i\leq k}\Gamma_{i})$ where $\Gamma_{0}$ is the maximal clique factor and each $\Gamma_{i}$ ($1\leq i\leq k$) is an indecomposable factor with at least two vertices (see Definition 2.18). We set $V_{i}=V(\Gamma_{i})$ and $E_{i}=E(\Gamma_{i})$ for each $i\in\{0,1,\ldots,k\}$. Also, we set $V_{\ast}=V(\Gamma_{\ast})$. Then, for every $i\neq 0$, the complement graph $(\Gamma_{i})^{c}$ of $\Gamma_{i}$ is connected and the complement graph $(\Gamma_{\ast})^{c}$ of $\Gamma_{\ast}$ is the disjoint union of connected components: $(\Gamma_{\ast})^{c}=\bigsqcup_{1\leq i\leq k}(\Gamma_{i})^{c}$, by Lemma 2.17. When $k=1$ and $\Gamma_{0}=\emptyset$, $C_{\Gamma,\Gamma_{0}}=C_{\Gamma}$ and a WPD contracting element is given in [14]. When $k>1$ and $\Gamma_{0}=\emptyset$, a WPD contracting element is given in [29]. Our goal in this section is to give a WPD contracting element with respect to the action on the reduced clique-cube complex $C_{\Gamma,\Gamma_{0}}$ in our general setting. Indeed, we explicitly construct three elements: $\gamma_{\natural}\in A_{V_{\ast}}$, and $\gamma_{\sharp},\gamma_{\flat}\in A_{\Gamma}$ which have some features aligned with Theorem 2.9, and show that their product $\gamma_{\natural\sharp\flat}=\gamma_{\natural}\gamma_{\sharp}\gamma_{\flat}\in A_{\Gamma}$ is indeed a WPD contracting element. It may be useful for understanding the construction of $\gamma_{\natural\sharp\flat}$ to look at Section 5, which treats a concrete example.
We define $Q(\Gamma,\Gamma_{0})$ as a finite simple graph with
(4.1) $&V(Q(\Gamma,\Gamma_{0}))=\{V_{i}\mid 1\leq i\leq k\},\\ &E(Q(\Gamma,\Gamma_{0}))=\left\{(V_{i},V_{j})\middle|\begin{array}[]{l}1\leq i,\ j\leq k,\ i\neq j,\\ (\Gamma_{i})^{t}\text{ and }(\Gamma_{j})^{t}\text{ are connected in }(\Gamma_{0}\ast\Gamma_{i}\ast\Gamma_{j})^{t}\end{array}\right\}.\end{aligned}$
Note that for $i\neq 0$, $(\Gamma_{i})^{t}$ is connected because $(\Gamma_{i})^{c}$ is connected, $V((\Gamma_{i})^{c})=V((\Gamma_{i})^{t})=V_{i}$, and $E((\Gamma_{i})^{c})\subset E((\Gamma_{i})^{t})$. Note that the following are equivalent for different $i,j\in\{1,\ldots,k\}$:
  • (I)
    $(\Gamma_{i})^{t}$ and $(\Gamma_{j})^{t}$ are connected in $(\Gamma_{0}\ast\Gamma_{i}\ast\Gamma_{j})^{t}$;
  • (II)
    there exists a path from a vertex of $\Gamma_{i}$ and a vertex of $\Gamma_{j}$ on $\Gamma_{0}\ast\Gamma_{i}\ast\Gamma_{j}$, consisting of edges with label greater than $2$.
Because $A_{\Gamma}$ is irreducible, i.e., $\Gamma^{t}$ is connected, $Q(\Gamma,\Gamma_{0})$ is connected.
We take a spanning tree $T$ of $Q(\Gamma,\Gamma_{0})$. We regard $T$ as a rooted tree with the root $V_{1}$. By trading indices of $V_{2},\ldots,V_{k}$ if necessary, we suppose that $i<j$ only if $V_{i}$ is not farther than $V_{j}$ from $V_{1}$ in $T$.
For each $i,j$ with $i<j$ and $(V_{i},V_{j})\in E(T)$, take a path on $\Gamma$;
(4.2) $p_{i,j}=(w_{i,j,0},w_{i,j,1},\ldots,w_{i,j,d_{i,j}})$
consisting of edges with label greater than $2$ such that $w_{i,j,0}\in V_{i}$, $w_{i,j,d_{i,j}}\in V_{j}$ and $w_{i,j,1},\ldots,w_{i,j,d_{i,j}-1}\in V_{0}$. Moreover, set $d_{j,i}:=d_{i,j}$ and a path on $\Gamma$;
(4.3) $p_{j,i}=(w_{j,i,0},w_{j,i,1},\ldots,w_{j,i,d_{j,i}}):=(w_{i,j,d_{i,j}},\ldots,w_{i,j,1},w_{i,j,0}).$
For any $i,j\in\{1,\ldots,k\}$ with $(V_{i},V_{j})\in E(T)$, set
(4.4) $s_{i,j}:=w_{i,j,0}\in V_{i},\ t_{i,j}:=w_{i,j,d_{i,j}}\in V_{j}$
and
(4.5) $\tau_{i,j}=w_{i,j,0}w_{i,j,1}\cdots w_{i,j,d_{i,j}}\in A_{\Gamma}.$
Then we have $s_{j,i}=t_{i,j}$ in $V_{j}$, $t_{j,i}=s_{i,j}$ in $V_{i}$ and $\tau_{j,i}=w_{i,j,d_{i,j}}\cdots w_{i,j,1}w_{i,j,0}\in A_{\Gamma}.$
When $k\geq 2$, we take a closed path
(4.6) $(V_{i_{1}},\dots,V_{i_{r}},V_{i_{r+1}})$
with $i_{1}=i_{r+1}=1$ on the spanning tree $T$ of $Q(\Gamma,\Gamma_{0})$ passing through every vertex at least once. Since $r$ is the length of the closed path on the tree, $r$ is an even number. If $k=1$, then $T$ has no edges, and we set $r=0$.
To construct $\gamma_{\natural}$ and $\gamma_{\flat}$, we start with the following lemma, which is almost the same as [29, Lemma 6.1]. See Section 2.6 for notations related to matrices.
Lemma 4.2.
There exist $n\in\mathbb{N}$, a matrix $M_{\natural}=(v_{i,l})\in\mathcal{M}(k,n;V)$, and a map
(4.7) $l_{\flat}:\left\{(i,j)\middle|\begin{array}[]{l}1\leq i,j\leq k,\\ (V_{i},V_{j})\in E(T)\end{array}\right\}\to\{1,\ldots,n\}$
which satisfy the following:
  • (i)
    For each $i\in\{1,\ldots,k\}$, $(v_{i,1},\ldots,v_{i,n},v_{i,1})$ is a closed path on $(\Gamma_{i})^{c}$ that passes through every vertex at least once;
  • (ii)
    For each $i,j\in\{1,\ldots,k\}$ with $(V_{i},V_{j})\in E(T)$, $v_{i,l_{\flat}(i,j)}=s_{i,j}$, $v_{j,l_{\flat}(i,j)}=t_{i,j}$;
  • (iii)
    $l_{\flat}({j,i})=l_{\flat}(i,j)$.
Proof.
For any $i\in\{1,\ldots,k\}$, we consider the length $n_{i}$ of a closed path on $(\Gamma_{i})^{c}$ passing through every vertex at least once. Set $n=\prod_{1\leq i\leq k}n_{i}$. Then, for any $i\in\{1,\ldots,k\}$, by concatenating $\frac{n}{n_{i}}$ copies of a closed path of length $n_{i}$ on $(\Gamma_{i})^{c}$ passing through every vertex at least once, we have a closed path $(v^{\prime}_{i,1},\ldots,v^{\prime}_{i,n},v^{\prime}_{i,n+1})$ with $v^{\prime}_{i,1}=v^{\prime}_{i,n+1}$ on $(\Gamma_{i})^{c}$ passing through every vertex at least $\frac{n}{n_{i}}$ times (in particular, at least once).
We set $(v_{1,1},\ldots,v_{1,n},v_{1,n+1}):=(v^{\prime}_{1,1},\ldots,v^{\prime}_{1,n},v^{\prime}_{1,n+1})$. For any $j\in\{2,\ldots,k\}$, we define $(v_{j,1},\ldots,v_{j,n},v_{j,n+1})$ inductively as follows. Take $j\in\{2,\ldots,k\}$. Suppose that $(v_{i,1},\ldots,v_{i,n},v_{i,n+1})$ is defined for $i\in\{1,\ldots,j-1\}$ with $(V_{i},V_{j})\in E(T)$. Note that such $i$ is unique for $j$. Then, we define $l_{\flat}(i,j)$ as the minimum $l$ such that $v_{i,l}=s_{i,j}$. By a cyclic permutation of $v^{\prime}_{j,1},\ldots,v^{\prime}_{j,n}$, we have $v_{j,1},\ldots,v_{j,n}$ such that $v_{j,l_{\flat}(i,j)}=t_{i,j}$. By setting $v_{j,n+1}:=v_{j,1}$, we have $(v_{j,1},\ldots,v_{j,n},v_{j,n+1})$. Finally, for any $i,j$ with $1\leq i<j\leq k$ and $(V_{i},V_{j})\in E(T)$, we set $l_{\flat}(j,i):=l_{\flat}(i,j)$. ∎
We take $n\in\mathbb{N}$, $M_{\natural}=(v_{i,l})\in\mathcal{M}(k,n;V)$ and a map $l_{\flat}$ as in Lemma 4.2. In the case where $k=1$, $E(T)=\emptyset$, and thus $l_{\flat}$ and the conditions (ii), (iii) are meaningless. Nonetheless, the condition (i) remains meaningful.
Now we define $\gamma_{\natural}\in A_{V_{\ast}}$ as
(4.8) $\gamma_{\natural}:=prod(M_{\natural})=(v_{1,1}v_{2,1}\cdots v_{k,1})\cdots(v_{1,n}v_{2,n}\cdots v_{k,n}).$
When $\Gamma_{0}=\emptyset$ and $k=1$, this $\gamma_{\natural}$ is a WPD contracting element on $C_{\Gamma,\Gamma_{0}}=C_{\Gamma}$ [14]. We do not directly need this fact. Even in the general setting, we can confirm that $\gamma_{\natural}$ fits the condition (i) of Theorem 2.9. Unfortunately, it is unclear whether $\gamma_{\natural}$ fits the conditions (ii) and (iii). To more easily approach these conditions, we prepare two modifications $\gamma_{\sharp}$ and $\gamma_{\flat}$ based on powers of $\gamma_{\natural}$.
We prepare for the definition of $\gamma_{\flat}$. Let $e$ be the identity in $A_{\Gamma}$. When $k=1$ (for example, $\Gamma$ in Figure 4), we define $\gamma_{\flat}=e$. Suppose that $k\geq 2$. Recall that we have $n\in\mathbb{N}$, $M_{\natural}=(v_{i,l})\in\mathcal{M}(k,n;V)$, a map $l_{\flat}$ as in Lemma 4.2, and a closed path $(V_{i_{1}},\dots,V_{i_{r}},V_{i_{r+1}})$ with $i_{1}=i_{r+1}=1$, defined in (4.6). Note that $r\geq 2$, since $k\geq 2$. We define supplementary matrices $M^{\prime}_{\flat}(a),M^{\prime\prime}_{\flat}(a)\in\mathcal{M}(k,n;A_{\Gamma})$ $(a\in\{1,\ldots,r\})$ by
(4.9) $\left(M^{\prime}_{\flat}(a)\right)[i,l]=\begin{cases}v_{i,l}^{-1}\quad&(i\in\{i_{a},i_{a+1}\},l=l_{\flat}(i_{a},i_{a+1}))\\ e\quad&(\text{otherwise}),\end{cases}$
(4.10) $\left(M^{\prime\prime}_{\flat}(a)\right)[i,l]=\begin{cases}\tau_{i_{a},i_{a+1}}\quad&(i=1,l=l_{\flat}(i_{a},i_{a+1}))\\ e\quad&(\text{otherwise}).\end{cases}$
Using $M^{\prime}_{\flat}(a)$ and $M^{\prime\prime}_{\flat}(a)$, we define $M_{\flat}(a)\in\mathcal{M}(k,n;A_{\Gamma})$ by
(4.11) $M_{\flat}(a):=M^{\prime\prime}_{\flat}(a)\odot M^{\prime}_{\flat}(a)\odot M_{\natural},$
[…]
(4.12) $M_{\flat}=cbind\left(M_{\flat}(1),\ldots,M_{\flat}(r)\right).$
For the definition of $cbind$, see (2.2).
We now define $\gamma_{\flat}\in A_{\Gamma}$ as
(4.13) $\gamma_{\flat}:=prod(M_{\flat})=prod(M_{\flat}(1))\cdots prod(M_{\flat}(r)).$
For the definition of the map $prod$, see (2.5) and (2.6). $\gamma_{\flat}$ is a modification of $\gamma_{\natural}^{r}$. When $k\geq 2$ and $\Gamma_{0}=\emptyset$, we can take $\gamma_{\flat}$ as a WPD element [29]. This fact is not used directly. To deal with the case where $\Gamma_{0}\neq\emptyset$, we need the other modification, $\gamma_{\sharp}$.
To explain the construction of $\gamma_{\sharp}$, we introduce the “height” $h$ of elements of $V_{0}$ in the case $V_{0}\neq\emptyset$. Otherwise, we define $\gamma_{\sharp}=e$. For $h\geq 1$, we set
(4.14) $W(h):=\{w\in V_{0}\mid\min_{v\in V_{\ast}}d_{\Gamma^{t}}(w,v)=h\}.$
Since $\Gamma^{t}$ is connected, we have $\mathfrak{h}\in\mathbb{N}$ such that $W(\mathfrak{h})\neq\emptyset$ and $V_{0}=W(1)\sqcup\cdots\sqcup W(\mathfrak{h})$. Suppose that $\mathfrak{h}\geq 2$. Otherwise, we define $\gamma_{\sharp}=e$. We define two maps
(4.15) $i_{\sharp}:V_{0}\to\{1,\ldots,k\},$
(4.16) $l_{\sharp}:V_{0}\to\{1,\ldots,n\}$
such that $(i_{\sharp}(w),l_{\sharp}(w))$ is the minimum of $(i,l)$ indexing $v_{i,l}\in V_{*}$ with
$d_{\Gamma^{t}}(w,v_{i,l})=\min_{v\in V_{\ast}}d_{\Gamma^{t}}(w,v)$
with respect to the order $<$ on $\{1,\ldots,k\}\times\{1,\ldots,n\}$ defined by
$(i,l)<(i^{\prime},l^{\prime})\Leftrightarrow(l<l^{\prime})\text{ or }(l=l^{\prime}\text{ and }i<i^{\prime}).$
Note that this is a slightly unusual dictionary order, which considers the second factor first. In particular, for $w\in W(1)$, $(i_{\sharp}(w),l_{\sharp}(w))$ is the minimum of $(i,l)$ with $\mu(w,v_{i,l})>2$ with respect to the above order.
We set
(4.17) $W(0):=\{v_{i_{\sharp}(w),l_{\sharp}(w)}\in V_{\ast}\mid w\in W(1)\},$
which equals $\{v_{i_{\sharp}(w),l_{\sharp}(w)}\in V_{\ast}\mid w\in V_{0}\}$. Obviously, we have $W(0)\cap V_{0}=\emptyset$. We extend two maps $i_{\sharp}$ and $l_{\sharp}$ on $W(0)\sqcup V_{0}$ by defining $(i_{\sharp}(w),l_{\sharp}(w))$ for $w\in W(0)$ as the minimum of $(i,l)$ with $w=v_{i,l}$ with respect to the order $<$ on $\{1,\ldots,k\}\times\{1,\ldots,n\}$. Note that for $w\in W(0)$, $i$ and $l$ with $w=v_{i,l}$, $i$ is unique, but $l$ is not necessarily unique.
Next, by using the order $<$ on $\{1,\ldots,k\}\times\{1,\ldots,n\}$ and the two maps $i_{\sharp}$, $l_{\sharp}$ on $W(0)\sqcup V_{0}=W(0)\sqcup W(1)\sqcup\cdots\sqcup W(\mathfrak{h})$, we define a total order $\prec$ on $W(0)\sqcup V_{0}$ as follows. First, we take a total order $\prec$ on $W(0)$ such that for any $w,w^{\prime}\in W(0)$, $w\prec w^{\prime}$ if and only if $(i_{\sharp}(w),l_{\sharp}(w))<(i_{\sharp}(w^{\prime}),l_{\sharp}(w^{\prime}))$. Next, suppose that for $1\leq h\leq\mathfrak{h}$, $W(0)\sqcup\cdots\sqcup W(h-1)$ is equipped with a total order $\prec$. Then we define a map
$\pi_{h}:W(h)\to W(h-1)$
as $\pi_{h}(w)$ $(w\in W(h))$ is the minimum of $w^{\prime}\in W(h-1)$ with $\mu(w,w^{\prime})>2$ with respect to the given total order $\prec$. We take a total order $\prec$ on $W(h)$ such that for any different $w,w^{\prime}\in W(h)$, $w\prec w^{\prime}$ only if $\pi(w)\preceq\pi(w^{\prime})$. Moreover, for any $w\in W(0)\sqcup\cdots\sqcup W(h-1)$ and $w^{\prime}\in W(h)$, we set $w\prec w^{\prime}$. Then $\prec$ on $W(0)\sqcup\cdots\sqcup W(h-1)$ is extended on $W(0)\sqcup\cdots\sqcup W(h)$. Finally, we have the order $\prec$ on $W(0)\sqcup V_{0}=W(0)\sqcup\cdots\sqcup W(\mathfrak{h})$. We define a map
(4.18) $\pi:W(1)\sqcup\cdots\sqcup W(\mathfrak{h})\to W(0)\sqcup\cdots\sqcup W(\mathfrak{h}-1),$
by $\pi(w)=\pi_{h}(w)\in W(h-1)$ for $w\in W(h)$. For any $w\in W(0)\sqcup V_{0}$, let $h\in\{0,\ldots,\mathfrak{h}\}$ be the height of $w$; that is, $w\in W(h)$. We take a path from $w$ to $v_{i_{\sharp}(w),l_{\sharp}(w)}$ on $\Gamma^{t}$;
(4.19) $(w(h),w(h-1),\ldots,w(0)):=(w,\pi(w),\pi^{2}(w),\ldots,\pi^{h}(w)).$
We note that $(i_{\sharp}(w),l_{\sharp}(w))=(i_{\sharp}(w(h^{\prime})),l_{\sharp}(w(h^{\prime})))$ for any $0\leq h^{\prime}\leq h$. In particular, $w(0)=\pi^{h}(w)=v_{i_{\sharp}(w),l_{\sharp}(w)}$. See Figure 7.
Figure 7. In the figure, $w$ is a vertex in $W(1)$, and thin solid lines indicate edges labeled by $2$. A thick solid line, which connects $w$ and $v_{i_{\sharp}(w),l_{\sharp}(w)}$, indicates an edge labeled by an integer greater than $2$. For all edges which are presented as dotted lines, we do not care about their label.
Now we name all elements of $W(1)\sqcup\cdots\sqcup W(\mathfrak{h}-1)=V_{0}\setminus W(\mathfrak{h})$ as
(4.20) $w_{1},w_{2},\ldots,w_{\mathfrak{m}}$
such that $w_{1}\prec w_{2}\prec\cdots\prec w_{\mathfrak{m}}$ where
(4.21) $\mathfrak{m}:=\#(W(1)\sqcup\cdots\sqcup W(\mathfrak{h}-1)).$
Moreover, we set the heights $h_{1},h_{2},\ldots,h_{\mathfrak{m}}$ for $w_{1},w_{2},\ldots,w_{\mathfrak{m}}$, respectively. We have $11=h_{1}\leq h_{2}\leq\cdots\leq h_{\mathfrak{m}}=\mathfrak{h}-$ 1 . The following is clear by the construction of the order $\prec$ and the map $\pi$.
Lemma 4.3.
Let $h\in\{1,\ldots,\mathfrak{h}\}$ and $w\in W(h)$.
  1. (1)
    Let $h=1$. Let $w^{\prime}\in V_{0}$ with $w^{\prime}\succeq w$. For any $(i,l)\in\{1,\ldots,k\}\times\{1,\ldots,n\}$ with $(i,l)<(i_{\sharp}(w),l_{\sharp}(w))$, we have $\mu(v_{i,l},w^{\prime})=2$.
  2. (2)
    Let $h\geq 2$. Let $w^{\prime}\in V_{0}$ with $w^{\prime}\succeq w$.
    1. (i)
      For any $v\in V_{\ast}$, we have $\mu(v,w^{\prime})=2$.
    2. (ii)
      For any $w^{\prime\prime}\in V_{0}$ with $w^{\prime\prime}\prec\pi(w)$, we have $\mu(w^{\prime\prime},w^{\prime})=2$.
See Figure 7.
We now explain the construction of $\gamma_{\sharp}$. If $V_{0}=\emptyset$ or $\mathfrak{h}=1$, then we have already defined $\gamma_{\sharp}=e$. Suppose that $V_{0}\neq\emptyset$ and that $\mathfrak{h}\geq 2$. We define supplementary matrices $M^{\prime}_{\sharp}(\mathfrak{j}),M^{\prime\prime}_{\sharp}(\mathfrak{j})\in\mathcal{M}(k,n;A_{\Gamma})$ $(\mathfrak{j}\in\{1,\ldots,\mathfrak{m}\})$ by
(4.22) $\left(M^{\prime}_{\sharp}(\mathfrak{j})\right)[i,l]=\begin{cases}v_{i,l}^{-1}&\quad(i=i_{\sharp}(w_{\mathfrak{j}}),l=l_{\sharp}(w_{\mathfrak{j}}))\\ e&\quad(\text{otherwise})\end{cases}$
and
(4.23) $\left(M^{\prime\prime}_{\sharp}(\mathfrak{j})\right)[i,l]=\begin{cases}w_{\mathfrak{j}}\left(h_{\mathfrak{j}}\right)w_{\mathfrak{j}}\left(h_{\mathfrak{j}}-1\right)\cdots w_{\mathfrak{j}}(0)&\quad(i=k,l=l_{\sharp}(w_{\mathfrak{j}}))\\ e&\quad(\text{otherwise}),\end{cases}$
where $h_{\mathfrak{j}}$ is the height of $w_{\mathfrak{j}}$. Using $M^{\prime}_{\sharp}(\mathfrak{j})$ and $M^{\prime\prime}_{\sharp}(\mathfrak{j})$, we define $M_{\sharp}(\mathfrak{j})\in\mathcal{M}(k,n;A_{\Gamma})$ by
(4.24) $M_{\sharp}(\mathfrak{j}):=M_{\natural}\odot M^{\prime}_{\sharp}(\mathfrak{j})\odot M^{\prime\prime}_{\sharp}(\mathfrak{j}),$
[…]
(4.25) $M_{\sharp}=cbind\left(M_{\sharp}(1),\ldots,M_{\sharp}(\mathfrak{m})\right).$
For the definition of $cbind$, see (2.2). We now define $\gamma_{\sharp}\in A_{\Gamma}$ as
(4.26) $\gamma_{\sharp}:=prod(M_{\sharp})=prod(M_{\sharp}(1))\cdots prod(M_{\sharp}(\mathfrak{m})).$
$\gamma_{\sharp}$ is a modification of $\gamma_{\natural}^{\mathfrak{m}}$.
We define $M_{\natural\sharp\flat}\in\mathcal{M}(k,(1+\mathfrak{m}+r)n;A_{\Gamma})$ by
(4.27) $M_{\natural\sharp\flat}:=cbind(M_{\natural},M_{\sharp},M_{\flat}).$
Now, we define $\gamma_{\natural\sharp\flat}\in A_{\Gamma}$ by
(4.28) $\gamma_{\natural\sharp\flat}:=prod(M_{\natural\sharp\flat})=prod(M_{\natural})prod(M_{\sharp})prod(M_{\flat})=\gamma_{\natural}\gamma_{\sharp}\gamma_{\flat}.$
For later reference, we define supplementary elements $\lambda_{l}$ $(l\in\{1,\ldots,n\})$, $\lambda_{\sharp}(\mathfrak{j})$, $\gamma_{\sharp}(\mathfrak{j})$ $(\mathfrak{j}\in\{1,\ldots,\mathfrak{m}\})$, $\lambda_{\flat}(a)$, $\gamma_{\flat}(a)$ $(a\in\{1,\ldots,r\})$, $\gamma_{\natural\sharp\flat}[d]\in A_{\Gamma}$ $(d\in\{1,\ldots,(1+r+\mathfrak{m})n\})$. For every $l\in\{1,\ldots,n\}$, we define $\lambda_{l}\in A_{V_{\ast}}\subset A_{\Gamma}$ by
(4.29) $\lambda_{l}:=prod(M_{\natural}[\ ,l])=v_{1,l}v_{2,l}\cdots v_{i,l}\cdots v_{k,l}.$
For every $\mathfrak{j}\in\{1,\ldots,\mathfrak{m}\}$, we define $\lambda_{\sharp}(\mathfrak{j}),\gamma_{\sharp}(\mathfrak{j})\in A_{\Gamma}$ by
(4.30) $\lambda_{\sharp}(\mathfrak{j})$ $:=prod(M_{\sharp}(\mathfrak{j})[\ ,l_{\sharp}(w_{\mathfrak{j}})]),$
(4.31) $\gamma_{\sharp}(\mathfrak{j})$ $:=prod(M_{\sharp}(\mathfrak{j})).$
For every $a\in\{1,\ldots,r\}$, we define $\lambda_{\flat}(a),\gamma_{\flat}(a)\in A_{\Gamma}$ by
(4.32) $\lambda_{\flat}(a)$ $:=prod(M_{\flat}(a)[\ ,l_{\flat}(i_{a},i_{a+1})]),$
(4.33) $\gamma_{\flat}(a)$ $:=prod(M_{\flat}(a)).$
Note that $\lambda_{\sharp}(\mathfrak{j})$ and $\lambda_{\flat}(a)$ can be regarded as modifications of $\lambda_{l_{\sharp}(w_{\mathfrak{j}})}$ and $\lambda_{l_{\flat}(i_{a},i_{a+1})}$, respectively. Also note that $\gamma_{\sharp}(\mathfrak{j})$ and $\gamma_{\flat}(a)$ can be regarded as modifications of $\gamma_{\natural}$. For every $d\in\{1,\ldots,(1+\mathfrak{m}+r)n\}$, let $\gamma_{\natural\sharp\flat}[d]$ be the product of the components of the first $d$ columns of $M_{\natural\sharp\flat}\in\mathcal{M}(k,(1+\mathfrak{m}+r)n;A_{\Gamma})$. That is,
(4.34) $\gamma_{\natural\sharp\flat}[d]:=prod(M_{\natural\sharp\flat}[\ ,1:d]).$
We recall the following for convenience:
  1. (1)
    $k$ is the number of indecomposable factors of the unique join decomposition of $\Gamma_{\ast}$;
  2. (2)
    $r$ is the length of the closed path (4.6) on $T$;
  3. (3)
    $n$ is the common length of the closed paths on $(\Gamma_{i})^{c}$ for all $i\in\{1,\ldots,k\}$ taken in Lemma 4.2;
  4. (4)
[…]
For $1\leq l\leq n$, set
$U_{l}:=V_{0}\sqcup\{v_{1,l},v_{2,l},\ldots,v_{k,l}\}.$
Then, $U_{l}$ spans a clique containing $\Gamma_{0}$ in $\Gamma$; that is, $U_{l}\in\Delta_{\Gamma,\Gamma_{0}}$. Hence, we have a $k$-dimensional cube $[A_{V_{0}},A_{U_{l}}]$ in $C_{\Gamma,\Gamma_{0}}$. The hyperplanes dual to edges of $[A_{V_{0}},A_{U_{l}}]$ are of $v_{i,l}$-type ($i\in\{1,\ldots,k\}$). We denote such hyperplanes as $H_{i,l}$ ($i\in\{1,\ldots,k\}$).
The following is a variant for $C_{\Gamma,\Gamma_{0}}$ of [29, Lemma 6.2] for $C_{\Gamma}$.
Lemma 4.4.
For $i\in\{1,\ldots,k\}$, $l\in\{1,\ldots,n\}$, $\mathfrak{j}\in\{1,\ldots,\mathfrak{m}\}$ and $a\in\{1,\ldots,r\}$, we have the following:
  1. (1)
    $H_{i,l}\cap H_{i,l+1}=\emptyset$, where we set $H_{i,n+1}:=H_{i,1}$;
  2. (2)
    $H_{i,l}\cap\lambda_{l}H_{i,l}=\emptyset$;
  3. (3)
    $H_{i,l_{\flat}(i_{a},i_{a+1})}\cap\lambda_{\flat}(a)H_{i,l_{\flat}(i_{a},i_{a+1})}=\emptyset$;
  4. (4)
    $H_{i,l_{\sharp}(w_{\mathfrak{j}})}\cap\lambda_{\sharp}(\mathfrak{j})H_{i,l_{\sharp}(w_{\mathfrak{j}})}=\emptyset$;
  5. (5)
    $[A_{V_{0}},A_{U_{l}}]\cap[A_{V_{0}},A_{U_{l+1}}]=\{A_{V_{0}}\}$, where we set $U_{n+1}:=U_{1}$;
  6. (6)
    $[A_{V_{0}},A_{U_{l}}]\cap\lambda_{l}[A_{V_{0}},A_{U_{l}}]=\{A_{U_{l}}\}$;
  7. (7)
    $[A_{V_{0}},A_{U_{l_{\flat}(i_{a},i_{a+1})}}]\cap\lambda_{\flat}(a)[A_{V_{0}},A_{U_{l_{\flat}(i_{a},i_{a+1})}}]=\{A_{U_{l_{\flat}(i_{a},i_{a+1})}}\}$.
  8. (8)
    $[A_{V_{0}},A_{U_{l_{\sharp}(w_{\mathfrak{j}})}}]\cap\lambda_{\sharp}(\mathfrak{j})[A_{V_{0}},A_{U_{l_{\sharp}(w_{\mathfrak{j}})}}]=\{A_{U_{l_{\sharp}(w_{\mathfrak{j}})}}\}$.
Proof.
(1) $H_{i,l}$ and $H_{i,l+1}$ are of $v_{i,l}$-type and $v_{i,l+1}$-type, respectively. Note that $v_{i,l}\neq v_{i,l+1}$ implies $H_{i,l}\neq H_{i,l+1}$. Because $(v_{i,l},v_{i,l+1})\notin E$, we have $H_{i,l}\cap H_{i,l+1}=\emptyset$ (see Remark 2.14).
(2) Because $[A_{U_{l}\setminus\{v_{i,l}\}},A_{U_{l}}]\subset N(H_{i,l})$, we have $[\lambda_{l}A_{U_{l}\setminus\{v_{i,l}\}},\lambda_{l}A_{U_{l}}]\subset\lambda_{l}N(H_{i,l})$. Note that $\lambda_{l}A_{U_{l}}=A_{U_{l}}$ and $\lambda_{l}A_{U_{l}\setminus\{v_{i,l}\}}=v_{1,l}\cdots v_{i,l}A_{U_{l}\setminus\{v_{i,l}\}}$. Assume that $H_{i,l}\cap\lambda_{l}H_{i,l}\neq\emptyset$. Because $H_{i,l}$ and $\lambda_{l}H_{i,l}$ are of $v_{i,l}$-type, we see that $H_{i,l}=\lambda_{l}H_{i,l}$ (see Remark 2.14). Then, we have
$[\lambda_{l}A_{U_{l}\setminus\{v_{i,l}\}},\lambda_{l}A_{U_{l}}]=[v_{1,l}\cdots v_{i,l}A_{U_{l}\setminus\{v_{i,l}\}},A_{U_{l}}]\subset N(H_{i,l}).$
Hence, we obtain $A_{U_{l}\setminus\{v_{i,l}\}}=v_{1,l}\cdots v_{i,l}A_{U_{l}\setminus\{v_{i,l}\}}$. Hence, $A_{U_{l}\setminus\{v_{i,l}\}}=v_{i,l}A_{U_{l}\setminus\{v_{i,l}\}}$. This means that $v_{i,l}\in A_{U_{l}\setminus\{v_{i,l}\}}$. However, from Lemma 2.10 (2), $\{v_{i,l}\}\cap(U_{l}\setminus\{v_{i,l}\})=\emptyset$ implies that $A_{\{v_{i,l}\}}\cap A_{U_{l}\setminus\{v_{i,l}\}}=A_{\emptyset}=\{e\}$. This contradicts $v_{i,l}\neq e$ in $A_{\Gamma}$.
(3) When $i\neq i_{a},i_{a+1}$, by the same argument as in (2), we see that
$H_{i,l_{\flat}(i_{a},i_{a+1})}\cap\lambda_{\flat}(a)H_{i,l_{\flat}(i_{a},i_{a+1})}=\emptyset.$
Now, assume that $i=i_{a}$ or $i=i_{a+1}$ and
$H_{i,l_{\flat}(i_{a},i_{a+1})}\cap\lambda_{\flat}(a)H_{i,l_{\flat}(i_{a},i_{a+1})}\neq\emptyset.$
Note that $\lambda_{\flat}(a)A_{U_{l_{\flat}(i_{a},i_{a+1})}}=A_{U_{l_{\flat}(i_{a},i_{a+1})}}$ and
$\lambda_{\flat}(a)A_{U_{l_{\flat}(i_{a},i_{a+1})}\setminus\{v_{i,l_{\flat}(i_{a},i_{a+1})}\}}=\tau_{i_{a},i_{a+1}}A_{U_{l_{\flat}(i_{a},i_{a+1})}\setminus\{v_{i,l_{\flat}(i_{a},i_{a+1})}\}}.$
Because both $H_{i,l_{\flat}(i_{a},i_{a+1})}$ and $\lambda_{\flat}(a)H_{i,l_{\flat}(i_{a},i_{a+1})}$ are of $v_{i,l_{\flat}(i_{a},i_{a+1})}$-type, we see that $H_{i,l_{\flat}(i_{a},i_{a+1})}=\lambda_{\flat}(a)H_{i,l_{\flat}(i_{a},i_{a+1})}$. Then, we have
$\begin{split}[\lambda_{\flat}(a)A_{U_{l_{\flat}(i_{a},i_{a+1})}\setminus\{v_{i,l_{\flat}(i_{a},i_{a+1})}\}},\lambda_{\flat}(a)A_{U_{l_{\flat}(i_{a},i_{a+1})}}]\\ =[\tau_{i_{a},i_{a+1}}A_{U_{l_{\flat}(i_{a},i_{a+1})}\setminus\{v_{i,{l_{\flat}(i_{a},i_{a+1})}}\}},A_{U_{l_{\flat}(i_{a},i_{a+1})}}]\subset N(H_{i,{l_{\flat}(i_{a},i_{a+1})}}).\end{split}$
Hence, we have
$\tau_{i_{a},i_{a+1}}A_{U_{l_{\flat}(i_{a},i_{a+1})}\setminus\{v_{i,l_{\flat}(i_{a},i_{a+1})}\}}=A_{U_{l_{\flat}(i_{a},i_{a+1})}\setminus\{v_{i,l_{\flat}(i_{a},i_{a+1})}\}}.$
This contradicts Lemma 3.9.
(4) is shown by a similar manner to (2).
Parts (5), (6), (7) and (8) follow from (1), (2), (3) and (4), respectively. ∎
The following is a variant for $C_{\Gamma,\Gamma_{0}}$ of [29, Lemma 6.3] for $C_{\Gamma}$, and its proof is the same. In fact, Lemma 3.9 plays the same role that [29, Lemma 5.2] did in the proof of [29, Lemma 6.3].
Lemma 4.5.
For $a\in\{1,\ldots,r\}$, we have
$H_{i_{a},l_{\flat}(i_{a},i_{a+1})}\cap\lambda_{\flat}(a)H_{i_{a+1},l_{\flat}(i_{a},i_{a+1})}=\emptyset.$
Figure 8. All hyperplanes in Lemma 4.6 for the case where $\Gamma=\Gamma_{0}\ast\Gamma_{1}\ast\Gamma_{2}$. Hyperplanes $J_{1,d}$ and $J_{2,d}$ intersecting $K_{d}$ are represented as the solid and the dotted lines, respectively.
Noting Lemma 4.4, for any $i\in\{1,\ldots,k\}$, we define a sequence of hyperplanes
$\ldots,J_{i,-1},J_{i,0},J_{i,1},\ldots,J_{i,2(1+\mathfrak{m}+r)n},J_{i,2(1+\mathfrak{m}+r)n+1},\ldots,$
and a sequence of $k$-dimensional cubes
$\ldots,K_{-1},K_{0},K_{1},\cdots,K_{2(1+\mathfrak{m}+r)n},K_{2(1+\mathfrak{m}+r)n+1},\ldots$
as follows. First, for $a\in\{1,\ldots,1+\mathfrak{m}+r\}$ and $l\in\{1,\dots,n\}$, we define
$J_{i,2((a-1)n+l)-1}$ $:=\gamma_{\natural\sharp\flat}[(a-1)n+l-1]H_{i,l},$
$J_{i,2((a-1)n+l)}$ $:=\gamma_{\natural\sharp\flat}[(a-1)n+l]H_{i,l},$
$K_{2((a-1)n+l)-1}$ $:=\gamma_{\natural\sharp\flat}[(a-1)n+l-1][A_{V_{0}},A_{U_{l}}]$
$=[\gamma_{\natural\sharp\flat}[(a-1)n+l-1]A_{V_{0}},\gamma_{\natural\sharp\flat}[(a-1)n+l-1]A_{U_{l}}],$
$K_{2((a-1)n+l)}$ $:=\gamma_{\natural\sharp\flat}[(a-1)n+l][A_{V_{0}},A_{U_{l}}]$
$=[\gamma_{\natural\sharp\flat}[(a-1)n+l]A_{V_{0}},\gamma_{\natural\sharp\flat}[(a-1)n+l]A_{U_{l}}]$
$=[\gamma_{\natural\sharp\flat}[(a-1)n+l]A_{V_{0}},\gamma_{\natural\sharp\flat}[(a-1)n+l-1]A_{U_{l}}].$
Here, we note that both $J_{i,2((a-1)n+l)-1}$ and $J_{i,2((a-1)n+l)}$ are of the same $v_{i,l}$-type, and that we have $\gamma_{\natural\sharp\flat}[(a-1)n+l-1]A_{U_{l}}=\gamma_{\natural\sharp\flat}[(a-1)n+l]A_{U_{l}}$. See Figure 8.
Second, for any $b\in\{1,\dots,2(1+\mathfrak{m}+r)n\}$ and $c\in\mathbb{Z}$, we set
$J_{i,2(1+\mathfrak{m}+r)nc+b}:={(\gamma_{\natural\sharp\flat})}^{c}J_{i,b},$
$K_{2(1+\mathfrak{m}+r)nc+b}:={(\gamma_{\natural\sharp\flat})}^{c}K_{b}.$
Additionally, we have the two connected components $J_{i,2(1+\mathfrak{m}+r)nc+b}^{-}$ and $J_{i,2(1+\mathfrak{m}+r)nc+b}^{+}$ such that
$C_{\Gamma}\!\setminus\!\!\setminus J_{i,2(1+\mathfrak{m}+r)nc+b}=J_{i,2(1+\mathfrak{m}+r)nc+b}^{-}\sqcup J_{i,2(1+\mathfrak{m}+r)nc+b}^{+},$
and the “left endpoint” of $K_{2(1+\mathfrak{m}+r)nc+b}$ is in $J_{i,2(1+\mathfrak{m}+r)nc+b}^{-}$ while the “right endpoint” of $K_{2(1+\mathfrak{m}+r)nc+b}$ is in $J_{i,2(1+\mathfrak{m}+r)nc+b}^{+}$. Precisely, we have
${(\gamma_{\natural\sharp\flat})}^{c}\left(\gamma_{\natural\sharp\flat}\left[\frac{b-1}{2}\right]\right)A_{V_{0}}\in J_{i,2(1+\mathfrak{m}+r)nc+b}^{-},$
${(\gamma_{\natural\sharp\flat})}^{c}\left(\gamma_{\natural\sharp\flat}\left[\frac{b-1}{2}\right]\right)A_{U_{l}}\in J_{i,2(1+\mathfrak{m}+r)nc+b}^{+}$
if $b=2((a-1)n+l)-1$ with some $a\in\{1,\ldots,1+\mathfrak{m}+r\}$ and $l\in\{1,\dots,n\}$, and we have
${(\gamma_{\natural\sharp\flat})}^{c}\left(\gamma_{\natural\sharp\flat}\left[\frac{b}{2}\right]\right)A_{U_{l}}\in J_{i,2(1+\mathfrak{m}+r)nc+b}^{-},$
${(\gamma_{\natural\sharp\flat})}^{c}\left(\gamma_{\natural\sharp\flat}\left[\frac{b}{2}\right]\right)A_{V_{0}}\in J_{i,2(1+\mathfrak{m}+r)nc+b}^{+}$
if $b=2((a-1)n+l)$ with some $a\in\{1,\ldots,1+\mathfrak{m}+r\}$ and $l\in\{1,\dots,n\}$. Note that in the former case,
$K_{2(1+\mathfrak{m}+r)nc+b}$ $={(\gamma_{\natural\sharp\flat})}^{c}\left(\gamma_{\natural\sharp\flat}\left[\frac{b-1}{2}\right]\right)[A_{V_{0}},A_{U_{l}}]$
$=\left[{(\gamma_{\natural\sharp\flat})}^{c}\left(\gamma_{\natural\sharp\flat}\left[\frac{b-1}{2}\right]\right)A_{V_{0}},{(\gamma_{\natural\sharp\flat})}^{c}\left(\gamma_{\natural\sharp\flat}\left[\frac{b-1}{2}\right]\right)A_{U_{l}}\right]$
and in the latter case,
$K_{2(1+\mathfrak{m}+r)nc+b}=\left[{(\gamma_{\natural\sharp\flat})}^{c}\left(\gamma_{\natural\sharp\flat}\left[\frac{b-2}{2}\right]\right)A_{U_{l}},{(\gamma_{\natural\sharp\flat})}^{c}\left(\gamma_{\natural\sharp\flat}\left[\frac{b}{2}\right]\right)A_{V_{0}}\right].$
Then, Lemma 4.4 implies that, for any $i\in\{1,\ldots,k\}$,
$\cdots\subsetneq J_{i,-1}^{-}\subsetneq J_{i,0}^{-}\subsetneq J_{i,1}^{-}\subsetneq\cdots\subsetneq J_{i,2(1+\mathfrak{m}+r)n}^{-}\subsetneq J_{i,2(1+\mathfrak{m}+r)n+1}^{-}\subsetneq\cdots,$
$\cdots\supsetneq J_{i,-1}^{+}\supsetneq J_{i,0}^{+}\supsetneq J_{i,1}^{+}\supsetneq\cdots\supsetneq J_{i,2(1+\mathfrak{m}+r)n}^{+}\supsetneq J_{i,2(1+\mathfrak{m}+r)n+1}^{+}\supsetneq\cdots.$
Note that
$J_{i,0}^{-}\not\ni A_{V_{0}}\in J_{i,1}^{-},$
$J_{i,2(1+\mathfrak{m}+r)n}^{+}\ni\gamma_{\natural\sharp\flat}A_{V_{0}}\not\in J_{i,2(1+\mathfrak{m}+r)n+1}^{+}.$
The following is a variant for $C_{\Gamma,\Gamma_{0}}$ of [29, Lemma 6.4] for $C_{\Gamma}$.
Lemma 4.6.
The set $\{J_{i,d}\}_{i\in\{1,\ldots,k\},d\in\{1,\ldots,2(1+\mathfrak{m}+r)n\}}$ is the set of all hyperplanes separating $A_{V_{0}}$ and $\gamma_{\natural\sharp\flat}A_{V_{0}}$.
See Figure 8.
To prove the lemma, we consider a path $\ell$ from $A_{V_{0}}$ to $\gamma_{\natural\sharp\flat}A_{V_{0}}$ that diagonally penetrates each of the cubes $K_{1},\ldots,K_{2(1+\mathfrak{m}+r)n}$ in order. The set of all hyperplanes intersecting the path $\ell$ is $\{J_{i,d}\}_{i\in\{1,\ldots,k\},d\in\{1,\ldots,2(1+\mathfrak{m}+r)n\}}$.
We now state the final lemma required to complete the proof of Theorem 4.1. Recall Definition 2.7 and Remark 2.8 for sequences of separating hyperplanes. The following is a variant for $C_{\Gamma,\Gamma_{0}}$ of [29, Lemma 6.5] for $C_{\Gamma}$.
Lemma 4.7.
  1. (1)
    For any $i\in\{1,\ldots,k\}$, the sequence of hyperplanes
    $J_{i,1},\ldots,J_{i,2(1+\mathfrak{m}+r)n}$
    is a sequence of separating hyperplanes from $J_{i,0}$ to $J_{i,2(1+\mathfrak{m}+r)n+1}$, where $J_{i,0}$ and $J_{i,2(1+\mathfrak{m}+r)n+1}$ are of $v_{i,n}$-type and $v_{i,1}$-type, respectively (in particular, from $A_{V_{0}}$ to $\gamma_{\natural\sharp\flat}A_{V_{0}}$). See Figure 8.
  2. (2)
    1. (2-i)
      The sequence of hyperplanes
      $J_{1,1},J_{1,2},\ldots,J_{1,2n}$
      is a sequence of separating hyperplanes from
      $\gamma_{\natural\sharp\flat}[0]A_{V_{0}}=A_{V_{0}}\text{ to }\gamma_{\natural\sharp\flat}[n]A_{V_{0}}=\gamma_{\natural}A_{V_{0}}.$
    2. (2-ii)
      The sequence of hyperplanes
      $J_{1,2n+1},J_{1,2n+2},\ldots,J_{1,2(1+\mathfrak{m})n}$
      is a sequence of separating hyperplanes from
      $\gamma_{\natural\sharp\flat}[n]A_{V_{0}}=\gamma_{\natural}A_{V_{0}}\text{ to }\gamma_{\natural\sharp\flat}[(1+\mathfrak{m})n]A_{V_{0}}=\gamma_{\natural}\gamma_{\sharp}A_{V_{0}}.$
    3. (2-iii-$a$)
      For any $a\in\{1,\ldots,r\}$, the sequence of hyperplanes
      $J_{i_{a},2(1+\mathfrak{m})n+2(a-1)n+1},J_{i_{a},2(1+\mathfrak{m})n+2((a-1)n+1)},\ldots,$
      $J_{i_{a},2(1+\mathfrak{m})n+2((a-1)n+l_{\flat}(i_{a},i_{a+1}))-1},J_{i_{a+1},2(1+\mathfrak{m})n+2((a-1)n+l_{\flat}(i_{a},i_{a+1}))},$
      $J_{i_{a+1},2(1+\mathfrak{m})n+2((a-1)n+l_{\flat}(i_{a},i_{a+1}))+1},\ldots,J_{i_{a+1},2(1+\mathfrak{m}+a)n}$
      is a sequence of separating hyperplanes from
      $\gamma_{\natural\sharp\flat}[(1+\mathfrak{m}+a-1)n]A_{V_{0}}\text{ to }\gamma_{\natural\sharp\flat}[(1+\mathfrak{m}+a)n]A_{V_{0}}.$
      See Figure 9.
  3. (3)
    The sequence of hyperplanes composed of the sequences in (2-i), (2-ii), (2-iii-$1$), (2-iii-$2$), $\ldots$, (2-iii-$r$), in order, is a sequence of separating hyperplanes from $J_{1,0}$ to $J_{1,2(1+\mathfrak{m}+r)n+1}$, where $J_{1,0}$ and $J_{1,2(1+\mathfrak{m}+r)n+1}$ are of $v_{1,n}$-type and $v_{1,1}$-type, respectively (in particular, from $A_{V_{0}}$ to $\gamma_{\natural\sharp\flat}A_{V_{0}}$). Here, note that $i_{1}=1$. Moreover, the sequence contains a hyperplane of $v_{i}$-type for any $i\in\{1,\ldots,k\}$ and any $v_{i}\in V_{i}$.
Figure 9. The sequence of hyperplanes (2-iii-1) in Lemma 4.7 for the case where $\Gamma=\Gamma_{0}\ast\Gamma_{1}\ast\Gamma_{2}$. The two hyperplanes $J_{i_{1},2(1+\mathfrak{m})n+2l_{\flat}(i_{1},i_{2})-1}$ and $J_{i_{2},2(1+\mathfrak{m})n+2l_{\flat}(i_{1},i_{2})}$ do not intersect each other.
Proof.
(1) The assertion is clear from parts (1), (2), (3) and (4) of Lemma 4.4.
(2) Parts (1), (2), (3) and (4) of Lemma 4.4 and Lemma 4.5 imply that sequences in (2-i), (2-ii) and (2-iii-$a$) are sequences of separating hyperplanes.
(3) The sequence in the assertion is a sequence of separating hyperplanes. We need to show that the sequence contains a hyperplane of $v_{i}$-type for any $i\in\{1,\ldots,k\}$ and any $v_{i}\in V_{i}$. Recall that $(V_{i_{1}},\dots,V_{i_{r}},V_{i_{r+1}})$ (4.6) is a closed path on the tree $T$. Thus, if an edge $(V_{i},V_{j})\in E(T)$ is contained in the closed path, then so is the inverse edge $(V_{j},V_{i})$. In addition, because $T$ is a spanning tree of $Q(\Gamma)$ and the closed path $(V_{i_{1}},\dots,V_{i_{r}},V_{i_{r+1}})$ passes through every vertex at least once, for any $i\in\{1,\ldots,k\}$, $V_{i}$ is contained in the closed path as a vertex. Additionally, recall that $T$ has the root $V_{1}$ and that $i<j$ only if $V_{i}$ is not farther than $V_{j}$ from $V_{1}$ in $T$.
Now, take any $i\in\{1,\ldots,k\}$. We consider the two cases of $i=1$ and $i\neq 1$.
First, suppose that $i=1$. Note that $i_{1}=1$ and set $j=i_{2}$. Take $a\in\{1,2,\ldots,r\}$ such that $a\neq 1$, $i_{a}=1$, and $i_{a-1}=j$. Then, the sequence in the assertion contains the first half of the sequence of (2-iii-$1$):
$J_{i_{1},2(1+\mathfrak{m})n+1},J_{i_{1},2(1+\mathfrak{m})n+2},\ldots,J_{i_{1},2(1+\mathfrak{m})n+2l_{\flat}(i_{1},i_{2})-1},$
which are of $v_{1,1}$-type, $v_{1,1}$-type, $\ldots$, $v_{1,l_{\flat}(1,j)}$-type, and the second half of the sequence of (2-iii-$(a-1)$):
$J_{i_{a},2(1+\mathfrak{m})n+2((a-2)n+l_{\flat}(i_{a-1},i_{a}))},J_{i_{a},2(1+\mathfrak{m})n+2((a-2)n+l_{\flat}(i_{a-1},i_{a}))+1},\ldots,$
$J_{i_{a},2(1+\mathfrak{m})n+2(a-1)n},$
which are of $v_{1,l_{\flat}(j,1)}$-type, $v_{1,l_{\flat}(j,1)+1}$-type, $\ldots$, $v_{1,n}$-type. Note that $l_{\flat}(j,1)=l_{\flat}(1,j)$ by Lemma 4.2. Take any vertex $v_{1}\in V_{1}$. Then, there exists some $l\in\{1,\ldots,n\}$ such that $v_{1,l}=v_{1}$ by Lemma 4.2. Hence, the sequence in the assertion contains a hyperplane of $v_{1}$-type.
Next, suppose that $i\neq 1$. Take the smallest $a\in\{1,\ldots,r\}$ such that $i_{a}=i$. Because $i_{1}=1$ and $i\neq 1$, we have $a\neq 1$. We set $j=i_{a-1}$. Then, we have $a^{\prime}\in\{1,\ldots,r\}$ such that $a\leq a^{\prime}$, $i_{a^{\prime}}=i$, and $i_{a^{\prime}+1}=j$. The sequence in the assertion contains the second half of the sequence of (2-iii-$(a-1)$):
$J_{i_{a},2(1+\mathfrak{m})n+2((a-2)n+l_{\flat}(i_{a-1},i_{a}))},J_{i_{a},2(1+\mathfrak{m})n+2((a-2)n+l_{\flat}(i_{a-1},i_{a}))+1},\ldots,$
$J_{i_{a},2(1+\mathfrak{m})n+2(a-1)n},$
which are of $v_{i,l_{\flat}(j,i)}$-type, $v_{i,l_{\flat}(j,i)+1}$-type, $\ldots$, $v_{i,n}$-type, and the first half of the sequence of (2-iii-$a^{\prime}$):
$J_{i_{a^{\prime}},2(1+\mathfrak{m})n+2(a^{\prime}-1)n+1},J_{i_{a^{\prime}},2(1+\mathfrak{m})n+2((a^{\prime}-1)n+1)},\ldots,$
$J_{i_{a^{\prime}},2(1+\mathfrak{m})n+2((a^{\prime}-1)n+l_{\flat}(i_{a^{\prime}},i_{a^{\prime}+1}))-1},$
which are of $v_{i,1}$-type, $v_{i,1}$-type, $\ldots$, $v_{i,l_{\flat}(j,i)}$-type. Note that $l_{\flat}(j,i)=l_{\flat}(i,j)$ by Lemma 4.2. Take any vertex $v_{i}\in V_{i}$. Then, there exists some $l\in\{1,\ldots,n\}$ such that $v_{i,l}=v_{i}$ by Lemma 4.2. Hence, the sequence in the assertion contains a hyperplane of $v_{i}$-type. ∎
We can now complete the proof of Theorem 4.1.
Proof of $(1)\Rightarrow(2)$ in Theorem 4.1.
We show $(1)\Rightarrow(2)$ in Theorem 4.1. Consider $\gamma_{\natural\sharp\flat}\in A_{\Gamma}$ defined by (4.28) and hyperplanes
$J:=J_{1,0}\text{ and }J^{\prime}:=J_{1,2(1+\mathfrak{m}+r)n+1},$
which are of $v_{1,n}$-type and $v_{1,1}$-type, respectively. We now verify conditions (i), (ii) and (iii) in Theorem 2.9.
(i) $\gamma_{\natural\sharp\flat}$ skewers $(J,J^{\prime})$. Indeed, it is clear that
$J^{+}\supsetneq\gamma_{\natural\sharp\flat}^{-1}(J^{\prime+})\supsetneq\gamma_{\natural\sharp\flat}(J^{+})\supsetneq J^{\prime+}\supsetneq\gamma_{\natural\sharp\flat}^{2}(J^{+}).$
(ii) We show that $J$ and $J^{\prime}$ are strongly separated. Take any hyperplane $H$ with $J\cap H\neq\emptyset$. When $H$ is of $v_{i}$-type for some $i$ and $v_{i}\in V_{i}$, take a hyperplane $H^{\prime}$ of $v_{i}$-type separating $A_{V_{0}}$ and $\gamma_{\natural\sharp\flat}A_{V_{0}}$ such that $J\cap H^{\prime}=\emptyset$ by part (3) of Lemma 4.7. Then, $H\cap H^{\prime}=\emptyset$ by Remark 2.14. Because $J\subset H^{\prime-},H^{\prime+}\supset J^{\prime}$, $J\cap H\neq\emptyset$ and $H\cap H^{\prime}=\emptyset$, we have $J^{\prime}\cap H=\emptyset$.
(iii) We show $Stab(J)\cap Stab(J^{\prime})=\{1\}$. We divide the proof into three parts: (iii-1) we show that $Stab(J)\cap Stab(J^{\prime})\subset A_{V_{0}}$; (iii-2) we show that $Stab(J)\cap Stab(J^{\prime})\subset\gamma_{\natural\sharp\flat}A_{V_{0}}\gamma_{\natural\sharp\flat}^{-1}$; (iii-3) we show that $Stab(J)\cap Stab(J^{\prime})=\{1\}$ using (iii-1) and (iii-2).
(iii-1) Note that for any $i\in\{1,\ldots,k\}$ and any $v_{i}\in V_{i}$, we have at least one sequence of separating hyperplanes $P^{\prime}_{1},\ldots,P^{\prime}_{M^{\prime}}$ from $A_{V_{0}}$ to $\gamma_{\natural\sharp\flat}A_{V_{0}}$ such that $P^{\prime}_{1}$ is of $v_{i}$-type and $P^{\prime}_{M^{\prime}}$ is of $v_{1,n}$-type. For example, we can take such a sequence by considering a subsequence of the sequence in part (3) of Lemma 4.7. For any $i\in\{1,\ldots,k\}$ and any $v_{i}\in V_{i}$, we can take a longest sequence of separating hyperplanes $P_{1},\ldots,P_{M}$ from $A_{V_{0}}$ to $\gamma_{\natural\sharp\flat}A_{V_{0}}$ such that $P_{1}$ is of $v_{i}$-type and $P_{M}$ is of $v_{1,n}$-type, where
$A_{V_{0}}\in P_{1}^{-},P_{1}^{+}\supsetneq P_{2}^{+}\supsetneq\cdots\supsetneq P_{M-1}^{+}\supsetneq P_{M}^{+}\ni\gamma_{\natural\sharp\flat}A_{V_{0}}$
by taking decompositions into appropriate connected components
$C_{\Gamma,\Gamma_{0}}\!\setminus\!\!\setminus P_{1}=P_{1}^{-}\sqcup P_{1}^{+},\ldots,C_{\Gamma,\Gamma_{0}}\!\setminus\!\!\setminus P_{M}=P_{M}^{-}\sqcup P_{M}^{+}.$
Note that $P_{1},\ldots,P_{M}\in\{J_{j,d}\}_{j\in\{1,\ldots,k\},d\in\{1,\ldots,2(1+\mathfrak{m}+r)n\}}$ by Lemma 4.6. By noting $P_{M}\in\{J_{1,d}\}_{d\in\{1,\ldots,2(1+\mathfrak{m}+r)n\}}$ and applying part (1) of Lemma 4.7 for the case $i=1$, we have $P_{M}^{+}\supsetneq J^{\prime+}$.
Now assume that $Stab(J)\cap Stab(J^{\prime})\not\subset A_{V_{0}}$. Take $g\in Stab(J)\cap Stab(J^{\prime})$ with $g\not\in A_{V_{0}}$. Note that $g^{-1}\in Stab(J)\cap Stab(J^{\prime})$. Then, we have a hyperplane $H$ of $v_{i}$-type for some $i\in\{1,\ldots,k\}$ and some $v_{i}\in V_{i}$, separating $A_{V_{0}}$ and $gA_{V_{0}}$. We take a longest sequence of separating hyperplanes $P_{1},\ldots,P_{M}$ from $A_{V_{0}}$ to $\gamma_{\natural\sharp\flat}A_{V_{0}}$ such that $P_{1}$ is of $v_{i}$-type and $P_{M}$ is of $v_{1,n}$-type. Then we have
$g\gamma_{\natural\sharp\flat}A_{V_{0}},g^{-1}\gamma_{\natural\sharp\flat}A_{V_{0}}\in P_{M}^{+}$
by $P_{M}^{+}\supsetneq J^{\prime+}$. We take a connected component $H^{-}$ of $C_{\Gamma,\Gamma_{0}}\!\setminus\!\!\setminus H$ such that $A_{V_{0}}\in H^{-}$. Then the other connected component $H^{+}$ satisfies $gA_{V_{0}}\in H^{+}$. Because $H\cap J\neq\emptyset$, we have $H\cap J^{\prime}=\emptyset$ by (ii). Thus, $H$ cannot separate $\gamma_{\natural\sharp\flat}A_{V_{0}}$ and $g\gamma_{\natural\sharp\flat}A_{V_{0}}$. Hence we have either
(iii-1-$a$) $A_{V_{0}},\gamma_{\natural\sharp\flat}A_{V_{0}},g\gamma_{\natural\sharp\flat}A_{V_{0}}\in H^{-}\text{ and }gA_{V_{0}}\in H^{+}$ or
(iii-1-$b$) $A_{V_{0}}\in H^{-}\text{ and }gA_{V_{0}},\gamma_{\natural\sharp\flat}A_{V_{0}},g\gamma_{\natural\sharp\flat}A_{V_{0}}\in H^{+}$.
Assume that case (iii-1-$a$) occurs. Then $H$ does not separate $A_{V_{0}}$ and $\gamma_{\natural\sharp\flat}A_{V_{0}}$ and thus $H\notin\{J_{i,d}\}_{d\in\{1,\ldots,2(1+\mathfrak{m}+r)n\}}$ by Lemma 4.6. Hence, $H$ and $P_{1}$ are different hyperplanes of the same $v_{i}$-type, and thus they cannot intersect by Remark 2.14. Then we have either $H^{-}\supsetneq P_{1}^{+}$ or $H^{+}\supsetneq P_{1}^{+}$. However, $H^{+}\supsetneq P_{1}^{+}$ cannot occur because $\gamma_{\natural\sharp\flat}A_{V_{0}}\notin H^{+}$ and $\gamma_{\natural\sharp\flat}A_{V_{0}}\in P_{1}^{+}$. Therefore,
$H^{-}\supsetneq P_{1}^{+}.$
Then we have
$gA_{V_{0}}\in H^{+},H^{-}\supsetneq P_{1}^{+}\supsetneq P_{2}^{+}\supsetneq\cdots\supsetneq P_{M-1}^{+}\supsetneq P_{M}^{+}\ni g\gamma_{\natural\sharp\flat}A_{V_{0}}.$
Then, $Q_{1}=H,Q_{2}=P_{1},\ldots,Q_{M+1}=P_{M}$ is a sequence of separating hyperplanes from $gA_{V_{0}}$ to $g\gamma_{\natural\sharp\flat}A_{V_{0}}$ such that $Q_{1}$ is of $v_{i}$-type and $Q_{M+1}$ is of $v_{1,n}$-type. Thus, $g^{-1}Q_{1},g^{-1}Q_{2},\ldots,g^{-1}Q_{M+1}$ is a sequence of separating hyperplanes from $A_{V_{0}}$ to $\gamma_{\natural\sharp\flat}A_{V_{0}}$ such that $g^{-1}Q_{1}$ is of $v_{i}$-type and $g^{-1}Q_{M+1}$ is of $v_{1,n}$-type, which contradicts the fact that the sequence $P_{1},\ldots,P_{M}$ is the longest.
Next, assume that the case (iii-1-$b$) occurs. Then $H$ does not separate $gA_{V_{0}}$ and $g\gamma_{\natural\sharp\flat}A_{V_{0}}$; that is, $g^{-1}H$ does not separate $A_{V_{0}}$ and $\gamma_{\natural\sharp\flat}A_{V_{0}}$ and thus $g^{-1}H\notin\{J_{i,d}\}$ $({d\in\{1,\ldots,2(1+\mathfrak{m}+r)n\}})$ by Lemma 4.6. Hence, $g^{-1}H$ and $P_{1}$ are different hyperplanes of the same $v_{i}$-type, and thus they cannot intersect by Remark 2.14. Then we have either $g^{-1}H^{+}\supsetneq P_{1}^{+}$ or $g^{-1}H^{-}\supsetneq P_{1}^{+}$. However $g^{-1}H^{-}\supsetneq P_{1}^{+}$ can not occur because $\gamma_{\natural\sharp\flat}A_{V_{0}}\notin g^{-1}H^{-}$ and $\gamma_{\natural\sharp\flat}A_{V_{0}}\in P_{1}^{+}$. Therefore
$g^{-1}H^{+}\supsetneq P_{1}^{+}.$
Then we have
$g^{-1}A_{V_{0}}\in g^{-1}H^{-},g^{-1}H^{+}\supsetneq P_{1}^{+}\supsetneq P_{2}^{+}\supsetneq\cdots\supsetneq P_{M-1}^{+}\supsetneq P_{M}^{+}\ni g^{-1}\gamma_{\natural\sharp\flat}A_{V_{0}}.$
Then, $Q_{1}=H,Q_{2}=gP_{1},\ldots,Q_{M+1}=gP_{M}$ is a sequence of separating hyperplanes from $A_{V_{0}}$ to $\gamma_{\natural\sharp\flat}A_{V_{0}}$ such that $Q_{1}$ is of $v_{i}$-type and $Q_{M+1}$ is of $v_{1,n}$-type, which contradicts the fact that the sequence $P_{1},\ldots,P_{M}$ is the longest.
We now see that $Stab(J)\cap Stab(J^{\prime})\subset A_{V_{0}}$.
(iii-2) We show $Stab(J)\cap Stab(J^{\prime})\subset\gamma_{\natural\sharp\flat}A_{V_{0}}\gamma_{\natural\sharp\flat}^{-1}$. This can be shown in a similar way as (iii-1). Indeed we reach a contradiction by assuming $Stab(J)\cap Stab(J^{\prime})\not\subset\gamma_{\natural\sharp\flat}A_{V_{0}}\gamma_{\natural\sharp\flat}^{-1}$.
[…]
Suppose that $g\in A_{V_{0}\setminus\{w^{\prime}\in V_{0}\mid w^{\prime}\prec w\}}=A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}$. We show that $g\in A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succ w\}}$.
Figure 10. Cubes penetrated by $\ell$ when $k=3$ and $l_{\sharp}(\mathfrak{j})=1$. In this case, $g$ fixes cubes $[A_{V_{0}},A_{U_{1}}]$ and $[v_{1,1}v_{2,1}v_{3,1}A_{V_{0}},A_{U_{1}}]$, which are penetrated by $\ell$. Then $g$ fixes three edges $[A_{V_{0}\sqcup\{v_{1,1},v_{2,1}\}},A_{U_{1}}]$, $[A_{V_{0}\sqcup\{v_{1,1},v_{3,1}\}},A_{U_{1}}]$ and $[v_{1,1}A_{V_{0}\sqcup\{v_{2,1},v_{3,1}\}},A_{U_{1}}]$, and thus three squares $[A_{V_{0}\sqcup\{v_{1,1}\}},A_{U_{1}}]$, $[v_{1,1}A_{V_{0}\sqcup\{v_{2,1}\}},A_{U_{1}}]$ and $[v_{1,1}A_{V_{0}\sqcup\{v_{3,1}\}},A_{U_{1}}]$. Hence $g$ fixes the cube $[v_{1,1}A_{V_{0}},A_{U_{1}}]$. By a similar argument, $g$ fixes the cube $[v_{1,1}v_{2,1}A_{V_{0}},A_{U_{1}}]$. Therefore, $g$ fixes not only $A_{V_{0}}$ and $v_{1,1}v_{2,1}v_{3,1}A_{V_{0}}$, but also $v_{1,1}A_{V_{0}}$ and $v_{1,1}v_{2,1}A_{V_{0}}$.
(iii-3-$a$) We consider the case where $h=1$; that is, $w\in W(1)$. Then we have $v_{i_{\sharp}(w),l_{\sharp}(w)}\in V_{\ast}$. We consider a vertex on $\ell$ of the form
(4.35) $\lambda_{1}\cdots\lambda_{l_{\sharp}(w)-1}\lambda_{l_{\sharp}(w)}A_{V_{0}},$
[…]
$\lambda_{l_{\sharp}(w)}=v_{1,l_{\sharp}(w)}\cdots v_{i_{\sharp}(w),l_{\sharp}(w)}\cdots v_{k,l_{\sharp}(w)}.$
Since the action of $A_{\Gamma}$ on $C_{\Gamma,\Gamma_{0}}$ preserves the cube structure, all cubes penetrated by $\ell$ are fixed by $g$. Therefore, $g$ fixes
$\lambda_{1}\cdots\lambda_{l_{\sharp}(w)-1}v_{1,l_{\sharp}(w)}\cdots v_{i_{\sharp}(w),l_{\sharp}(w)}A_{V_{0}},$
see Figure 10. Then
$g\in(\lambda_{1}\cdots\lambda_{l_{\sharp}(w)-1}v_{1,l_{\sharp}(w)}\cdots v_{i_{\sharp}(w),l_{\sharp}(w)})A_{V_{0}}(\lambda_{1}\cdots\lambda_{l_{\sharp}(w)-1}v_{1,l_{\sharp}(w)}\cdots v_{i_{\sharp}(w),l_{\sharp}(w)})^{-1}.$
By $g\in A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}$ and Lemma 4.3 (1), $\lambda_{1}\cdots\lambda_{l_{\sharp}(w)-1}v_{1,l_{\sharp}(w)}\cdots v_{i_{\sharp}(w)-1,l_{\sharp}(w)}$ commutes with $g$, see Figure 7. Therefore, $g\in v_{i_{\sharp}(w),l_{\sharp}(w)}A_{V_{0}}v_{i_{\sharp}(w),l_{\sharp}(w)}^{-1}$, and thus $g\in A_{V_{0}}\cap v_{i_{\sharp}(w),l_{\sharp}(w)}A_{V_{0}}v_{i_{\sharp}(w),l_{\sharp}(w)}^{-1}$. According to Lemma 2.11, $A_{V_{0}}\cap v_{i_{\sharp}(w),l_{\sharp}(w)}A_{V_{0}}v_{i_{\sharp}(w),l_{\sharp}(w)}^{-1}\subset A_{V_{0}\setminus\{w\}}$. Hence
$g\in A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}\cap A_{V_{0}\setminus\{w\}}=A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succ w\}}.$
(iii-3-$b$) We consider the case where $w\in W(h)$ where $h\geq 2$. Then we have
$(w(h),w(h-1),\ldots w(0))=(w,\pi(w),\ldots\pi^{h}(w))$
[…]
(4.36) $\gamma_{\natural}\gamma_{\sharp}(1)\cdots\gamma_{\sharp}(\mathfrak{j}-1)\lambda_{1}\cdots\lambda_{l_{\sharp}(w_{\mathfrak{j}})-1}\lambda_{\sharp}(\mathfrak{j})A_{V_{0}},$
which is fixed by $g$. We note that $\lambda_{\sharp}(\mathfrak{j})$ equals
$v_{1,l_{\sharp}(w_{\mathfrak{j}})}\cdots v_{i_{\sharp}(w_{\mathfrak{j}})-1,l_{\sharp}(w_{\mathfrak{j}})}v_{i_{\sharp}(w_{\mathfrak{j}})+1,l_{\sharp}(w_{\mathfrak{j}})}\cdots v_{k,l_{\sharp}(w_{\mathfrak{j}})}w_{\mathfrak{j}}(h-1)\cdots w_{\mathfrak{j}}(1)w_{\mathfrak{j}}(0)$
$=v_{1,l_{\sharp}(w_{\mathfrak{j}})}\cdots v_{i_{\sharp}(w_{\mathfrak{j}})-1,l_{\sharp}(w_{\mathfrak{j}})}v_{i_{\sharp}(w_{\mathfrak{j}})+1,l_{\sharp}(w_{\mathfrak{j}})}\cdots v_{k,l_{\sharp}(w_{\mathfrak{j}})}w(h-1)\cdots w(1)w(0).$
Hence
$\gamma_{\natural}\gamma_{\sharp}(1)\cdots\gamma_{\sharp}(\mathfrak{j}-1)\lambda_{1}\cdots\lambda_{l_{\sharp}(w_{\mathfrak{j}})-1}\lambda_{\sharp}(\mathfrak{j})$
$=$ $\left(\gamma_{\natural}\gamma_{\sharp}(1)\cdots\gamma_{\sharp}(\mathfrak{j}-1)\lambda_{1}\cdots\lambda_{l_{\sharp}(w_{\mathfrak{j}})-1}\right)\cdot$
$\left(v_{1,l_{\sharp}(w_{\mathfrak{j}})}\cdots v_{i_{\sharp}(w_{\mathfrak{j}})-1,l_{\sharp}(w_{\mathfrak{j}})}v_{i_{\sharp}(w_{\mathfrak{j}})+1,l_{\sharp}(w_{\mathfrak{j}})}\cdots v_{k,l_{\sharp}(w_{\mathfrak{j}})}w(h-1)\cdots w(1)w(0)\right).$
Note that $\gamma_{\natural}\gamma_{\sharp}(1)\cdots\gamma_{\sharp}(\mathfrak{j}-1)$ is a product of elements of $V_{\ast}\sqcup\{w_{1},\ldots w_{\mathfrak{j}-1}\}$. By Lemma 4.3 (2) with $w_{1}\prec\cdots\prec w_{\mathfrak{j}-1}\prec w_{\mathfrak{j}}=\pi(w)$, any $w^{\prime\prime}\succeq w$ commutes with
$\gamma_{\natural}\gamma_{\sharp}(1)\cdots\gamma_{\sharp}(\mathfrak{j}-1)\lambda_{1}\cdots\lambda_{l_{\sharp}(w_{\mathfrak{j}})-1}v_{1,l_{\sharp}(w_{\mathfrak{j}})}\cdots v_{i_{\sharp}(w_{\mathfrak{j}})-1,l_{\sharp}(w_{\mathfrak{j}})}v_{i_{\sharp}(w_{\mathfrak{j}})+1,l_{\sharp}(w_{\mathfrak{j}})}\cdots v_{k,l_{\sharp}(w_{\mathfrak{j}})}.$
Hence so does $g$. Then $g\in(w(h-1)\cdots w(0))A_{V_{0}}(w(h-1)\cdots w(0))^{-1}$. Next, by noting $w(0)\prec w(1)\prec\cdots\prec w(h-1)\prec w(h)=w$, we can show $g\in A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succ w\}}$. Indeed, it follows from Lemma 2.11 that $g$ belongs to
$A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}$
$\cap(w(h-1)\cdots w(0))A_{V_{0}}(w(h-1)\cdots w(0))^{-1}$
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}\cap A_{V_{0}}$
$\cap(w(h-1)\cdots w(1)w(0))A_{V_{0}}(w(h-1)\cdots w(1)w(0))^{-1}$
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}\cap(w(h-1)\cdots w(1))A_{V_{0}}(w(h-1)\cdots w(1))^{-1}$
$\cap(w(h-1)\cdots w(1)w(0))A_{V_{0}}(w(h-1)\cdots w(1)w(0))^{-1}$
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}$
$\cap(w(h-1)\cdots w(1))(A_{V_{0}}\cap w(0)A_{V_{0}}w(0)^{-1})(w(h-1)\cdots w(1))^{-1}$
$\subset$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}$
$\cap(w(h-1)\cdots w(1))A_{V_{0}\setminus\{w(1)\}}(w(h-1)\cdots w(1))^{-1}$
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}\cap A_{V_{0}\setminus\{w(1)\}}$
$\cap(w(h-1)\cdots w(2)w(1))A_{V_{0}\setminus\{w(1)\}}(w(h-1)\cdots w(2)w(1))^{-1}$
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}\cap(w(h-1)\cdots w(2))A_{V_{0}\setminus\{w(1)\}}(w(h-1)\cdots w(2))^{-1}$
$\cap(w(h-1)\cdots w(2)w(1))A_{V_{0}\setminus\{w(1)\}}(w(h-1)\cdots w(2)w(1))^{-1}$
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}$
$\cap(w(h-1)\cdots w(2))(A_{V_{0}\setminus\{w(1)\}}\cap w(1)A_{V_{0}\setminus\{w(1)\}}w(1)^{-1})(w(h-1)\cdots w(2))^{-1}$
$\subset$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}$
$\cap(w(h-1)\cdots w(2))A_{V_{0}\setminus\{w(1),w(2)\}}(w(h-1)\cdots w(2))^{-1}$
$\cdots$
$\subset$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}$
$\cap w(h-1)A_{V_{0}\setminus\{w(1),\ldots,w(h-1)\}}w(h-1)^{-1}$
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}$
$\cap(A_{V_{0}\setminus\{w(1),\ldots,w(h-1)\}}\cap w(h-1)A_{V_{0}\setminus\{w(1),\ldots,w(h-1)\}}w(h-1)^{-1})$
$\subset$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}$
$\cap A_{V_{0}\setminus\{w(1),\ldots,w(h-1),w(h)\}}$
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succ w\}}.$
In the above deformation, all $\subset$ are consequences of Lemma 2.11, and the final equality follows from $w(h)=w$. The other equalities are trivial deformations.
Since $\prec$ is a total order on a finite set $V_{0}$, by applying (iii-3-$a$) or (iii-3-$b$) for $w\in V_{0}$ in order from the smallest element of $V_{0}$ to the largest one, we have $g\in A_{\emptyset}=\{e\}$ and thus $g=e$. ∎
Proof of Corollaries 1.4 and 1.5.
Let $A_{\Gamma}$ be an irreducible Artin group associated with $\Gamma$ that is not a clique. Then, Theorem 1.2 implies that $A_{\Gamma}$ is acylindrically hyperbolic. Therefore, the center of $A_{\Gamma}$ is finite by acylindrical hyperbolicity ([33, Corollary 7.3]). Proposition 3.8 then implies Corollary 1.4. We can show Corollary 1.5 in a similar way. ∎
Remark 4.8.
There are several versions of the so-called center conjecture. Note that the center of any irreducible Artin group of spherical type is known to be infinite cyclic [4], [17]. Then, as one version, the center conjecture containing this known fact can be stated as follows: the center of any irreducible Artin group of spherical type and of infinite type is infinite cyclic and trivial, respectively. Since the center of any possibly reducible Artin group with each irreducible factor satisfying the center conjecture is trivial or torsion-free, the last part of Corollary 1.4 clearly implies the following: in the setting of Corollary 1.4, if each irreducible factor of $A_{\Gamma_{0}}$ satisfies the center conjecture, then $A_{\Gamma}$ satisfies the center conjecture. This claim is independently given in [27] by a different method.
5. Example
Let $\Gamma$ be the graph with $V(\Gamma)=\{x,y,u_{1},u_{2},v_{1},v_{2}\}$ as shown in Figure 1. In this section, we explain the construction of $\gamma_{\natural\sharp\flat}$ for this example.
Remark 5.1.
Theorems 1.2 and 4.1 can be applied to Artin groups associated with $\Gamma$ in Figures 4 and 5. However, these examples may be too simple to illustrate the construction of $\gamma_{\natural\sharp\flat}$. Indeed, in the leftmost defining graph in Figure 4, $\gamma_{\flat}=e\in A_{\Gamma}$. For the middle and right $\Gamma$, $\gamma_{\sharp}=\gamma_{\flat}=e\in A_{\Gamma}$. Also, for all $\Gamma$ in Figure 5, $\gamma_{\sharp}=e\in A_{\Gamma}$. The graph $\Gamma$ in Figure 1 is one of the smallest that has non-trivial $\gamma_{\sharp}$ and $\gamma_{\flat}$.
We note that $\Gamma$ is irreducible; that is, $\Gamma^{t}$ is connected. Then we have the join decomposition $\Gamma=\Gamma_{0}\ast\Gamma_{\ast}=\Gamma_{0}\ast(\Gamma_{1}\ast\Gamma_{2})$, where we set subgraphs $\Gamma_{i}$ $(i=0,1,2)$ of $\Gamma$ by
(5.1) $V_{0}:=V(\Gamma_{0})=\{x,y\},\ E_{0}:=E(\Gamma_{0})=\{(x,y),\ (y,x)\},$
(5.2) $V_{1}:=V(\Gamma_{1})=\{u_{1},u_{2}\},\ E_{1}:=E(\Gamma_{1})=\emptyset,$
(5.3) $V_{2}:=V(\Gamma_{2})=\{v_{1},v_{2}\},\ E_{2}:=E(\Gamma_{2})=\emptyset.$
Clearly, $\Gamma_{0}$ is the maximal clique factor, and $\Gamma_{1}$ and $\Gamma_{2}$ are join-indecomposable factors. Note that $k=2$. We give an explicit construction of $\gamma_{\natural\sharp\flat}=\gamma_{\natural}\gamma_{\sharp}\gamma_{\flat}$ for this $\Gamma$ following Section 4.
First, we construct $\gamma_{\natural}$. In our $\Gamma$, $Q(\Gamma,\Gamma_{0})$ is the tree consisting of one edge and two vertices $V_{1}$ and $V_{2}$ (see (LABEL:QGamma)). Here, the spanning tree $T$ coincides with $Q(\Gamma,\Gamma_{0})$. We choose $V_{1}$ as the root of $T$. Take paths $p_{1,2}$ and $p_{2,1}$ in $\Gamma^{t}$ as follows:
(5.4) $p_{1,2}=(w_{1,2,0},w_{1,2,1},w_{1,2,2})=(u_{1},x,v_{1}),$
(5.5) $p_{2,1}=(w_{2,1,0},w_{2,1,1},w_{2,1,2})=(v_{1},x,u_{1})$
(see (4.2) and (4.3)). Their end vertices are
(5.6) $s_{1,2}=w_{1,2,0}=u_{1}\in V_{1},\ t_{1,2}=w_{1,2,2}=v_{1}\in V_{2},$
(5.7) $s_{2,1}=w_{2,1,0}=v_{1}\in V_{2},\ t_{2,1}=w_{2,1,2}=u_{1}\in V_{1},$
respectively (see (4.4)). Set $\tau_{1,2},\tau_{2,1}\in A_{\Gamma}$ as
(5.8) $\tau_{1,2}=w_{1,2,0}w_{1,2,1}w_{1,2,2}=u_{1}xv_{1},$
(5.9) $\tau_{2,1}=w_{2,1,0}w_{2,1,1}w_{2,1,2}=v_{1}xu_{1}$
(see (4.5)). We fix a closed path on $T=Q(\Gamma,\Gamma_{0})$;
(5.10) $(V_{i_{1}},V_{i_{2}},V_{i_{3}})=(V_{1},V_{2},V_{1})$
(see (4.6)). We note that the length $r$ of the path is $2$.
According to Lemma 4.2, we construct $n\in\mathbb{N}$, $M_{\natural}\in\mathcal{M}(2,n;V(\Gamma))$ and a map $l_{\flat}$. Take closed paths of length $n=2$ on $\Gamma_{1}^{c}$, $\Gamma_{2}^{c}$:
(5.11) $(u_{1},u_{2},u_{1}),\ (v_{1},v_{2},v_{1}),respectively.$
Set
(5.12) $M_{\natural}=\begin{pmatrix}u_{1}&u_{2}\\ v_{1}&v_{2}\end{pmatrix},$
and
(5.13) $l_{\flat}(1,2)=l_{\flat}(2,1)=1.$
Then,
(5.14) $M_{\natural}[1,l_{\flat}(1,2)]=u_{1},\ M_{\natural}[2,l_{\flat}(2,1)]=v_{1}.$
We have
(5.15) $\gamma_{\natural}=u_{1}v_{1}u_{2}v_{2},$
[…]
(5.16) $M^{\prime}_{\flat}(1)=M^{\prime}_{\flat}(2)=\begin{pmatrix}u_{1}^{-1}&e\\ v_{1}^{-1}&e\end{pmatrix},$
[…]
(5.17) $M^{\prime\prime}_{\flat}(1)=\begin{pmatrix}\tau_{1,2}&e\\ e&e\end{pmatrix}=\begin{pmatrix}u_{1}xv_{1}&e\\ e&e\end{pmatrix}$
and
(5.18) $M^{\prime\prime}_{\flat}(2)=\begin{pmatrix}\tau_{2,1}&e\\ e&e\end{pmatrix}=\begin{pmatrix}v_{1}xu_{1}&e\\ e&e\end{pmatrix},$
[…]
(5.19) $M_{\flat}(1)=\begin{pmatrix}u_{1}xv_{1}&u_{2}\\ e&v_{2}\end{pmatrix},$
and
(5.20) $M_{\flat}(2)=\begin{pmatrix}v_{1}xu_{1}&u_{2}\\ e&v_{2}\end{pmatrix},$
[…]
(5.21) $M_{\flat}=\begin{pmatrix}u_{1}xv_{1}&u_{2}&v_{1}xu_{1}&u_{2}\\ e&v_{2}&e&v_{2}\end{pmatrix},$
[…]
(5.22) $\gamma_{\flat}=(u_{1}xv_{1})(u_{2}v_{2})(v_{1}xu_{1})(u_{2}v_{2}),$
[…]
(5.23) $W(1)=\{x\},\ W(2)=\{y\}.$
Note that $d_{\Gamma^{t}}(x,V_{\ast})=1$, $d_{\Gamma^{t}}(y,V_{\ast})=2$. Then $V_{0}=W(1)\sqcup W(2)$ and $\mathfrak{h}=2$. Here, $d_{\Gamma^{t}}(x,V_{\ast})=d_{\Gamma^{t}}(x,u_{1})=d_{\Gamma^{t}}(x,v_{1})$. Therefore,
(5.24) $(i_{\sharp}(x),l_{\sharp}(x))=(i_{\sharp}(y),l_{\sharp}(y))=(1,1)$
[…]
(5.25) $w_{1}=w_{\mathfrak{m}}=x$
where
(5.26) $\mathfrak{m}=\#\left(W(1)\sqcup\cdots\sqcup W(\mathfrak{h}-1)\right)=\#W(1)=1$
[…]
(5.27) $l_{\sharp}(w_{1})=l_{\sharp}(x)=1.$
[…]
(5.28) $M^{\prime}_{\sharp}(1)=\begin{pmatrix}u_{1}^{-1}&e\\ e&e\end{pmatrix},$
[…]
(5.29) $M^{\prime\prime}_{\sharp}(1)=\begin{pmatrix}e&e\\ w_{1}(1)w_{1}(0)&e\end{pmatrix}=\begin{pmatrix}e&e\\ xu_{1}&e\end{pmatrix},$
[…]
(5.30) $M_{\sharp}(1)=\begin{pmatrix}e&u_{2}\\ v_{1}xu_{1}&v_{2}\end{pmatrix},$
[…]
(5.31) $M_{\sharp}=M_{\sharp}(1)=\begin{pmatrix}e&u_{2}\\ v_{1}xu_{1}&v_{2}\end{pmatrix},$
[…]
(5.32) $\gamma_{\sharp}=(v_{1}xu_{1})(u_{2}v_{2}),$
[…]
Then
(5.33) $\gamma_{\natural\sharp\flat}=\gamma_{\natural}\gamma_{\sharp}\gamma_{\flat}=(u_{1}v_{1}u_{2}v_{2})(v_{1}xu_{1}u_{2}v_{2})(u_{1}xv_{1}u_{2}v_{2}v_{1}xu_{1}u_{2}v_{2})$
[…]
Acknowledgments
The authors would like to thank Antoine Goldsborough, Nicolas Vaskou, Kasia Jankiewicz and MurphyKate Montee for useful conversations. Especially the authors thank Antoine Goldsborough and Nicolas Vaskou for their pointing out to us that Theorem 1.2 has an application to random Artin groups.